### v2

# A Systematic Literature Review on Detecting Software Vulnerabilities with Large Language Models

Abstract.
The increasing adoption of Large Language Models (LLMs) in software engineering has sparked interest in their use for software vulnerability detection. However, the rapid development of this field has resulted in a fragmented research landscape, with diverse studies that are difficult to compare due to differences in, e.g., system designs and dataset usage. This fragmentation makes it difficult to obtain a clear overview of the state-of-the-art or compare and categorize studies meaningfully. In this work, we present a comprehensive systematic literature review (SLR) of LLM-based software vulnerability detection. We analyze 263 studies published between January 2020 and November 2025, categorizing them by task formulation, input representation, system architecture, and techniques. Further, we analyze the datasets used, including their characteristics, vulnerability coverage, and diversity. We present a fine-grained taxonomy of vulnerability detection approaches, identify key limitations, and outline actionable future research opportunities. By providing a structured overview of the field, this review improves transparency and serves as a practical guide for researchers and practitioners aiming to conduct more comparable and reproducible research. We publicly release all artifacts and maintain a living repository of LLM-based software vulnerability detection studies at https://github.com/hs-esslingen-it-security/Awesome-LLM4SVD.
1. Introduction
Vulnerability detection plays a critical role in the software development life cycle, identifying security vulnerabilities before they can be exploited in deployed systems. The growing complexity of modern software, combined with an ever-increasing threat landscape, has led to a surge in reported vulnerabilities. The Common Vulnerabilities and Exposures (CVE) records (CVE, 2025b) provide unique identifiers for publicly known software vulnerabilities. Over 40,000 CVEs have been published in 2024 alone, with over 12,000 reported in the first quarter of 2025 (CVE, 2025a). Complementing CVEs, the Common Weakness Enumeration (CWE) (MITRE, 2025a) offers a hierarchical classification of vulnerability types, serving as a higher-level abstraction for understanding underlying vulnerability patterns.
Vulnerability assessment pipelines as found in industrial development workflows typically consist of multiple steps: detection (identifying whether code contains a vulnerability), localization (pinpointing the exact location in the code), severity estimation (evaluating potential impact), and repair (patching the code to mitigate the vulnerability), often initiated reactively when a new CVE is disclosed. Despite advances in tooling, e.g., static application security testing (SAST) tools, these processes remain largely manual or only partially automated due to the dependency on expert-defined rules (cf. Taghavi and Feyzi (Taghavi Far and Feyzi, 2025) for a comparison between LLMs and established software vulnerability detection methods). Further, these tools suffer from high false positive rates (Croft et al., 2023). As such, existing workflows are insufficient to meet the scale of remediation demands.
At the same time, the adoption of Large Language Models (LLMs)111Due to the lack of a formal definition for LLMs, in this work, we use the term LLM broadly to encompass both large-scale language models (e.g., GPT-4) and smaller pre-trained models, such as BERT and CodeBERT (Zhao et al., 2024b). We differentiate architectures and scales of LLMs later on in the body of this work. for software engineering tasks has grown rapidly since the introduction of BERT (Devlin et al., 2019) in 2018. LLMs are deep neural network models, typically based on Transformer architectures (Vaswani et al., 2017), that are pre-trained on massive corpora of natural language. LLMs have demonstrated impressive capabilities in understanding and generating code. Industry forecasts further emphasize the transformative potential of LLMs in code generation (Amodei, 2025).
However, while LLMs hold promise in code generation, they also introduce new risks. Generated code often lacks awareness of existing libraries or internal codebases, which may lead to redundancies and the creation of syntactically different but semantically equivalent code. This complicates the detection of vulnerabilities, especially when they are subtle or semantically nuanced. Additionally, LLMs are prone to hallucinations and the generation of insecure code (Pearce et al., 2025; Liu et al., 2024d, c; Fu et al., 2025a). Development teams may be unaware of vulnerabilities introduced by generated code, whether in their codebase, third-party components, or libraries.
The combination of LLM-generated code and the increasing volume of vulnerabilities calls for automated, scalable, and reliable vulnerability assessment pipelines. LLMs’ ability to generalize across programming languages and representations, as well as their high flexibility through prompt adaptation, make them a promising approach for software vulnerability assessment. Therefore, recent research has begun to explore the use of LLMs. In a prior study (Kaniewski et al., 2024), we surveyed early work on LLM-based vulnerability handling, including detection, localization, and repair, and highlighted open challenges with regard to scalability, data diversity, and model reliability. We continued to extensively document the progress in this field, with a focus on vulnerability detection studies as a prerequisite for all subsequent remediation steps. As the number of studies on LLM-based vulnerability detection continues to grow, we observe a wide variety of system architectures, adaptation techniques, and evaluation methodologies. This heterogeneity leads to two pressing challenges: First, there is currently no comprehensive survey that maps the landscape of LLM-based software vulnerability detection methods, their system designs, and dataset usage. As a result, researchers face difficulties identifying trends, gaps, or best practices in this rapidly evolving area. Second, studies frequently adopt custom datasets, metrics, and data splits, making it difficult to assess progress, reproduce results, or compare models fairly. To address these challenges, we contribute this systematic literature review (SLR) of LLM-based software vulnerability detection studies. Specifically, our contributions are as follows:
We conduct a SLR of studies covering LLM-based software vulnerability detection in the time from January 2020 to November 2025, covering 263 studies.
We present a comprehensive taxonomy that categorizes existing approaches based on task formulation (classification and generation), input representation (code and auxiliary information), system architecture (LLM usage), adaptation and orchestration technique, as well as dataset usage. This taxonomy allows for meaningful comparison of studies.
We perform an in-depth analysis of commonly used and selected datasets, investigating class balance, CWE coverage as well as diversity, and limitations. Based on this analysis, we provide actionable insights and best practices for future dataset selection and evaluation design to improve cross-study comparability.
We discuss overarching limitations and challenges of current LLM-based vulnerability detection approaches and outline future research opportunities.
To support the community, we provide the artifacts and maintain a regularly updated list of reviewed studies at https://github.com/hs-esslingen-it-security/Awesome-LLM4SVD (Awesome-LLM4SVD) (Kaniewski et al., 2025).
The remainder of this paper is structured as follows: We review related surveys in Section 2, and describe the SLR methodology in Section 3. Section 4 presents the comprehensive taxonomy for LLM-based vulnerability detection. We extend this taxonomy and analyze datasets used in Section 5. We discuss current limitations and future research opportunities in Section 6. Finally, we address limitations of this SLR in Section 7 and conclude this review in Section 8.
2. Related Work
With the growing adoption of LLMs across various domains, numerous works have emerged to explore their applications. In this section, we discuss related works, covering related SLRs and surveys that intersect most closely with the topic of LLM-based software vulnerability detection, as well as works focusing on vulnerability datasets.
2.1. Related SLRs and Surveys
In Table 1, we provide an overview of related surveys, with particular attention to their coverage of the main aspects of this SLR, i.e., focus on LLM-based studies, input representations, adaptation techniques, analysis of datasets, and a comprehensive taxonomy. In the following, we discuss these surveys and how this SLR differs in coverage from previous works.
Several surveys analyze the application of LLMs in the domain of cyber security (Xu et al., 2025; Ferrag et al., 2025a), addressing vulnerability detection as one of several sub-tasks. Xu et al. (Xu et al., 2025), for example, cover the domains of software and system security, network security, information and content security, hardware security, and blockchain security. The authors focus on what LLMs have been employed to support these security tasks, and discuss domain-specific adaptation techniques, i.e., fine-tuning, prompt engineering, and external augmentation. Further, they address dataset collection and pre-processing techniques. Ferrag et al. (Ferrag et al., 2025a) focus on LLMs employed for cyber security and their optimization strategies, briefly discussing advanced training and fine-tuning techniques. The authors also provide an overview of software vulnerability datasets used for fine-tuning.
Surveys in the domain of software engineering (Hou et al., 2024; Zhang et al., 2024a; Fu et al., 2025b) categorize LLM usage by phases of the software development life cycle. Vulnerability detection is typically part of software testing (Zhang et al., 2024a) or software quality assurance (Hou et al., 2024). Fu et al. (Fu et al., 2025b) focus on AI for DevSecOps, covering software vulnerability detection and classification. Due to the broader scope of AI, the survey covers fewer studies applying LLM. Zhang et al. (Zhang et al., 2024a) focus on the LLM perspective, providing a comprehensive overview of code LLMs, discussing pre-training tasks, and downstream applications. Hou et al. (Hou et al., 2024) discuss techniques used to optimize LLMs, including prompt engineering and fine-tuning techniques. The authors also provide an overview of input forms used for diverse software engineering tasks (i.e., token- and tree/graph-based) and highlight data collection, selection, and pre-processing processes. While these surveys provide overviews of the application of LLMs in broader domains, we focus on insights into the state-of-the-art adaptation and orchestration techniques, input representations, and datasets that are specific to software vulnerability detection.
| Coverage | ||||||||||
| Reference | Published | Focus | Time Frame | #Studies (all — SVD) |
SLR |
LLM Focus |
Input Repres. |
Adaptation Techniques |
Dataset Analysis |
Taxonomy |
| Cyber Security | ||||||||||
| Xu et al. (Xu et al., 2025) | Sep 2025 | Cyber security | 2020-2025 | 185 — 22 | ✓ | ● | ○ | ● | ◐ | ○ |
| Ferrag et al. (Ferrag et al., 2025a) | May 2025 | Cyber security | n/s | n/s — 9 | ✗ | ● | ○ | ◐ | ◐ | ○ |
| Software Engineering | ||||||||||
| Zhang et al. (Zhang et al., 2024a) | Sep 2024 | Software engineering | 2020-2024 | 947 — 71 | ✓ | ● | ○ | ○ | ○ | ○ |
| Hou et al. (Hou et al., 2024) | Dec 2024 | Software engineering | 2017-2024 | 395 — 18 | ✓ | ● | ● | ● | ◐ | ○ |
| Fu et al. (Fu et al., 2025b) | Apr 2025 | DevSecOps | 2017-2023 | 99 — 9 | ✓ | ○ | ○ | ○ | ◐* | ○ |
| Vulnerability Detection | ||||||||||
| Shiri Harzevili et al. (Shiri Harzevili et al., 2024) | Nov 2024 | Autom. software vulnerability detection | 2011-2024 | 138 — 7 | ✓ | ○ | ● | ○ | ◐ | ○ |
| Shereen et al. (Shereen et al., 2024) | Dec 2024 | Autom. vulnerability detection | 2006-2024 | 79 — 10 | ✗ | ○ | ◐ | ○ | ◐* | ○ |
| Shimmi et al. (Shimmi et al., 2025) | Jun 2025 | Software vulnerability detection | 2018-2023 | 98 — 13 | ✓ | ○ | ● | ○ | ◐* | ● |
| Taghavi Far and Feyzi (Taghavi Far and Feyzi, 2025) | Feb 2025 | Software vulnerability detection | n/s | n/s — n/s | ✓ | ● | ○ | ◐ | ◐ | ○ |
| Basic and Giaretta (Basic and Giaretta, 2025) | Apr 2025 | Code security | n/s | n/s — n/s | ✓ | ● | ○ | ◐ | ○ | ○ |
| Zhou et al. (Zhou et al., 2025a) | May 2025 | Vulnerability detection and repair | 2018-2024 | 58 — 40 | ✓ | ● | ○ | ● | ○ | ○ |
| Germano et al. (Germano et al., 2025) | Nov 2025 | Code vulnerability detection, repair, expl. | 2018-2025 | 208 — 190 | ✓ | ● | ○ | ○ | ◐* | ○ |
| Sheng et al. (Sheng et al., 2025) | Sep 2025 | Software security, vulnerability detection | 2017-2025 | 58 — 58 | ✗ | ● | ◐ | ● | ◐ | ○ |
| *List only | ||||||||||
| This SLR | Dec 2025 | Software vulnerability detection | 2020-2025 | 263 | ✓ | ● | ● | ● | ● | ● |
Shiri Harzevili et al. (Shiri Harzevili et al., 2024), Shereen et al. (Shereen et al., 2024), and Shimmi et al. (Shimmi et al., 2025) focus on the application of machine learning and deep learning techniques for vulnerability detection, discussing seven, ten, and thirteen studies applying LLMs, respectively. These studies discuss machine learning and deep learning specific input (tree, graph, token) (Shiri Harzevili et al., 2024; Shimmi et al., 2025) or program representations (source, binary, intermediate) (Shereen et al., 2024) and provide an overview of common datasets. Shimmi et al. (Shimmi et al., 2025) also provide a taxonomy of methods, feature representations, and embedding techniques. In contrast to these surveys, we focus explicitly on LLM-based studies, discussing and categorizing LLM-specific input representations and adaptation techniques, as well as challenges with existing datasets and dataset usage. Taghavi Far and Feyz (Taghavi Far and Feyzi, 2025) consider a broader scope of software vulnerability detection, including, e.g., phishing detection, threat detection, and patch recommendation. The authors compare the application of LLMs to traditional program analysis techniques, discuss various fine-tuning techniques, and present selected pre-training and fine-tuning datasets. The surveys by Basic and Giaretta (Basic and Giaretta, 2025), Zhou et al. (Zhou et al., 2025a), and Germano et al. (Germano et al., 2025) are closely aligned with software vulnerability detection, extending to code security and vulnerability repair. Specifically, Basic and Giaretta (Basic and Giaretta, 2025) focus on vulnerabilities introduced by LLM-generated code as well as LLMs for detecting and fixing vulnerabilities (in a unified process), discussing various prompting strategies. Zhou et al. (Zhou et al., 2025a) discuss LLMs for vulnerability detection and repair, focusing on early fine-tuning, prompt engineering (zero-shot, few-shot), and retrieval augmentation techniques. Germano et al. (Germano et al., 2025) survey studies for vulnerability detection, repair, and explanation with LLMs, discussing the different models and programming languages applied as well as the distribution of datasets. Most closely to this SLR is the survey by Sheng et al. (Sheng et al., 2025), covering vulnerability detection techniques and insights. The authors discuss code pre-processing techniques, cover prompt engineering as well as fine-tuning techniques, and provide an overview of languages and granularities in benchmarks and datasets. In contrast, this SLR further offers a fine-grained taxonomy of prompting and fine-tuning techniques, incorporating learning paradigms and orchestration strategies. Furthermore, we introduce a temporal analysis of dataset usage and critically evaluate the vulnerability diversity and coverage of common benchmarks. Finally, we deepen the discussion on input representations and the role of auxiliary information in both prompting and tuning contexts, as well as classification and generation tasks.
Overall, this SLR extends the field through a dedicated and in-depth overview of 263 studies on LLM-based software vulnerability detection. The further this research field evolves, the more patterns and divergences can be identified. By systematically classifying studies using a comprehensive taxonomy, this SLR reveals focused insights that remain hidden in broader surveys, including how vulnerabilities are represented, which (new) techniques have been adapted and are effective for detection, and how study design influences applicability and comparability.
2.2. Vulnerability Datasets
A further distinguishing aspect of this SLR from previous surveys is its in-depth analysis of vulnerability detection datasets. High-quality, realistic, and diverse datasets play a crucial role in learning-based software vulnerability detection, prompting studies to analyze, classify, and evaluate their characteristics, quality, and impact on model performance. Discussed surveys have touched upon vulnerability detection datasets: Xu et al. (Xu et al., 2025) and Hou et al. (Hou et al., 2024) categorize dataset collection, i.e., open-source, collected, constructed, or industrial, as well as dataset type, i.e., code-based, text-based, graph-based, software repository-based, or combined. Shiri Harzevili et al. (Shiri Harzevili et al., 2024) categorize dataset sources into benchmark, hybrid, open source software, and repository. Further, they differentiate code, text, numerical, and hybrid dataset types. Shereen et al. (Shereen et al., 2024) compare frequently used datasets with respect to properties such as language, size, annotation, and granularity. Similarly, other discussed surveys (Ferrag et al., 2025a; Fu et al., 2025b; Shimmi et al., 2025; Taghavi Far and Feyzi, 2025; Germano et al., 2025; Sheng et al., 2025) list used datasets with respect to, e.g., size, programming language, granularity, or source. In contrast to previous surveys that primarily list datasets, we present a taxonomy and comprehensive analysis. We categorize datasets along type, granularity, source, and labeling methodology, and investigate additional quality dimensions, such as class balance, CWE diversity, and distribution, offering an in-depth analysis of represented vulnerabilities. Further, we add a time component and discuss emerging trends and use cases for datasets in LLM-based software vulnerability detection.
Other related studies focus explicitly on vulnerability detection datasets: Hu et al. (Hu et al., 2025) analyze benchmarks for evaluating LLM-based solutions, focusing on construction methods, programming languages, and metrics. Lin et al. (Lin et al., 2022b) examine dataset construction methodologies, comparing datasets across six dimensions, such as granularity, vulnerability type, and labeling approach. Guo et al. (Guo et al., 2024a) analyze selected datasets in terms of vulnerability distributions and types. Jain et al. (Jain et al., 2023) provide a code-centric evaluation of commonly used C/C++datasets for deep learning-based vulnerability detection. Moreover, the works (Guo and Bettaieb, 2023; Croft et al., 2023; Risse et al., 2025; Yadav and Wilson, 2024) provide in-depth analyses of quality issues in selected vulnerability datasets, such as data imbalance, low vulnerability coverage, biased vulnerability distribution, high duplication rates, and label noise. Despite this growing body of work, a targeted and comprehensive overview of vulnerability detection datasets commonly used for LLM-based software vulnerability detection approaches, along with their limitations and suitability, remains missing. This survey addresses this gap, offering a taxonomy and systematic analysis of datasets, their CWE coverage, diversity, and usage, thereby facilitating better comparability and benchmarking in future research.
3. Methodology
We followed established guidelines by Kitchenham et al. (Kitchenham and Charters, 2007; Petersen et al., 2015) for conducting SLRs in software engineering, i.e., defining research questions, conducting the search, defining selection criteria, and analyzing the selected studies. In the following, we outline the research focus of this SLR as well as the three phases of literature search, study selection, and study analysis, as visualized in Figure 1.
3.1. Research Focus
In this SLR, we address the following research questions RQ1–RQ5:
RQ1
How is the vulnerability detection task formulated and approached in LLM-based systems? Traditional deep-learning-based approaches predominantly frame the detection task as binary classification (vulnerable vs. non-vulnerable) (Shereen et al., 2024). While this formulation simplifies evaluation, it falls short in reflecting real-world software engineering workflows, where developers additionally require information on, e.g., vulnerability types and root causes. As LLMs are capable of producing structured reasoning steps and conceptual explanations, understanding whether LLM-based approaches move beyond this limited formulation is essential. We investigate how existing LLM-based approaches formulate the task, i.e., do they maintain the binary framing or do they extend or reformulate it, e.g., as multi-label vulnerability type prediction or vulnerability report generation (cf. Section 4.1).
RQ2
How is input, particularly the semantics of vulnerabilities, represented and encoded for LLMs? Existing machine learning and deep learning approaches convert code into suitable representations to capture semantics, for example, by constructing graph structures for GNNs or vectorizing code embeddings for CNNs (Shiri Harzevili et al., 2024; Shereen et al., 2024). For LLMs with their world knowledge, it remains unclear whether code alone is sufficient to capture specific vulnerability semantics. Investigating how studies represent input and whether they incorporate auxiliary information is crucial to understanding what information LLMs can leverage for effective detection (cf. Section 4.2).
RQ3
What are the predominant techniques for adapting LLMs to the task of software vulnerability detection? LLMs can be adapted via prompting, fine-tuning, or pre-training. Each technique has different implications for data requirements, computational costs, and practical deployment. Since these constraints affect the feasibility of integrating LLM-based methods into software engineering pipelines, this question investigates which adaptation techniques are used and how well they align with practical constraints (cf. Section 4.4.2).
RQ4
What datasets are used to train, tune, or evaluate LLM-based vulnerability detection approaches, and how does dataset selection impact applicability? Dataset selection influences the types and variety of vulnerabilities under detection. In terms of applicability, synthetic or oversimplified datasets lead to models with poor generalization. Therefore, understanding the characteristics (such as covered vulnerability types, size, and label quality) of the datasets used in current research is essential to enable reliable benchmarking and assess model generalization to real-world applications (cf. Section 5).
RQ5
How does dataset selection impact comparability across studies? Dataset selection influences whether results are comparable across studies. Inconsistent evaluation setups as well as not adhering to open science practices make it impossible to compare study results. Identifying commonly used datasets as part of evaluation setups enables researchers to select datasets to improve comparability across studies (cf. Section 5.3).
A joint investigation of RQ1-RQ5 is crucial to obtain a holistic understanding of LLM-based vulnerability detection research, from input representations to system design and evaluation, and to identify key levers to enhance robustness, practicability, and comparability.
3.2. Literature Search
To identify relevant studies, we conducted a systematic literature search across three major databases: IEEE Xplore, ACM Digital Library, and arXiv. We selected IEEE Xplore and the ACM Digital Library, as they host peer-reviewed publications from conferences and journals. We further include arXiv for preprints to capture ongoing research that is still in the submission process, given the rapidly evolving nature of this field.
The search string construction followed a two-step process.
First, we manually identified a set of relevant primary studies in our prior publication (Kaniewski et al., 2024).
We used these studies as a starting point to derive vulnerability detection-related terms. Second, to ensure systematic coverage of LLM-related terminology, we incorporated keywords used in related SLRs (Hou et al., 2024; Zhang et al., 2024a).
Specifically, we considered the keywords LLM, Large Language Model, Pre-trained, Pre-training, BERT, Codex, GPT, T5, ChatGPT, and GPT-*. We iteratively refined the search string by adding observed major model families and used wildcard extensions to ensure systematic coverage. The final search string is as follows:
Search String: (LLM OR LLMs OR Large Language Model OR Large Language Models OR pre-train* OR GPT* OR ChatGPT OR T5* OR CodeT5* OR Llama* OR CodeLlama* OR Codex OR BERT OR CodeBERT OR GraphCodeBERT OR Bard OR Claude OR Gemma* OR DeepSeek* OR Gemini* OR Mistral* OR Mixtral* OR Phi* or Qwen* OR CodeQwen* OR WizardCoder*) AND ((vulnerability OR vulnerabilities OR CVE OR CWE) AND (detect* OR identif* OR classif* OR analy* OR discover* OR assess*) AND (software OR program OR code))
3.3. Study Selection
Figure 1 illustrates the study selection pipeline used in this SLR. Using the final search string, we ran an automated search across the selected databases to identify studies whose titles or abstracts matched the specified keywords. We carried out the initial search in December 2024. After the initial crawl, we continuously integrated newly published material through database-specific search alerts and scholar.inbox (Flicke et al., 2025), ensuring that further relevant studies were considered and included.
Following this automated search, we applied a selection process to filter out irrelevant studies based on predefined inclusion (✓) and exclusion (✗) criteria, cf. Table 2. We retrieved each accessible study written in English and first filtered by title and abstract before performing a full text screen to assess if it is in scope. Specifically, we focus on studies that apply LLMs to actively detect software vulnerabilities, rather than applying LLMs to work with outputs of other detection tools, e.g., SAST tools. Further, we specifically focus on software vulnerability detection rather than detection in binary code, smart contracts, network logs, or detection of malware. We further exclude any duplicates or minor variations of multi- or extended-version papers as well as studies classified as reviews, surveys, theses, abstracts, posters, demonstrations, or tutorials.
Given the inclusion of preprints from arXiv, we conducted a manual quality assessment to determine study eligibility. We assessed whether the studies provided a clear contribution, adequately described the proposed workflow and implementation, outlined the evaluation setup (including datasets, baselines, and metrics), and presented coherent findings that support their central claims.
After filtering, we performed backward and forward snowballing. We manually reviewed the reference lists of the selected studies (backward snowballing) and used citation tracking via Google Scholar to identify additional papers citing the included studies (forward snowballing). Studies retrieved through snowballing also pass through the pipeline of inclusion and exclusion criteria.
Finally, at the time of writing this SLR, we cover 263 studies. Further studies added through the continuous search will be made available as part of the maintained repository, cf. (Kaniewski et al., 2025).
todo.
| Inclusion Criteria |
| ✓ Studies with title or abstract matching the specified search string. |
| ✓ Studies that use and evaluate LLMs for software vulnerability detection, introducing novel techniques or systems. |
| Exclusion Criteria |
| ✗ Studies not written in English. |
| ✗ Studies not accessible in full text. |
| ✗ Studies classified as reviews, surveys, theses, abstracts, posters, demonstrations, or tutorials. |
| ✗ Studies that focus on domains outside source code vulnerability detection, such as binary code, smart contracts, network analysis, or malware detection. |
| ✗ Studies that do not use LLMs to actively detect vulnerabilities but solely to work on outputs produced by other tools. |
| ✗ Studies that mention LLMs only in passing or use them merely as baselines. |
| ✗ Studies identified as duplicates or minor variations of multi- or extended-version papers from the same authors. |
3.4. Study Analysis
Following the selection process, we identified 263 papers that met the selection criteria. During the full-text review of each selected study, we extracted key data points, recording how the vulnerability detection task was defined, which LLMs were used, and what adaptation techniques were employed. Additionally, we extracted information on input representation formats, datasets used for fine-tuning and evaluation, targeted vulnerability types (e.g., specific CWEs), and evaluation metrics. This structured analysis enabled us to derive a comprehensive taxonomy of LLM-based software vulnerability detection approaches, address the research questions, and develop a deeper understanding of current trends, open challenges, and emerging directions in LLM-based software vulnerability detection.
The earliest included study dates back to 2020 (cf. Fig 2 (a)). Since then, we can observe a rapid increase in LLM-based vulnerability detection studies. Notably, 78 (29.7%) of the selected studies were sourced from arXiv, cf. Fig 2 (b), reflecting both the fast-paced development of this field and the prevalence of recent works still undergoing peer review. For the peer-reviewed studies, Fig 2 (c) visualizes the distribution of the selected studies across venue tiers. In addition, Table 3 lists the respective A*/A conferences as well as Q1 journals with their study counts.
| Type | Venue (Count) |
| Conf. A* | ICSE (13), EMNLP☼ (4), USENIX Security† (3), ASE (3), ESEC/FSE (2), INFOCOM◦ (2), ACL◦ (2), NDSS† (1), SP† (1), ICLR☼ (1) |
| Conf. A | ISSRE (3), ESORICS† (3), ACSAC† (2), ISSTA (2), SANER (2), MSR (2), DSN◦ (1), EASE (1), EuroS&P† (1), ASIACCS† (1), ICPC (1), ICSME (1), ICST (1), NAACL◦ (1), RAID† (1) |
| Jour. Q1 | TSE (7), IEEE Access◦ (4), COSE† (4), EMSE (4), TOSEM (3), JISA† (3), JSS (3), IST (1), Info. Fusion◦ (1), Cluster Computing◦ (1), TDSC† (1), Neural Comput. Appl.◦ (1) |
4. Vulnerability Detection with LLMs
Recent research has adopted a variety of techniques for detecting software vulnerabilities using LLMs. In this chapter, we present a comprehensive taxonomy that categorizes the different ways LLMs are applied to vulnerability detection. Specifically, we examine task formulations, input representations, system architectures, and techniques. The overall taxonomy is illustrated in Figure 3. We discuss each category on the basis of the surveyed studies in the following subsections.
4.1. Task Formulation (F)
Software vulnerability detection can be framed in multiple ways, depending on how the task is formulated and what objectives are prioritized beyond detection (e.g., explanation). With respect to the objectives, the integration of LLMs into vulnerability detection introduces unique advantages over traditional approaches. Unlike rule-based approaches, which require definitions of vulnerable patterns, the language understanding capabilities of LLMs allow for highly flexible goal setting; detection objectives can be refined or expanded, e.g., adding reasoning or vulnerability localization, simply by adjusting natural language prompts. LLMs further have extensive pre-trained world knowledge, such as software architecture patterns, programming language-specific constructs, and broad cybersecurity concepts. Such domain knowledge is difficult to encode explicitly in traditional tools or classical ML approaches, but serves as a versatile starting point for LLM-based detection. In the following, we outline and discuss the most common detection task formulations used with LLMs in the surveyed studies: classification and generation tasks.
4.1.1. Classification (F1)
Vulnerability detection is most commonly formulated as a classification problem, differentiating binary and multi-class forms. Binary Classification (F1.1) determines whether a given code contains a security vulnerability or not (’Yes’/’No’). Vulnerability-Specific Classification (F1.1.1) (Shereen et al., 2024), as a refinement of the binary formulation, determines whether a given code contains a specific vulnerability, represented, e.g., by CWE-ID. Multi-Class Classification (F1.2) determines which specific type of vulnerability the given code contains (type prediction), often using CWE-IDs as class labels. This formulation may use a given pre-defined list of vulnerability types in a prompt engineering setting. Similarly, Multi-Label Classification (F1.3) aims to determine any vulnerability types present in the code, recognizing that a code snippet may contain multiple distinct vulnerabilities simultaneously.
Binary classification (F1.1, 195 studies) simplifies the detection task as it is reduced to a decision of whether a code sample is vulnerable or not. While this approach facilitates model training and deployment, it lacks fine-grained insight into the specific type or cause of the vulnerability, limiting its use in real-world vulnerability remediation.
Vulnerability-specific classification (F1.1.1, 20 studies) is more targeted, which is valuable in contexts with high-impact CWEs or regulatory compliance. However, it depends on curated datasets and may struggle to generalize to related vulnerability types not present in the training data.
Multi-class (F1.2, 61 studies) and multi-label (F1.3, 23 studies) classification provide the most detailed perspective by distinguishing among various types of vulnerabilities. These formulations align more closely with real-world vulnerability remediation, where identifying the exact nature of a vulnerability is essential for vulnerability patching. However, these formulations introduce additional complexity in both modeling and evaluation. Their effectiveness also depends heavily on high-quality, CWE-labeled, and well-distributed datasets.
Some studies combine or cascade these formulations. For example, a binary classifier may act as a preliminary filter or router, forwarding likely vulnerable samples to a CWE-specific classifier for further analysis (Yang et al., 2025a).
4.1.2. Generation (F2)
While classification is the primary formulation, it offers limited explanations and insights into the root causes of vulnerabilities. Therefore, several studies extend the scope of classification and leverage the language understanding and generation capabilities of LLMs to generate additional outputs. This generative approach takes advantage of the model’s flexibility, allowing researchers to request diverse detection outputs, such as reasoning chains or fix suggestions, through simple natural language instructions. We categorize generative tasks as (i) vulnerability description, (ii) root cause reasoning, and (iii) structured vulnerability report.
Vulnerability Description (F2.1) focuses on generating a high-level natural language summary or description of the specific detected vulnerability, providing developers with its context (Chan et al., 2023; Cipollone et al., 2025; Ezenwoye et al., 2024; Fu et al., 2023a; Gajjar et al., 2025a; Ingemann Tuffveson Jensen et al., 2024; Islam et al., 2023; Kouliaridis et al., 2024; Lin and Mohaisen, 2024; Li et al., 2025d, c; Mohajer et al., 2024; Safdar et al., 2025a; Shu et al., 2025; Sovrano et al., 2025; Steenhoek et al., 2025b; Tamberg and Bahsi, 2025; Wang et al., 2025g, d; Yıldırım et al., 2024; Yildiz et al., 2025).
Root Cause Reasoning (F2.2) moves beyond high-level descriptions to provide a code-level justification or reasoning of the specific vulnerability and its root cause (Ahmed et al., 2025; Dolcetti et al., 2025; Du et al., 2025a; Espinha Gasiba et al., 2024; Hairan and Şahman, 2025; Islam et al., 2024b; Jararweh et al., 2025; Keltek et al., 2024; Khare et al., 2025; Kholoosi et al., 2024; Li et al., 2025g, i, f, h; Liu et al., 2023, 2024b; Li et al., 2025c; Mao et al., 2025; Mathews et al., 2024; Mhatre et al., 2025; Ni et al., 2023b; Safdar et al., 2025a; Sayagh and Ghafari, 2025; Simoni et al., 2025; Steenhoek et al., 2025a; Sun et al., 2025b; Ullah et al., 2024; Wang et al., 2025a, b, e; Wei et al., 2025; Wen et al., 2025; Weyssow et al., 2025; Widyasari et al., 2024; Yang et al., 2025b, c; Zhu et al., 2025a; Zibaeirad and Vieira, 2025; Chen et al., 2025a; Nguyen et al., 2025a). However, the reliability of LLM-generated reasoning remains under scrutiny. For instance, Steenhoek et al. (Steenhoek et al., 2025a) perform an error analysis on LLM responses, demonstrating that the models struggle to reason about the code semantics relevant to identifying vulnerabilities, especially subtle semantic differences caused by small textual changes. Wen et al. (Wen et al., 2025) propose forward and backward reasoning, where forward reasoning aims to deduce the causes of vulnerabilities, while backward reasoning seeks to understand the code changes in the fixes. Their results show that LLMs can reason about vulnerabilities but often lack the domain-specific knowledge required to effectively invoke such reasoning.
Other approaches request a Structured Vulnerability Report (F2.3) (Dubniczky et al., 2025; Gajjar et al., 2025b; Gnieciak and Szandala, 2025; Keltek et al., 2024; Khare et al., 2025; Mahmud et al., 2025; Noever, 2023; Pelofske et al., 2024; Peng et al., 2025; Shashwat et al., 2024; Sun et al., 2025b; Wei et al., 2025; Weyssow et al., 2025; Widyasari et al., 2025; Xia et al., 2025; Yang et al., 2025c; Yu et al., 2025; Zhou et al., 2025b; Zhu et al., 2025a; Zibaeirad and Vieira, 2025; Jiao et al., 2025; Tsai et al., 2025b; Nguyen et al., 2025a): Studies often aggregate multiple predictive sub-tasks, such as vulnerability localization, severity estimation, type justification, and repair fixes, into a single structured output (e.g., JSON, structured text, or standardized static analysis formats (Gnieciak and Szandala, 2025)). For example, Zibaeirad and Vieira (Zibaeirad and Vieira, 2025) prompt models to generate a report containing the CWE-ID, fix suggestion, vulnerability severity, and exploitability. Similarly, Yang et al. (Yang et al., 2025c) structure the output to include the label, confidence score, vulnerability type, and affected components. Zhou et al. (Zhou et al., 2025b) produce reports detailing function names, namespaces, source/sink type, and dangerous parameter indices.
These generative tasks reflect a broader shift in the role of LLMs from simple classifiers to end-to-end vulnerability assistants that support developers interactively throughout the software development life cycle. Figure 4 maps the relationships between classification and generation tasks across the surveyed studies. Each node is annotated with the absolute number of studies using that task formulation, providing a quantitative overview of the landscape. The flows further visualize studies that combine classification with generative tasks. These combined tasks create opportunities for multitask benchmarks that evaluate detection accuracy alongside, e.g., reasoning quality, localization precision, and severity estimation, offering a more holistic perspective on vulnerability remediation.
4.2. Input Representation (I)
The representation of the input code and additional context plays a pivotal role in effective vulnerability detection. Distinct from traditional static analysis tools that require language-specific rules, LLMs offer unique capabilities in how they process input. First, their extensive pre-trained world knowledge allows them to interpret diverse input forms, e.g., code and natural language instructions. This flexibility enables the processing of multi-language source code with varied vulnerability types and auxiliary information within a unified framework. Second, their semantic understanding allows LLMs to generalize; by reasoning about the input rather than matching fixed vulnerable patterns, LLMs may identify rare or novel vulnerabilities. Considering these capabilities, we categorize input representations along the code and auxiliary information axes.
4.2.1. Code (I1)
Raw (I1.1) input representations of code involve feeding source code directly into the model during inference as plain text sequences (used by 139 studies). This approach relies entirely on the model’s understanding and pre-learned knowledge of programming syntax to detect vulnerable patterns. This representation is widely used for training or fine-tuning LLMs. However, raw code inputs may struggle to capture the deeper semantics and context necessary for detecting subtle or complex vulnerabilities. As such, this approach may underperform in scenarios requiring precise reasoning about program logic, data flow, or code structure.
Combining raw code with or expanding representations to Structured (I1.2) code enhances the model’s understanding of code by capturing deeper syntactic and semantic structures of code and vulnerabilities, thereby addressing the limitations of relying solely on textual patterns. Code is transformed into graph-based or other structured formats that expose control flow, data flow, and dependency relationships relevant to vulnerability semantics. Common graph representations include Abstract Syntax Trees (ASTs) (Ahmed et al., 2025; Bahaa et al., 2024; JianJie and Le, 2023; Li et al., 2025j; Wen et al., 2024a; Zhang et al., 2025a; Yu and Zhang, 2025), specifically flattened (Li et al., 2025j) or decomposed (Zhang et al., 2025a), Control Flow Graphs (CFGs) (Gajjar et al., 2025a; Liu et al., 2025a; Lu et al., 2024; Sun et al., 2024; Zhang et al., 2023a), Data Flow Graphs (DFGs) (Jiang et al., 2024b; Liu et al., 2024f; Wang et al., 2024b; Yu et al., 2022), forms of Program Dependency Graphs (PDGs) (Du et al., 2024; Hin et al., 2022; Lu et al., 2024; Sun et al., 2025c; Wu and Zhang, 2024), combining data and control dependencies, and Code Property Graphs (CPGs) (Chang et al., 2024; Feng et al., 2025; Lekssays et al., 2025; Li et al., 2025e; Liu et al., 2025b; Ngan et al., 2025; Ni et al., 2023a; Rangapuram and Ratha, 2025; Liang et al., 2024; Yu and Zhang, 2025), which integrate multiple code views (AST, CFG, PDG). Some studies employ composite or multi-layered graph structures that integrate several of the above (Jiang et al., 2022; Liu et al., 2024a; Islam et al., 2023, 2024a; Zheng et al., 2025, 2024; Liu et al., 2025a; Lu et al., 2024), which can outperform other graphs (Zhang et al., 2023b), or combine multiple modalities including text, graphs, and visual representations (Ni et al., 2023a).
In addition to graph structures, other structured inputs aim to isolate semantically relevant code regions, such as program or code slices (Wu and Xiao, 2025; Tian and Zhang, 2025; Ding et al., 2023; Huang et al., 2022; Sun et al., 2023; Zhu et al., 2022), chunks (Halder et al., 2025), or code gadgets (Pham et al., 2024; Gujar, 2024; G”̈urfidan, 2024; Kim et al., 2022; Li et al., 2021; Omar, 2023; Omar and Shiaeles, 2023; Peng et al., 2023; Purba et al., 2023; Quan et al., 2025; Thapa et al., 2022; Zhao and Guan, 2025; Zhao et al., 2023; Chen et al., 2025a). Code gadgets, in particular, consist of program statements that are semantically related to each other through data or control dependencies (Li et al., 2018), isolating vulnerability-relevant code regions. Other studies include program traces, i.e., execution-aware code structure (Ding et al., 2024b; Mächtle et al., 2025). Mächtle et al. (Mächtle et al., 2025) specifically build trace gadgets, i.e., trace the code before slicing. By making structural and semantic dependencies explicit, structure-aware representations enable models to reason more effectively about vulnerability-related behaviors, especially those that span across the codebase or require understanding of program logic and execution flow.
4.2.2. Auxiliary Information (I2)
To enhance the effectiveness of vulnerability detection, studies move beyond source code by augmenting the input with auxiliary information, particularly in prompt-engineering settings. This additional information provides the model with domain knowledge or code context. We categorize this auxiliary information into (i) vulnerability information, (ii) semantic artifacts, (iii) execution artifacts, and (iv) tool output.
Vulnerability Information (I2.1) involves explicitly equipping the model with domain vulnerability knowledge. Most commonly, studies include a pre-defined list of CWEs to the prompt to narrow the classification scope (Bakhshandeh et al., 2023; Cao and Jun, 2024; Dozono et al., 2024; Fu et al., 2023b; Gnieciak and Szandala, 2025; Li et al., 2025i; Liu et al., 2024b; Sayagh and Ghafari, 2025; Shashwat et al., 2024; Steenhoek et al., 2025a; Wang et al., 2023a; Xia et al., 2025; Yu et al., 2025) or provide labeled examples (few-shot settings) (Chan et al., 2023; Ding et al., 2025; Dolcetti et al., 2025; Farr et al., 2025; Hannan et al., 2025; Li et al., 2025j; Liu et al., 2024f, 2023; Mao et al., 2025; Ni et al., 2025; Nong et al., 2024; Ouchebara and Dupont, 2025; Safdar et al., 2025a; Shu et al., 2025; Sovrano et al., 2025; Trad and Chehab, 2025; Tsai et al., 2025a; Wang et al., 2025b; Widyasari et al., 2024; Yang et al., 2025c; Yildiz et al., 2025; Yin et al., 2024; Zhou et al., 2024a, b; Chen et al., 2025a; Jiao et al., 2025; Tsai et al., 2025b; Le et al., 2024; Lu et al., 2024; Sheng et al., 2024; Zhang et al., 2025b). Other studies include more detailed vulnerability (CWE) definitions and descriptions (Sun et al., 2025b; Shimmi et al., 2024b; Li et al., 2025g; Kouliaridis et al., 2024; Ullah et al., 2024; Zhu et al., 2025a; Mohajer et al., 2024; Widyasari et al., 2025; Ceka et al., 2025; Pelofske et al., 2024; Mathews et al., 2024; Du et al., 2025a), particularly as part of retrieval architectures, or full analysis reports (Luo et al., 2025).
Semantic Artifacts (I2.2) focus on the high-level intent and functionality of the code, bridging the gap between syntax and understanding. For example, studies provide natural language explanations of the code functionality (Gajjar et al., 2025b; Torkamani et al., 2025), questions regarding the code’s behavior (Jararweh et al., 2025), explicit descriptions of data sources and sinks (Li et al., 2025e), or code comments and docstrings (Huynh et al., 2025).
Execution Artifacts (I2.3) capture the behavior, structural dependencies, or runtime environment of the code. Studies enrich inference input with data and control dependencies, interprocedural interactions, and environment constraints (Ahmed et al., 2025). For instance, researchers include information on callee functions (Sun et al., 2025b; Li et al., 2025i), dynamic execution signals such as unit test results (Mhatre et al., 2025), or code-context (Li et al., 2025f; Sayagh and Ghafari, 2025; Li et al., 2025i), e.g., global variables, type declarations, and library imports. Data flow analysis techniques are also used (Li et al., 2025j; Tamberg and Bahsi, 2025; Yang et al., 2025b; Zhou et al., 2025b), e.g., to extract dependencies from call graphs (Wen et al., 2024b).
Tool Output (I2.4) augments the model context with the outputs of other tools, most commonly SAST tools, allowing the LLM to interpret the static analysis findings to refine further predictions (Bakhshandeh et al., 2023; Keltek et al., 2024; Munson et al., 2025; Wang et al., 2025e; Yang et al., 2025c). Yang et al. (Yang et al., 2025c) inputs the results of other deep learning models (i.e., the prediction probability) to inform the final classification.
4.3. System Architecture (S)
Vulnerability detection systems using LLMs differ significantly in which and how models are integrated into the overall system. Broadly, we categorize existing approaches into two groups: LLM-centric systems, where the LLM serves as the main predictive component, and hybrid systems, which integrate an LLM with other learning-based components.
4.3.1. LLM (S1)
In LLM-centric approaches, the LLM performs the core predictive tasks, such as classification and reasoning. The model is typically adapted to the detection task through prompt engineering or fine-tuning. Additional tools, such as SAST, only provide auxiliary information rather than predictive signals. The surveyed studies employ a variety of LLMs with different architectures, pre-training domains, and scales. In the following, we discuss LLMs within these categories to categorize systems in more detail.
Model Architecture (S1.1)
Most LLMs are based on the Transformer architecture, which includes encoder and decoder modules with a self-attention mechanism. We follow other works and categorize the LLMs used in the surveyed studies into the three architectural groups (i) encoder-only, (ii) encoder-decoder, and (iii) decoder-only, as introduced by Pan et al. (Pan et al., 2024).
Encoder-Only (S1.1.1) LLMs, e.g., models from the BERT (Devlin et al., 2019) family (CodeBERT (Feng et al., 2020), GraphCodeBERT (Guo et al., 2021), RoBERTa (Liu et al., 2019), DistilBERT (Sanh et al., 2020)), use only the encoder to encode the sentence and comprehend the relationships between words (Pan et al., 2024). The encoder maps the input code into a fixed-size dense vector representation (embedding) that captures semantic relationships. In vulnerability detection, these models are typically adapted via fine-tuning by adding a classification head (e.g., a multi-layer perceptron) to output a probability score or class label. Alternatively, they function as frozen feature extractors, generating embeddings that serve as input for external classifiers.
Encoder-Decoder (S1.1.2) LLMs, e.g., T5 (Raffel et al., 2020), CodeT5 (Wang et al., 2021, 2023c), or UniXcoder (Guo et al., 2022), use both the encoder and decoder module in a sequence-to-sequence framework (Pan et al., 2024). The encoder maps the input code into a latent hidden-state representation, which the decoder uses to generate a target output sequence. For vulnerability detection, this architecture allows the task to be framed as text generation, where the input is the source code and the output is a generated text string, e.g., the label ”vulnerable” or a natural language explanation.
Decoder-Only (S1.1.3) LLMs use only the decoder module, using auto-regressive mechanisms to predict the next token in a sequence. The input, typically a prompt containing instructions and the code snippet, is processed sequentially, and the output is generated as a continuation of this input, enabling the generation of reasoning chains and structured vulnerability reports. Many state-of-the-art LLMs follow this architecture, e.g., the GPT family (Radford et al., 2019; Brown et al., 2020; OpenAI et al., 2024) and the Llama family (Touvron et al., 2023a, b; Grattafiori et al., 2024). Decoder-only LLMs are the dominant architecture used for vulnerability detection in the surveyed studies.
Pre-Training Domain (S1.2)
Surveyed studies rely on LLMs from two main domains: general-purpose LLMs and code LLMs. General-Purpose LLMs (S1.2.1) are pre-trained on natural language corpora and tend to perform well on explanation-based or natural language-heavy tasks. Examples from this domain are the BERT family, the GPT family, the Llama family, the Qwen series (Bai et al., 2023; Yang et al., 2024b; Qwen et al., 2025), Gemma (Team et al., 2024b, c), DeepSeek (DeepSeek-AI et al., 2025, 2024a), Mistral (Jiang et al., 2023), Mixtral (Jiang et al., 2024a), Claude (Anthropic, 2024), Gemini (Team et al., 2024a), and Phi (Javaheripi et al., 2023; Abdin et al., 2024). Code LLMs (S1.2.2) are pre-trained specifically on large-scale code corpora for code-related tasks, enhancing performance on syntax-sensitive or program-structure-aware tasks. Examples include CodeBERT (Feng et al., 2020), GraphCodeBERT (Guo et al., 2021), StarCoder (Li et al., 2023; Lozhkov et al., 2024), UniXcoder (Guo et al., 2022), the GPT-based Codex (Chen et al., 2021), CodeLlama (Rozière et al., 2024), CodeQwen 1.5 (Bai et al., 2023), Qwen2.5-Coder (Hui et al., 2024), CodeGemma (Team et al., 2024d), DeepSeek-Coder (Guo et al., 2024c; DeepSeek-AI et al., 2024b), and CodeT5 (Wang et al., 2021, 2023c).
Model Scale (S1.3)
The parameter count of an LLM typically serves as a proxy for its capacity to learn complex patterns and encode domain knowledge. However, increasing scale imposes proportional demands on computational resources for training and inference. Notably, there exists no formal consensus regarding the exact parameter threshold that constitutes a ”large” language model. Despite this ambiguity, to analyze the trade-off between performance and resource efficiency, we categorize the models used in the surveyed studies into four distinct groups:
Tiny (S1.3.1): Models with fewer than 1 billion parameters (1B). This category primarily includes pre-trained language models such as CodeBERT or UniXcoder. Due to their lightweight nature, they are highly efficient for fine-tuning and inference, making them suitable for resource-constrained environments.
Small (S1.3.2): Models with 1 to 8 billion parameters (1B–8B), such as Gemma-2B, Llama2-7B, or Llama3.1-8B. This scale represents a trade-off for many researchers; these models provide a significant improvement in generative capabilities over tiny models while remaining small enough to be deployed or fine-tuned on high-end consumer-grade hardware.
Medium (S1.3.3): Models between 8 and 70 billion parameters (8B–70B), including CodeLlama-34B, Mixtral 8x7B, and DeepSeek-R1-32B. These models typically outperform smaller counterparts in complex reasoning tasks but exceed the memory capacity of standard consumer hardware, necessitating enterprise-grade GPUs or multi-GPU setups.
Large (S1.3.4): Models exceeding 70 billion parameters (70B), such as GPT-4, Llama3.1-405B, and DeepSeek-v3. These models achieve state-of-the-art performance on vulnerability detection benchmarks. However, due to their immense computational requirements, they are predominantly accessed via APIs or require extensive high-performance computing clusters for local deployment.
Figure 5 visualizes the relationship between model scale and adaptation techniques, e.g., fine-tuning, parameter-efficient fine-tuning, and prompt engineering; highlighting that tiny and small models are often fine-tuned in a full-parameter setting while large models are mainly prompted.
4.3.2. Hybrid (S2)
Hybrid architectures combine LLMs with other learning-based components to leverage complementary strengths. Often, the LLM is used to produce embeddings that are subsequently processed by neural networks such as Recurrent Neural Networks (RNNs), Convolutional Neural Networks (CNNs), or Graph Neural Networks (GNNs).
RNNs (S2.1) are neural architectures designed to handle sequential data by modeling temporal and contextual dependencies between elements, e.g., syntactic, semantic, and control-flow relationships (Allamanis et al., 2018). Prominent RNN variants include Long Short-Term Memory networks (LSTMs) and their bidirectional forms (BiLSTMs), which are designed to handle long-range dependencies effectively. In several studies, BERT embeddings are directly fed into (Bi-)LSTM networks (Pham et al., 2024; Ziems and Wu, 2021; Cheng et al., 2025; Li et al., 2021; Xiong and Dong, 2024).
CNNs (S2.2) are neural architectures specialized in identifying local spatial patterns or features within data through convolutional filters. Initially developed for image recognition tasks, CNNs have been successfully adapted to textual and sequential data analysis by treating sequences as one-dimensional spatial inputs. This capability allows CNNs to effectively capture localized syntactic and semantic patterns in code sequences, making them useful for tasks such as vulnerability detection (Lin et al., 2020). Studies apply CNNs directly to code or graph embeddings (Zhang et al., 2023a; Pham et al., 2024). Others use pre-trained embeddings within TextCNN architectures (Hanif and Maffeis, 2022), or integrate CNNs with attention mechanisms to enhance feature extraction (Wang et al., 2023b; Wu and Zhang, 2024).
GNNs (S2.3) are specialized neural architectures designed to operate directly on graph-structured data, effectively modeling relationships between nodes and their structural dependencies. Given structured representations of code, such as AST and CFG, several studies combine LLMs with GNN-based architectures to take advantage of both sequential and structural representations of code (Lin et al., 2020). A common approach is to pass code embeddings, e.g., generated by CodeBERT or GraphCodeBERT, into GNN-based classifiers (Yu and Zhang, 2025; Tian and Zhang, 2025; Kong et al., 2024; Liu et al., 2025b; Zheng et al., 2024; Hin et al., 2022; Rangapuram and Ratha, 2025; Sun et al., 2025c, 2023; Ngan et al., 2025). Other studies use LLMs to feed graph embeddings (Islam et al., 2023, 2024a; Jiang et al., 2024b; Quan et al., 2025) or multiple modalities of a function, including text, graph, and image representations (Ni et al., 2023a) into Graph Convolution Networks (GCNs). Yang et al. (Yang et al., 2024a) use a GNN as a light-weight adapter layer and concatenate its learned embedding with a fine-tuned LLM. Similarly, Gajjar et al. (Gajjar et al., 2025a) use GCN and GAT to extract structural features, subsequently fusing these graph embeddings with semantic embeddings derived from an LLM.
Several hybrid architectures combine multiple neural components, integrating RNNs, CNNs, and GNNs in sequence or parallel. For example, Chang et al. (Chang et al., 2024) process CodeBERT embeddings using a two-layer GCN and a BiLSTM. Bahaa et al. (Bahaa et al., 2024) use a multi-layer approach composed of a DistilBERT embedding layer followed by CNN and BiLSTM layers. Mei et al. (Mei et al., 2023) use CodeBERT embeddings in three hybrid models, i.e., CLSTM, CBiLSTM (sequential structure), and CNN-BiLSTM (parallel structure). JianJie and Le (JianJie and Le, 2023) present a multi-base classifier combining BERT-encoder, GCN, CNN, and BiLSTM; the outputs are fused via stacking ensemble learning. Sun et al. (Sun et al., 2023) propose a stacking ensemble strategy that integrates serialization-based and graph-based neural networks to exploit complementary code semantics.
Some studies (Jiang et al., 2025a; Peng et al., 2023; Xue et al., 2025; Melo et al., 2025; Pham et al., 2024) combine LLMs with Other (S2.4) learning-based components that do not rely on RNNs, CNNs, or GNNs. For example, Melo et al. (Melo et al., 2025) evaluate the effectiveness of Sparse Autoencoders (SAEs) when applied to representations from LLMs, which are then classified using traditional machine learning models such as random forests. Jiang et al. (Jiang et al., 2025a) pass embeddings to a Policy network for prediction. Peng et al. (Peng et al., 2023) use CodeBERT embeddings as input to a Transformer-based classifier. Similarly, Xue et al. (Xue et al., 2025) perform feature fusion from multiple pre-trained models and process the combined representation with a Transformer model.
4.4. Technique (T)
To use LLMs effectively for vulnerability detection, studies employ various techniques ranging from simple prompting to complex system and training designs. We broadly categorize techniques into three distinct strategies: (i) feature extraction, (ii) adaptation, and (iii) orchestration.
4.4.1. Feature Extraction
In the most simple approach, language models are used in their pre-trained form without any additional training to extract features by generating input embeddings (Feature Extraction (T1)) (Jiang et al., 2022; Yu and Zhang, 2025; Tian and Zhang, 2025; Pham et al., 2024; Wu and Zhang, 2024; Haque et al., 2025; Liang et al., 2024; Du et al., 2023; Gujar, 2024; Hin et al., 2022; Islam et al., 2023, 2024a; Jararweh et al., 2025; Jiang et al., 2024b, 2025a; Le et al., 2025; Mahmud et al., 2025; Melo et al., 2025; Ngan et al., 2025; Nguyen et al., 2025c; Ni et al., 2023a; Peng et al., 2023; Rangapuram and Ratha, 2025; Sultan et al., 2024; Sun et al., 2023, 2025c; Wang et al., 2025g, 2023b; Xue et al., 2025; Zhang et al., 2023a; Zhao et al., 2023; Zhu et al., 2025b). These high-dimensional vector representations capture the semantic and syntactic properties of the code and serve as input for downstream classifiers.
4.4.2. Adaptation Technique (T2)
Beyond using LLMs as static feature extractors, recent work increasingly focuses on adapting LLMs to the vulnerability detection task through various state-of-the-art techniques, which differ in how they modify the model’s parameters. The primary adaptation techniques are pre-training, fine-tuning, and prompt engineering. Each technique reflects different trade-offs between computational cost, model adaptation, and task-specific performance. In the following, we discuss these adaptation techniques as observed in the surveyed studies.
Prompt Engineering (T2.1)
Prompt engineering refers to the design and structuring of input prompts in a way that elicits desired behaviors or outputs without requiring any modifications to the model’s weights. There are typically two types of prompts: a system prompt (providing general context or behavioral instructions) and a user prompt (specifying the task and providing the code to be analyzed). As LLMs remain sensitive to the phrasing, context, and formatting of prompts, numerous prompting strategies have emerged to improve task alignment and performance.
In Zero-Shot Prompting (T2.1.1), the model is expected to rely solely on its pre-trained knowledge to perform vulnerability detection. The LLM is provided with the task description and input (i.e., a potentially vulnerable code snippet under detection) (Ambati et al., 2024; Bae et al., 2024; Bakhshandeh et al., 2023; Chan et al., 2023; Cheshkov et al., 2023; Cipollone et al., 2025; Dozono et al., 2024; Dubniczky et al., 2025; Eberhardt and Milánkovich, 2024; Espinha Gasiba et al., 2024; Ezenwoye et al., 2024; Farr et al., 2025; Gong et al., 2025; Guo et al., 2024b; Hairan and Şahman, 2025; Huynh et al., 2025; Ingemann Tuffveson Jensen et al., 2024; Khare et al., 2025; Kholoosi et al., 2024; Kouliaridis et al., 2024; Lin and Mohaisen, 2024; Li et al., 2025j; Liu et al., 2024f, b; Li et al., 2025c; Lu et al., 2024; Luo et al., 2025; Maligazhdarova et al., 2025; Ni et al., 2025; Noever, 2023; Ouchebara and Dupont, 2025; Ozturk et al., 2023; Purba et al., 2023; Ridley et al., 2023; Safdar et al., 2025a; Saimbhi and Akpinar, 2024; Shu et al., 2025; Simoni et al., 2025; Steenhoek et al., 2025b, a; Sun et al., 2025b; Trad and Chehab, 2025; Tsai et al., 2025a; Ullah et al., 2024; Wang et al., 2025a, d, b; Wei et al., 2025; Wu et al., 2023; Yıldırım et al., 2024; Yu et al., 2025; Zhang et al., 2025b; Zhou et al., 2024a, b; Zibaeirad and Vieira, 2025, 2024). Typical prompts include simple binary instructions such as ”Is the following code vulnerable? Yes or No,” targeted CWE-specific queries such as ”Does the following code have a CWE-xxx vulnerability?”, or openly phrased ”Does this code have any vulnerabilities?” Studies often experiment with arbitrary variations to the base prompt, such as modifying task phrasing (e.g., ”vulnerable”, ”buggy”, ”insecure”). In a role-based setting, studies include a role in the prompt (e.g., ”Act as a vulnerability detection system” (Liu et al., 2024f)) to guide model behavior. Due to its simplicity, zero-shot prompting is commonly used as a starting point and serves as a baseline for comparison with more advanced prompting techniques.
In-Context Learning (T2.1.2) refers to the prompt engineering technique of embedding additional contextual information (drawing upon the auxiliary information discussed, see Section 4.2) into the model’s input to improve detection without parameter updates. Researchers frequently provide vulnerability definitions, e.g., a candidate list of CWEs (Dozono et al., 2024; Bakhshandeh et al., 2023; Shashwat et al., 2024; Steenhoek et al., 2025a; Wang et al., 2023a; Xia et al., 2025; Yu et al., 2025; Fu et al., 2023b) or detailed vulnerability descriptions (Cao and Jun, 2024; Shimmi et al., 2024b; Gnieciak and Szandala, 2025; Mathews et al., 2024; Mohajer et al., 2024; Ullah et al., 2024). Adopting a code-centric perspective, studies also enrich prompts with execution context (Sayagh and Ghafari, 2025; Sun et al., 2025b; Tamberg and Bahsi, 2025; Zhang et al., 2025a; Ahmed et al., 2025; Mhatre et al., 2025; Li et al., 2025f) or descriptions of the code’s intended functionality (Liu et al., 2024f, 2023; Yang et al., 2025b; Torkamani et al., 2025), supporting the vulnerability detection task. Other approaches use a hybrid workflow by embedding prior detection results from static analysis or similar tools, asking the LLM to verify or refine these findings (Keltek et al., 2024; Munson et al., 2025; Wang et al., 2025e; Yang et al., 2025c).
Few-Shot Prompting (T2.1.3) (Brown et al., 2020), a specific in-context learning technique, provides the model with one or multiple labeled input-output examples relevant to the task. By presenting these examples in the prompt, the model is guided to infer the desired task behavior, enabling in-context learning without modifying model weights. The examples used for few-shot prompting can be static, i.e., manually selected and reused across prompts (Ding et al., 2025; Dolcetti et al., 2025; Le et al., 2024; Li et al., 2025g; Mao et al., 2025; Ni et al., 2025; Sovrano et al., 2025; Trad and Chehab, 2025; Wang et al., 2025b; Yildiz et al., 2025; Yin et al., 2024; Chen et al., 2025a), or vulnerability-specific, i.e., tailored examples targeting particular vulnerability types (Farr et al., 2025; Ouchebara and Dupont, 2025; Ullah et al., 2024). In another study, the examples are generated by the LLM itself (Chan et al., 2023). Other common strategies for example selection, as discussed by Steenhoek et al. (Steenhoek et al., 2025a), include using random vulnerable and non-vulnerable examples from the training dataset (Li et al., 2025j; Mohajer et al., 2024; Ni et al., 2025; Ouchebara and Dupont, 2025; Steenhoek et al., 2025a; Widyasari et al., 2024; Yang et al., 2025b; Zhou et al., 2024a, b), constructing contrastive pairs where both a vulnerable code snippet and its corresponding fixed version are provided within the same prompt (Ceka et al., 2025; Nong et al., 2024; Steenhoek et al., 2025a; Yang et al., 2025b; Yildiz et al., 2025), or selecting examples from the training dataset based on similarity to the input code (Hannan et al., 2025; Ni et al., 2025; Shu et al., 2025; Steenhoek et al., 2025a; Widyasari et al., 2024; Zhou et al., 2024b; Tsai et al., 2025b).
Retrieval-Augmented Generation (T2.1.4) (Lewis et al., 2020) extends the few-shot paradigm by dynamically augmenting the prompt with relevant external information retrieved from a private knowledge base. Retrieval-Augmented Generation (RAG) systems use a retriever component to identify and fetch contextually similar entries based on embedding similarity or other relevance metrics, e.g., BM25. The LLM then works with the provided knowledge to generate the output. A key advantage of RAG over few-shot prompting is its ability to dynamically incorporate domain-specific and up-to-date knowledge, making it especially useful for integrating evolving vulnerability knowledge. In the context of vulnerability detection, RAG-based approaches use a wide range of retrieval content. Some studies retrieve the most similar code snippets or code dependencies along with known vulnerability information (Liu et al., 2024f, 2023; Lu et al., 2024; Ouchebara and Dupont, 2025; Safdar et al., 2025a; Sheng et al., 2024; Steenhoek et al., 2025a; Tsai et al., 2025a; Wen et al., 2024b; Yang et al., 2025c; Zhang et al., 2025b; Zhou et al., 2025b; Zhu et al., 2025a; Jiao et al., 2025; Tsai et al., 2025b). Others rely on abstract code functionality and vulnerability descriptions to better capture the vulnerability semantics and context of the input code (Du et al., 2025a; Li et al., 2025g; Shimmi et al., 2024b; Sun et al., 2025b). Additionally, vulnerability reports and CWE descriptions are retrieved to guide the model’s reasoning using structured vulnerability information (Cao and Jun, 2024; Keltek et al., 2024; Kouliaridis et al., 2024; Mathews et al., 2024; Pelofske et al., 2024; Sayagh and Ghafari, 2025).
Chain-of-Thought (CoT) Prompting (T2.1.5) (Wei et al., 2022) aims at enhancing the reasoning capabilities of LLMs by encouraging intermediate reasoning steps before arriving at a final prediction. Reasoning behavior is typically triggered by simple cues such as ”Let’s think step by step” (Bae et al., 2024; Ding et al., 2025; Dolcetti et al., 2025; Huynh et al., 2025; Ni et al., 2025; Pelofske et al., 2024; Safdar et al., 2025a; Sheng et al., 2024; Steenhoek et al., 2025a; Ullah et al., 2024; Wang et al., 2025b; Wen et al., 2025; Zhou et al., 2024a; Zibaeirad and Vieira, 2025; Yildiz et al., 2025), requests for explanation (Li et al., 2025c; Yang et al., 2025b; Chen et al., 2025a; Khare et al., 2025), or specific step-by-step instructions (Ceka et al., 2025; Du et al., 2024; Li et al., 2025g, i, h; Liu et al., 2024f; Mao et al., 2025; Peng et al., 2025; Sheng et al., 2024; Simoni et al., 2025; Sun et al., 2025b; Ullah et al., 2024; Wang et al., 2025b; Widyasari et al., 2024; Yang et al., 2025c; Yu et al., 2025; Zhou et al., 2025b; Zhu et al., 2025a; Zibaeirad and Vieira, 2025). Automated CoT prompt construction has also been explored (Trad and Chehab, 2025). Several studies adapt CoT prompting with domain-specific techniques. For instance, Nong et al. (Nong et al., 2024) propose a vulnerability-semantics-guided prompt format to guide the model through relevant data and control flow facts. More advanced CoT-inspired prompting techniques include self-consistency (Wang et al., 2023d), where the prompt is run multiple times, and predictions are aggregated (e.g., a file is marked vulnerable if the model returns a positive classification in two out of three runs) (Tamberg and Bahsi, 2025), and Tree-of-Thought (ToT) prompting (Tamberg and Bahsi, 2025; Munson et al., 2025), which branches the reasoning process by generating multiple alternative steps per stage, followed by a selection of the most promising path.
Training (T2.2)
The adaptation of LLMs to software vulnerability detection tasks often follows the established pre-training and fine-tuning paradigm: models are trained from scratch (Pre-Training (T2.2.1)) using large code corpora to capture programming-specific syntax and semantics (Alrashedy et al., 2024; Ahmad et al., 2021; Buratti et al., 2020; Ding et al., 2023, 2024b; Hanif and Maffeis, 2022; Huang et al., 2022; Jararweh et al., 2025; Jiang et al., 2024c; Li et al., 2025e; Liu et al., 2024e; Wang et al., 2025f; Xu et al., 2023; Yu et al., 2022) and further fine-tuned for domain-specific applications. Fine-Tuning (T2.2.2) is the most widely adopted technique. Studies adopt a range of fine-tuning strategies that vary in their degree of parameter update and architectural design.
Full-Parameter Fine-Tuning (T2.2.2.1) refers to tuning all parameters of a pre-trained language model for the downstream task of vulnerability detection. Models are either fine-tuned by replacing the last output layer (Grishina et al., 2023) with a classification head (e.g., a multi-layer perceptron), which return the probability of vulnerability; or fine-tuned as a generative model where the response is matched to keywords such as “vulnerable” or “non vulnerable”. Full-parameter fine-tuning is widely adopted in the surveyed studies (Du et al., 2025b; Wu and Xiao, 2025; Liu et al., 2024a; Alam et al., 2025; Liu et al., 2026; Alrashedy et al., 2024; Ahmad et al., 2021; Aslan et al., 2025; Atiiq et al., 2024b, a; Bahaa et al., 2024; Buratti et al., 2020; Chan et al., 2023; Chen et al., 2024, 2023; Cheng et al., 2025; Cipollone et al., 2025; Curto et al., 2024b; d’Aloisio et al., 2025; Deng et al., 2024; Ding et al., 2023, 2024b, 2025; Ehrenberg et al., 2024; Ferrag et al., 2025b; Fu et al., 2024; Fu and Tantithamthavorn, 2022; Fu et al., 2023a; Gajjar et al., 2025a; Gao et al., 2023; Grishina et al., 2023; Prabith et al., 2024; Guo et al., 2024b; G”̈urfidan, 2024; Halder et al., 2025; Han and Zheng, 2025; Hanif and Maffeis, 2022; Huang et al., 2022; Ibanez-Lissen et al., 2025; Jiang et al., 2024c; JianJie and Le, 2023; Ji et al., 2024; Kaanan et al., 2024; Kalouptsoglou et al., 2024a, b; Kim et al., 2022; Kong et al., 2024; Lan et al., 2025; Le et al., 2024; Li et al., 2021, 2025k; Lin et al., 2022a; Li et al., 2025b, j; Liu et al., 2025a, 2024e, b; Lu et al., 2023; Luo et al., 2021; Mächtle et al., 2025; Maligazhdarova et al., 2025; Mamede et al., 2022; Mechri et al., 2025; Mei et al., 2023; Mock et al., 2025; Nguyen et al., 2025b; Ni et al., 2023b; Omar, 2023; Omar and Shiaeles, 2023; Panichella, 2025; Purba et al., 2023; Qi et al., 2024; Quan et al., 2025; Rahman et al., 2024; Ridoy et al., 2024; Risse and B”̈ohme, 2024; Russo et al., 2025; Safdar et al., 2025b; Shestov et al., 2025; Shi et al., 2024; Shimmi et al., 2024b, a; Simoni et al., 2025; Singh et al., 2022; Steenhoek et al., 2025b, 2023a, 2023b; Sultana et al., 2024; Sun et al., 2024; Su and Wu, 2023; Takieldeen et al., 2025; Thapa et al., 2022; Tian et al., 2025; Torkamani et al., 2025; Wang et al., 2025c, d, f; Wen et al., 2025, 2024a, 2024b, 2023; Xiong and Dong, 2024; Xu et al., 2023; Yang et al., 2025a; Yin et al., 2024; Yuan et al., 2022; Zeng et al., 2023; Zhang et al., 2024c; Zhao and Guan, 2025; Zhao et al., 2024a; Zheng et al., 2025, 2024; Zhou et al., 2024a; Zhu et al., 2025b, 2022; Ziems and Wu, 2021; Nguyen et al., 2025a), particularly with tiny models, see Figure 5. While full-parameter fine-tuning is computationally intensive and less reusable across tasks, it allows the model to specialize in vulnerability detection by adapting all internal representations.
A specific variant of fine-tuning is Instruction-Tuning (T2.2.2.1.1), where a language model is fine-tuned on labeled input-output pairs annotated with natural language instructions (Bappy et al., 2025; Du et al., 2024; Islam et al., 2024b; Jiang et al., 2025b; Li et al., 2025c; Mao et al., 2025; Shu et al., 2025; Steenhoek et al., 2025a; Tian et al., 2024; Widyasari et al., 2025; Yusuf and Jiang, 2024; Zhang et al., 2025b, 2024b), e.g., ”Instruction: Detect whether the following code contains vulnerabilities” (Jiang et al., 2025b).
This method allows the model to better follow human-written prompts by learning general patterns for task completion.
In (Widyasari et al., 2025), Widyasari et al. use GPT-4o for generating the instructions.
Instruction-tuning has shown promise in enhancing performance on unseen instructions and is often combined with parameter-efficient (rather than full) fine-tuning techniques to reduce training cost (Du et al., 2024; Jiang et al., 2025b; Li et al., 2025c; Mao et al., 2025; Shu et al., 2025; Tian et al., 2024; Yusuf and Jiang, 2024; Zhang et al., 2025b, 2024b).
Key drawbacks of full-parameter fine-tuning are its high computational cost and the potential harm to the model’s generalization ability. To address these drawbacks, several studies adopt Parameter-Efficient Fine-Tuning (PEFT) (T2.2.2.2) techniques, which adapt only a small subset of model parameters or learn external modules for new tasks while keeping most of the pre-trained parameters frozen. To categorize the PEFT methods, we adopt the taxonomy by Han et al. (Han et al., 2024), covering (i) selective, (ii) additive, and (iii) reparameterized PEFT techniques.
In Selective PEFT (T2.2.2.2.1), only a small subset of model parameters is selected for fine-tuning, while keeping the other parameters frozen. In the surveyed studies, the final (Zhang et al., 2025a) or upper model layer(s) (Chen et al., 2025b) were fine-tuned to generate contextual embeddings.
Additive PEFT (T2.2.2.2.2) techniques introduce a small number of trainable parameters that are strategically positioned within the model architecture. During fine-tuning, only the weights of these additional modules or parameters are updated, reducing storage, memory, and computational resource requirements. Prominent additive PEFT techniques are adapter-tuning and prompt-tuning.
Adapter-Tuning (T2.2.2.2.2.1) inserts small trainable neural modules, i.e., adapters, into a frozen pre-trained model. During fine-tuning, only the adapter parameters are updated, while the model’s parameters remain unchanged (Akli et al., 2025; Li et al., 2025e). Rather than employing default configurations that use the same adapter settings across all layers, Akli et al. (Akli et al., 2025) present AutoAdapt to automatically discover task-specific, layer-wide adapter configurations, allowing each layer to adopt distinct parameters.
Prompt-Tuning (T2.2.2.2.2.2) involves learning prompts to improve model performance on a downstream task. Two main prompt forms are distinguished: hard prompts, which are manually crafted sequences of tokens appended to the input, and soft prompts, which are learned continuous embeddings that serve as virtual tokens prepended to the input. Prompt-tuning keeps the model weights frozen while tuning only the parameters of the soft prompts (Chang et al., 2024; Chen and Zhang, 2025; Feng et al., 2025; Lu et al., 2023; Ren et al., 2024). Various adaptations of prompt-tuning have been investigated. Ren et al. (Ren et al., 2024) integrate prompt-tuning with a reinforcement learning framework. Feng et al. (Feng et al., 2025) explore graph-enhanced prompt-tuning, which incorporates structural code information by embedding graph features into soft prompts.
Apart from the methods mentioned above, another additive approach has been used that strategically trains additional parameters during the fine-tuning process (Additive-Other (T2.2.2.2.2.3)): Li et al. (Li et al., 2025d) introduce a vulnerability steering vector that represents the concept of vulnerability in the representation space. The vector is injected into the activation values of the corresponding layer, guiding the model’s behavior without modifying all its parameters.
Reparameterized PEFT (T2.2.2.2.3) techniques transform a model’s architecture by modifying how its parameters are represented and trained. This typically involves replacing large components of the model with smaller, low-rank versions during training, reducing the number of parameters that need to be updated. At inference time, the model can be converted to its original parameterization (Han et al., 2024).
The most widely adopted reparameterization technique in the surveyed studies is categorized as Low-Rank Decomposition (T2.2.2.2.3.1): LoRA (Low-Rank Adaptation) (Hu et al., 2022). LoRA injects trainable low-rank decomposition matrices into the attention layers of a frozen language model, enabling task-specific adaptation with significantly fewer parameters (Prasad et al., 2025; Alqarni and Azim, 2025; Aslan et al., 2025; Cao et al., 2024; Chen and Zhang, 2025; Curto et al., 2024a; Du et al., 2024; Gajjar et al., 2025b; Gonçalves et al., 2025; Guo et al., 2024b; Jararweh et al., 2025; Jiang et al., 2025b; Lekssays et al., 2025; Li et al., 2025j; Liu et al., 2025a; Li et al., 2025c; Luo et al., 2025; Mao et al., 2025; Shestov et al., 2025; Shu et al., 2025; Wang et al., 2024b; Wen et al., 2025; Weyssow et al., 2025; Yusuf and Jiang, 2024; Zhang et al., 2024b; Zhou et al., 2024a; Zhu et al., 2024).
LoRA Derivatives (T2.2.2.2.3.2) have also been explored to improve efficiency and adaptability. Tian et al. (Tian et al., 2024) use PiSSA (Principal Singular Values and Singular Vectors Adaptation) (Meng et al., 2024), which shares the same architecture as LoRA but adopts a different initialization method. Ibanez-Lissen et al. (Ibanez-Lissen et al., 2025) use GaLore (Gradient Low-Rank Projection) (Zhao et al., 2024c), a memory-efficient training strategy that allows full-parameter training while requiring less memory than common low-rank approaches. Another efficient PEFT design is QLoRA (Quantized LoRA) (Dettmers et al., 2023), which applies LoRA to a quantized version of the language model to further reduce memory consumption. QLoRA has been used in multiple surveyed studies to efficiently fine-tune large models for code vulnerability tasks (Zhang et al., 2025b; Sultana et al., 2024; Yang et al., 2024a; Ouchebara and Dupont, 2025; Sun et al., 2025a; Wei et al., 2025).
Learning Paradigms (T2.3)
Beyond standard supervised fine-tuning, several advanced learning paradigms have been employed to enhance the performance and generalization capabilities of vulnerability detection models, e.g., contrastive learning, causal learning, multi-task learning, knowledge distillation, continual learning, and reinforcement learning.
Contrastive Learning (T2.3.1) aims to learn more discriminative representations by encouraging the model to pull semantically similar instances closer and push dissimilar ones apart in the representation space, thereby improving the model’s ability to distinguish between structurally similar but semantically different code, i.e., secure versus vulnerable code. Several studies adapt contrastive learning techniques to the vulnerability detection domain (Jiang et al., 2022; Wu and Xiao, 2025; Ni et al., 2023b; Wang et al., 2024b; Du et al., 2023; Ding et al., 2025; Ji et al., 2024; Gajjar et al., 2025a; Li et al., 2025e; Wang et al., 2025f). Notably, Du et al. (Du et al., 2023) propose mutual nearest neighbor contrastive learning to align the source and target domains. Ding et al. (Ding et al., 2025) introduce class-aware contrastive learning, which minimizes similarity only between samples with different class labels to improve type-specific discrimination. Ji et al. (Ji et al., 2024) present a hierarchical contrastive learning framework to bring vector representations of related CWEs closer together, combining supervised and self-supervised contrastive losses to promote geometric spread and improved inter-class separation.
Causal Learning (T2.3.2) guides models to focus on genuine cause-and-effect relationships rather than relying on spurious correlations in the code, e.g., variable or API names that are not actual causes of vulnerabilities. Rahman et al. (Rahman et al., 2024) identify such spurious features via semantic-preserving perturbations (e.g., variable renaming or dead-code injection) and then systematically remove their influence using causal inference techniques. Specifically, they simulate interventions, i.e., asking how the model would behave if certain spurious features were actively changed, and block non-causal paths by conditioning on known spurious features. This strategy allows the model to rely on causally relevant features and, thereby, improve robustness and generalization to unseen or perturbed code scenarios, such as different projects or naming conventions.
Multi-Task Learning (T2.3.3) has emerged as an effective paradigm for improving generalization by jointly training on related tasks, thereby reducing overfitting and enhancing model robustness across a range of tasks. Several studies adopt multi-task learning to enhance vulnerability analysis: Fu et al. (Fu et al., 2024) fine-tune under the multi-task setting for predicting CWE-ID and CWE-type, while Chen et al. (Chen et al., 2025b) combine CWE-ID classification with line-level vulnerability detection. Curto et al. (Curto et al., 2024b) perform binary vulnerability detection alongside multi-class vulnerability categorization. Similarly, Steenhoek et al. (Steenhoek et al., 2025b) fine-tune for simultaneous binary classification, multi-class vulnerability type prediction, and localization. Islam et al. (Islam et al., 2023) combine vulnerability classification with description. Du et al. (Du et al., 2024) apply multi-task instruction-tuning for vulnerability detection, localization, and interpretation of root causes; while Yang et al. (Yang et al., 2024a) focus on multi-task fine-tuning for vulnerability detection, explanation, and repair. Russo et al. (Russo et al., 2025) leverage shared information between self-admitted technical debt (SATD) and vulnerabilities to jointly detect both issues, Alqarni and Azim (Alqarni and Azim, 2025) combine the analysis of code, system calls, and bitstreams, and Ding et al. (Ding et al., 2024b) use multi-task pre-training to jointly learn static and dynamic code properties.
Knowledge Distillation (T2.3.4) is a model compression strategy in which a smaller model, i.e., the student, is trained to replicate the behavior of a larger model, i.e., the teacher. By learning from the teacher’s output distributions, the student achieves comparable performance while significantly reducing computational cost, making it suitable for resource-constrained settings. Some studies directly adopt pre-trained distilled models such as DistilBERT for vulnerability detection (d’Aloisio et al., 2025; Liu et al., 2025a; Lin et al., 2022a), while other studies apply knowledge distillation explicitly, using teacher and student LLMs (Omar, 2023; Shi et al., 2024; Wei et al., 2025; Weyssow et al., 2025; Panichella, 2025; Liu et al., 2026), or a setup with CNN and GNN models to distill knowledge into or from a LLM backbone (Nguyen et al., 2025b; Liu et al., 2025b; Fu et al., 2023a; Han and Zheng, 2025). A more advanced approach is presented by Fu et al. (Fu et al., 2023a), who introduce a hierarchical knowledge distillation framework. The CWE label space is split into multiple sub-distributions based on semantic similarity, and individual CNN teacher models are trained on each subset. A student language model then learns to generalize across the teacher outputs, improving its ability to handle multi-class vulnerability detection tasks. Han et al. (Han and Zheng, 2025) follow this approach with an enhanced memory mechanism. Panichella (Panichella, 2025) combines metamorphic testing with many-objective optimization for distillation of LLMs for code, measuring model robustness across semantically equivalent code variants.
Retraining a model on all known vulnerabilities each time a new vulnerability is disclosed would demand considerable computational resources and time. Continual Learning (T2.3.5) enables models to incrementally learn new knowledge over time, such as new programming languages or up-to-date vulnerability type data (Wei et al., 2025), without forgetting what has already been learned (Catastrophic Forgetting). Gao et al. (Gao et al., 2023) propose an adaptive regularization strategy that preserves the most representative (i.e., informative and diverse) samples in each dataset for model retraining under project-level, language-level, and time-level continual learning settings. Similarly, Zhang et al. (Zhang et al., 2024c) present an incremental learning approach that allows the model to incorporate new programming languages without degrading performance on those it has already learned. Tian et al. (Tian et al., 2025) introduce a training-free parameter fusion method that merges independent classification heads fine-tuned on different vulnerability types, allowing the fused model to adapt to new categories while preserving multi-class classification capabilities.
Reinforcement Learning (RL) (T2.3.6) introduces a reward-based feedback loop, allowing models to optimize sequential decision-making processes. Several studies focus on refining the model’s reasoning capabilities. Weyssow et al. (Weyssow et al., 2025) employ Reinforcement Learning from AI Feedback (RLAIF) to distill structured reasoning, reinforcing informative and coherent while penalizing misleading reasoning. Similarly, Simoni et al. (Simoni et al., 2025) integrate an RL-based loss function to fine-tune instruction-based LLMs. Ren et al. (Ren et al., 2024) introduce a reward mechanism with prompt tuning that encourages the model to discover prompts that maximize detection performance. Beyond reasoning, RL is utilized to optimize input and feature extraction. Jiang et al. (Jiang et al., 2025a) propose an RL framework to learn vulnerability-relevant structures within code snippets. Zhao et al. (Zhao and Guan, 2025) use RL to generate high-quality adversarial examples, selecting optimal perturbations that minimize the model’s vulnerability prediction probability to create harder training samples.
Other Data-Centric (T2.3.7) learning techniques have been explored in the surveyed studies to improve model performance, e.g., Positive and Unlabeled (PU) learning, active learning, and adversarial training.
PU learning tackles the issue of limited and noisy labels by training models using only positively (vulnerable) labeled and unlabeled data (Wen et al., 2023; Kong et al., 2024).
Forms of active learning aim to reduce labeling costs by strategically selecting the most informative samples for annotation (Alrashedy et al., 2024; Lan et al., 2025; Cheng et al., 2025). For example, Lan et al. (Lan et al., 2025) combine dataset maps with active learning to prioritize valuable samples while filtering out those that may harm performance.
Similarly, in self-paced learning, the most suitable data is selected for each training epoch, from easy to hard learning samples (Cheng et al., 2025).
Adversarial training enhances model robustness by introducing crafted perturbed code examples during training, improving robustness to noise and attacks (Yu et al., 2022; Chen et al., 2025b; Zhao and Guan, 2025; Du et al., 2024; Liu et al., 2026).
4.4.3. Orchestration Technique (T3)
While adaptation techniques focus on optimizing the model’s internal parameters or input context, orchestration techniques focus on the design of the inference process. Rather than relying on a simple input-model-output pipeline, orchestration techniques embed the LLM into advanced system designs that utilize multiple generation steps, iterative refinement, or collaborative workflows. We categorize orchestration techniques into Multi-Step, Verification, Agentic, Ensemble, and Controller-based systems.
Some studies engage the model in a Multi-Step (T3.1) analysis, rather than issuing a single prompt for vulnerability detection (Espinha Gasiba et al., 2024; Huynh et al., 2025; Liu et al., 2024f, 2023; Ullah et al., 2024; Wang et al., 2025b; Yu et al., 2025; Jiao et al., 2025; Yang et al., 2025b; Mhatre et al., 2025; Peng et al., 2025; Tsai et al., 2025a; Ding et al., 2025). A common strategy within this approach is to decouple comprehension from detection, prompting the model to first analyze the code’s functionality or intent before querying for vulnerabilities (Espinha Gasiba et al., 2024; Huynh et al., 2025; Liu et al., 2024f, 2023; Jiao et al., 2025; Yang et al., 2025b; Peng et al., 2025; Wang et al., 2025b). Beyond functional analysis, Yu et al. (Yu et al., 2025) task the model to first generate supplementary code files based on commit messages, subsequently using this supplementary context for the final detection. Mhatre et al. (Mhatre et al., 2025) follow a debugging style, first identifying code issues, followed by a deeper, targeted analysis of specific issues. Tsai et al. (Tsai et al., 2025a) propose a sequential approach that dynamically selects the optimal strategy, choosing between direct zero-shot classification, RAG-based prediction, or multi-agent collaboration, guided by the model’s internal confidence and collaborative signals.
To mitigate hallucinations and improve precision, studies increasingly adopt Verification (T3.2) mechanisms where models are prompted to review and improve generated outputs. Verification extends the standard prompt-response interaction by re-prompting the model to evaluate, validate, or refine a prediction. In single-model setups, the model is typically prompted to assess its prior response by verifying the correctness of its output (Xia et al., 2025) or by recursively critiquing and improving (Tamberg and Bahsi, 2025; Du et al., 2024; Sayagh and Ghafari, 2025). In multi-model settings, verification is implemented as an interactive process between models (collaborative refinement), possibly under different roles. For example, using the output of one model to prompt another model with corrective cues such as ”An expert has found… please recheck it” (Wang et al., 2025d; Ambati et al., 2024; Tsai et al., 2025b; Widyasari et al., 2024; Cao and Jun, 2024; Eberhardt and Milánkovich, 2024).
Agentic (T3.3) systems embed LLMs into autonomous, decision-making loops, following a higher-level behavioral pattern involving multi-step planning, tool use, and self-directed reasoning. Agents may operate individually or as part of multi-agent systems (Cao and Jun, 2024; Yildiz et al., 2025; Widyasari et al., 2025; Ahmed et al., 2025; Eberhardt and Milánkovich, 2024; Li et al., 2025h; Tsai et al., 2025a; Wang et al., 2025e; Tsai et al., 2025b), where agents specialize in subtasks or roles, interact with tools, and collaboratively refine outputs. For example, Yildiz et al. (Yildiz et al., 2025) employ Reasoning and Acting (ReAct) agents (Yao et al., 2023), which follow a thought-action-observation cycle, for vulnerability detection in code repositories. Widyasari et al. (Widyasari et al., 2025) present a courtroom-inspired multi-agent framework with four role-specific agents: security researcher, code author, moderator, and review board. The agents collaborate to assess code vulnerabilities through iterative analysis, argumentation, summarization, and a final verdict based on collective reasoning. Ahmed et al. (Ahmed et al., 2025) propose a multi-agent pipeline consisting of a normalization, planning, context, detection, and validation agent. Sayagh and Ghafari (Sayagh and Ghafari, 2025) use multiple agentic teams, with a lister searching for potential CWEs in the input code, a reviewer requesting missing CWEs, further agents identifying external code context, and a final agent deciding on each CWE based on gathered context.
Ensemble (T3.4) approaches combine the representations or predictions of diverse models to mitigate individual model limitations and biases. To leverage complementary feature representations, e.g., semantic, contextual, and syntactic features, studies fuse representations from different pre-trained models (Du et al., 2025b; Wu and Xiao, 2025; Mahmud et al., 2025; Ridoy et al., 2024; Shimmi et al., 2024a; Xiong and Dong, 2024; Sun et al., 2023; JianJie and Le, 2023). For example, Ridoy et al. (Ridoy et al., 2024) employ an ensemble stacking approach to synthesize the semantic understanding of CodeBERT, the structural representations of GraphCodeBERT, and the cross-modal capabilities of UniXcoder. Distinct from standard feature fusion, Tian et al. (Tian et al., 2025) introduce a parameter fusion method that merges independent classification heads trained on distinct vulnerability datasets, allowing the model to adapt to new vulnerability types. Ensemble approaches further include decision-level aggregation via voting mechanisms to combine predictions from multiple LLMs (Widyasari et al., 2024; Zhang et al., 2025b) or ensembles of LLMs and other tools (Zhang et al., 2025b; Zhou et al., 2024a).
Moving beyond static inference pipelines, some studies use LLMs in a Controller (T3.5) or collaboration setting. A central model or routing mechanism analyzes the input code and coordinates downstream resources, selecting, e.g., the most appropriate expert models or prompts. Yang et al. (Yang et al., 2025a) and Wu and Xiao (Wu and Xiao, 2025) implement a Mixture-of-Experts (MoE) framework where a router dynamically assigns input code to the most appropriate vulnerability experts. Similarly, Wang et al. (Wang et al., 2025e) employ a meta-agent to route code to specialized analyzers. Peng et al. (Peng et al., 2025) introduce a Mixture-of-Prompts architecture that activates specific prompt experts (dynamic prompt pipelines) based on code features of the code under detection. Li et al. (Li et al., 2025c) simulate team collaboration involving a leader and various expert models specialized on syntactic features; the leader orchestrates the process and derives the final detection from the experts’ analyses. Farr et al. (Farr et al., 2025) extend the routing logic to human collaborators. They flag code as automatic quarantine, cleared for deployment, or human analyst evaluation, where code with high model uncertainty is routed to human experts.
5. Vulnerability Datasets
In this SLR, we set an emphasis on investigating the datasets used in the adaptation and evaluation of LLM-based vulnerability detection approaches. Datasets play a central role: Their quality, structure, and coverage of diverse vulnerability types directly affect model performance and generalization. Further, the choice of dataset does not merely benchmark performance. A critical challenge in this domain is data leakage, where data samples overlap with the corpora used to pre-train the model, potentially leading to inflated performance.
Furthermore, the use of datasets has evolved beyond standard training. With the rise of advanced prompting techniques, datasets now serve as a dynamic input context, providing exemplars for in-context prompting or knowledge retrieval. This shift also changes which parts of the data are used. While deep learning-based models often relied solely on code syntax, LLMs actively leverage the natural language components of datasets, such as commit messages, vulnerability descriptions, and code comments, to support high-level explanation and reasoning tasks. Therefore, systematically comparing datasets is essential to understand their suitability for LLM-based detection, assess their limitations, and ensure meaningful evaluations.
To systematically analyze the datasets used in the surveyed studies, we extend the presented taxonomy to cover datasets. We outline the taxonomy based on the datasets used most commonly for training or evaluation in the surveyed studies, discussing type, granularity, source, and labeling. Further, we investigate the CWE coverage and diversity of selected datasets, discussing limitations such as long-tail distributions and a lack of representative vulnerability coverage. Finally, we discuss trends in dataset usage, emphasizing the comparability of evaluations across different use cases.
5.1. Taxonomy and Commonly Used Datasets
We categorize vulnerability datasets according to type, granularity, source, and labeling, as shown in Figure 6. Similar to related surveys (Shereen et al., 2024; Ferrag et al., 2025a; Taghavi Far and Feyzi, 2025; Sheng et al., 2025; Guo et al., 2024a; Guo and Bettaieb, 2023), we provide a comprehensive overview of commonly used datasets in the surveyed studies as well as selected, more recent datasets suited to address open challenges of LLM-based approaches, refer to Table 4. We briefly introduce these datasets along with the taxonomy and discuss the resulting practical challenges and biases in dataset construction. In addition, we compare the selected datasets based on further characteristics such as their programming languages covered, as well as quantitative characteristics, i.e., dataset size, the share of samples representing vulnerabilities, and the number of distinct CWEs covered (cf. Table 5) as a proxy for label diversity and comprehensiveness.
5.1.1. Dataset Type
We distinguish three dataset types: (i) synthetic, (ii) real-world, and (iii) mixed.
Synthetic datasets are artificially generated, often by injecting predefined vulnerability patterns into source code. The synthetic datasets Juliet C/C++ (Metrics and Evaluation, 2017a) and Juliet Java (Metrics and Evaluation, 2017b) (subsets of the Software Assurance Reference Dataset (SARD) (Metrics and Evaluation, 2006)) provide comprehensive sets of test cases and entire test suites. These test suites are widely used to benchmark both traditional static analysis tools and deep learning-based vulnerability detectors. A key strength of synthetic datasets lies in their reliable labels, as they inject known vulnerable patterns into code, thereby ensuring clear ground truth annotations. However, this injection-based approach often results in limited code diversity and reduced realism, as the code samples tend to be simple and may not reflect the complexity of real-world source code. Further, the manual effort required to represent a diverse set of vulnerabilities can be labor-intensive.
Real-World datasets are sourced from open-source software projects, e.g., Devign (Zhou et al., 2019), or curated vulnerability databases, e.g., Big-Vul (Fan et al., 2020) and MegaVul (Ni et al., 2024). These datasets are more suited for a realistic evaluation, capturing the challenges posed by detecting software vulnerabilities in real-world projects. However, due to the efforts required for crawling and parsing software projects, real-world datasets often suffer from label noise.
Mixed datasets combine real-world code with synthetic data, leveraging augmentation to enrich or balance the dataset. For example, the SARD (Metrics and Evaluation, 2006) dataset contains a mix of production, synthetic, and academic vulnerable samples. VulDeePecker (Li et al., 2018) combines samples from SARD and the National Vulnerability Database (NVD) (NVD, 2025). Draper (Russell et al., 2018) combines synthetic functions from the Juliet test suite with open-source code from software packages and public repositories. Mixed datasets offer a practical trade-off between realism and control, making them particularly valuable for covering underrepresented or rare vulnerability types.
5.1.2. Dataset Granularity
Vulnerabilities can occur at various structural levels within the software, from single lines to entire files. Dataset granularity refers to the level at which code is labeled or segmented within the dataset. Following Shereen et al. (Shereen et al., 2024), we differentiate coarse, medium, and fine-grained levels.
Coarse-grained datasets operate at the Project, Commit, or File level. ReposVul (Wang et al., 2024a) focuses on extracting dependencies on multiple granularity levels, including repository-level and file-level. Similarly, CrossVul (Nikitopoulos et al., 2021b) labels samples at the file level, thereby aligning closely with practical use cases in vulnerability detection. While this level of granularity offers realistic scenarios for analysis, it also introduces a high risk of noise due to the large code context surrounding the vulnerability.
Medium-grained datasets segment code at the Program, Slice or Function level. The largest share of commonly used vulnerability datasets is labeled at the function level (Metrics and Evaluation, 2006; Russell et al., 2018; Zhou et al., 2019; Fan et al., 2020; Zheng et al., 2021; Chakraborty et al., 2022; Bhandari et al., 2021; Chen et al., 2023; Ding et al., 2024a; Du et al., 2025a; Ni et al., 2024; Li et al., 2025k), cf. Table 4. The VulDeePecker dataset, also known as the Code Gadget Database (CGD), represents program slices as code gadgets composed of a number of program statements that are semantically related to each other through data or control dependencies. While this granularity supports a more focused analysis, it may miss critical contextual information when vulnerabilities span multiple functions or require interprocedural analysis.
Fine-grained datasets label vulnerabilities at the Line level, enabling precise vulnerability localization. Big-Vul, D2A (Zheng et al., 2021), FormAI, and ReposVul provide annotations at the line-level. Such a narrow scope, however, omits the broader context required to detect certain vulnerabilities, especially those involving complex data flows or control dependencies.
| Dataset | Year | Type | Granularity | Source | Labeling | #Used | Resource |
| SARD (Metrics and Evaluation, 2006) | 2006 | Mixed | File | Open-Source | Synthetically | 32 | (Metrics and Evaluation, 2006) |
| Juliet C/C++(Metrics and Evaluation, 2017a) | 2017 | Synthetic | File | Open-Source | Synthetically | 9 | (Metrics and Evaluation, 2017a) |
| Juliet Java (Metrics and Evaluation, 2017b) | 2017 | Synthetic | File | Open-Source | Synthetically | 6 | (Metrics and Evaluation, 2017b) |
| VulDeePecker (Li et al., 2018) | 2018 | Mixed | Slice | Constructed | Security Vendor | 11 | (CGCL Codes, 2017) |
| Draper (Russell et al., 2018) | 2018 | Mixed | Function | Constructed | Tool, Synthetically | 11 | (Kim and Russell, 2018) |
| Devign (Zhou et al., 2019) | 2019 | Real | Function | Open-Source | Developer | 69 | (Pinconschi, 2020) |
| Big-Vul (Fan et al., 2020) | 2020 | Real | Function, Line | Collected | Security Vendor | 63 | (Fan, 2020) |
| D2A (Zheng et al., 2021) | 2021 | Real | Function, Line | Open-Source | Tool | 9 | (IBM, 2021) |
| ReVeal (Chakraborty et al., 2022) | 2021 | Real | Function | Collected | Developer | 40 | (VulDetProject, 2020) |
| CVEfixes (v1.0.8) (Bhandari et al., 2021) | 2021 | Real | Function, Commit, File | Collected | Security Vendor | 19 | (Moonen and Vidziunas, 2024) |
| CrossVul (Nikitopoulos et al., 2021b) | 2021 | Real | File | Collected | Security Vendor | 4 | (Nikitopoulos et al., 2021a) |
| SecurityEval (v2.1) (Siddiq and Santos, 2022) | 2022 | Mixed | Program | Constructed | Synthetically, Tool | 4 | (Security & Software Engineering Research Lab at University of Notre Dame, 2022) |
| DiverseVul (Chen et al., 2023) | 2023 | Real | Function | Collected | Developer | 27 | (Wagner Research Group, 2025) |
| SVEN (He and Vechev, 2023) | 2023 | Real | Program | Constructed | Developer | 8 | (SRI Lab, ETH Zurich, 2023) |
| FormAI* (Tihanyi et al., 2023) | 2023 | Synthetic | Program, Line | Constructed | Tool | 2 | (FormAI Dataset, 2023) |
| ReposVul* (Wang et al., 2024a) | 2024 | Real | Project, File, Function, Line | Collected | Tool | 3 | (Wang, 2024) |
| PrimeVul (Ding et al., 2025) | 2024 | Real | Function | Constructed | Security Vendor | 27 | (DLVulDet, 2024) |
| PairVul* (Du et al., 2025a) | 2024 | Real | Function | Collected | Security Vendor | 1 | (KnowledgeRAG4LLMVulD, 2024) |
| MegaVul* (2024/04) (Ni et al., 2024) | 2024 | Real | Function | Collected | Security Vendor | 2 | (Icyrockton, 2023) |
| CleanVul* (Li et al., 2025k) | 2024 | Real | Function | Open-Source | Developer, Tool | 3 | (Li, 2024) |
5.1.3. Dataset Source
To categorize the source of the dataset, we adapt the four categories defined by Hou et al. (Hou et al., 2024), i.e., (i) open-source, (ii) collected, (iii) constructed, and (iv) closed-source.
Open-Source datasets are derived from public data collections accessible through open-source platforms or repositories, e.g., Github repositories. Examples for open-source datasets are SARD, Devign, D2A, and CleanVul (Li et al., 2025k). Their accessibility ensures transparency and reproducibility, making them widely used in academic research. However, datasets sourced directly from GitHub repositories often lack reliable ground-truth labels and are sensitive to repository selection and curation quality, i.e., the specific projects chosen and how samples are filtered, labeled, and balanced, leading to biased, noisy, or unrepresentative data.
Collected datasets are those scraped, mined, and extracted by researchers from various sources such as security trackers or related databases (e.g., NVD (NVD, 2025), CVE (CVE, 2025b), or CWE database (MITRE, 2025a)). MegaVul, for example, is collected by crawling the CVE database along with CVE-related open-source projects hosted across Git-based platforms. While authoritative vulnerability records provide structured metadata and standardized vulnerability classifications, the mapping of records to source code can be non-trivial and may miss undocumented or undisclosed vulnerabilities.
Constructed datasets are created by modifying or augmenting one or multiple other (collected) datasets, either manually or using semi-automatic methods, to better align with domain-specific objectives, e.g., focusing on specific CWEs or programming languages. Constructed datasets include the mixed datasets VulDeePecker and Draper. Further, SecurityEval (Siddiq and Santos, 2022) is constructed from sources such as GitHub’s CodeQL documentation, the CWE database, Sonar static analyzer rules, and prior Copilot-based studies. Additional samples were crafted by the authors. SVEN (He and Vechev, 2023) refines and validates samples from Big-Vul, CrossVul, and VUDENC (Wartschinski et al., 2022), focusing on nine common CWE types. The PrimeVul dataset (Ding et al., 2024a) merges and de-duplicates data from Big-Vul, CrossVul, CVEfixes (Bhandari et al., 2021), and DiverseVul (Chen et al., 2023).
Closed-Source datasets are obtained from commercial or industrial entities. Closed-sourced datasets are absent from the listed commonly used datasets. Notably, only few surveyed studies (Ding et al., 2023; Mock et al., 2025; Peng et al., 2025; Xue et al., 2025) use an industrial, closed-sourced dataset. Although such datasets offer significant potential for research targeting real-world deployment scenarios, they are often subject to restrictions that limit the publication and sharing of company-internal data. Such restricted access limits reproducibility, public benchmarking, and broader adoption within the research community.
5.1.4. Dataset Labeling
Accurate labeling is essential for constructing high-quality vulnerability datasets and reliable evaluations. To categorize the labeling of datasets, we adopt the four main label origin categories by Croft et al. (Croft et al., 2023): (i) security vendor-provided, (ii) developer-provided, (iii) tool-created, and (iv) synthetically created.
Security Vendor-provided labels are derived from curated vulnerability databases maintained by security vendors, i.e., organizations that collect and standardize information on disclosed vulnerabilities from various advisories. An example database is the NVD. Many entries include links to corresponding patches on security trackers or GitHub sites, enabling the mapping of vulnerabilities to real-world source code. For example, the Big-Vul dataset was collected by crawling the public CVE database and linking the CVE entries to related source code changes in GitHub projects. This labeling strategy is reliable and consistent with real-world vulnerability disclosures. However, it is limited to publicly disclosed cases and includes historical data with labels that may not align with up-to-date vulnerability mappings (Wang et al., 2024a).
Developer-provided labels are extracted directly from a project’s development history or issue tracker systems on platforms such as GitHub. A commonly used strategy involves vulnerability fixing commits (VFCs), identifying vulnerable code before a fix and secure code after a fix. VFCs are often identified via project references by the NVD or keyword search. The Devign dataset (also referred to as FFmpeg and QEMU) consists of vulnerable functions extracted from VFCs identified via keyword-based filtering and manually labeled by security researchers. Devign is also part of the CodeXGLUE benchmark (Lu et al., 2021). Contrary to manually labeling VFCs, relying solely on VFCs often results in noisy labels due to the simplifying assumption that all pre-commit code is vulnerable and all post-commit code is secure (Croft et al., 2023). In practice, VFCs may contain unrelated changes such as refactoring or test updates, and fixes are sometimes distributed across multiple commits. Further, labeling based solely on modified lines may overlook vulnerable context and cross-line dependencies. To address these challenges, newer datasets propose refined VFC labeling strategies. The PrimeVul dataset (Ding et al., 2024a) uses two labeling strategies: PRIMEVUL-ONEFUNC, which marks a function as vulnerable only if it is the only function modified by a VFC, and PRIMEVUL-NVDCHECK, which cross-references CVE descriptions with modified functions. CleanVul addresses the label noise using an LLM, which assigns a confidence score from 0 (no vulnerability detected) to 4 (very high likelihood of vulnerability) to identify vulnerability-fixing changes and filter out unrelated modifications such as test updates or refactorings.
Tool-created labels are automatically generated using, e.g., static analyzers or formal verification tools. For instance, in the Draper dataset, labels are generated using static analyzers and then manually mapped to binary vulnerability labels and CWEs by a team of security experts. Similarly, the D2A dataset applies differential analysis to VFC version pairs from several open-source projects using static analysis tools, identifying disappearing bugs likely to present true vulnerabilities. Notably, the FormAI dataset (Tihanyi et al., 2023) consists of compilable and self-contained C programs generated using GPT-3.5-Turbo, incorporating varying levels of complexity to simulate a broad spectrum of coding patterns. Labels are derived using formal verification. ReposVul jointly employs LLMs for evaluating the relevance between code changes and vulnerability fixes and static analysis tools for checking the vulnerabilities in the changes. While tools enable large-scale dataset creation, they are often criticized for generating a high volume of false positives. Static analysis tools, in particular, produce numerous warnings, many of which originate from low-confidence detections (Croft et al., 2023). Formal verification offers higher detection reliability but at higher computational costs.
Synthetically created labels result from injecting known vulnerable patterns into code, generating controlled artificial samples, e.g., as used by the synthetic Juliet test suites. This method is useful for training on rare vulnerability types and balancing datasets. However, such patterns may oversimplify the complexity of real-world vulnerabilities.
| Dataset | Programming Language | Granularity | Size | %Vuln | Multi- Class | #CWEs |
| SARD (Metrics and Evaluation, 2006) | C/C++, C#, Java, PHP | File | 450,000+ | n/s | ✓ | 150+ |
| Juliet C/C++(Metrics and Evaluation, 2017a) | C/C++ | File | 64,099 | n/s | ✓ | 118 |
| Juliet Java (Metrics and Evaluation, 2017b) | Java | File | 28,881 | n/s | ✓ | 112 |
| VulDeePecker (Li et al., 2018) | C/C++ | Slice | 61,638 | 28.7 | ✓ | 2 |
| Draper (Russell et al., 2018) | C/C++ | Function | 1,274,366 | 6.5 | ✓ | 4 |
| Devign (Zhou et al., 2019) | C/C++ | Function | 48,687 | 48.0 | ✗ | n/a |
| Big-Vul (Fan et al., 2020) | C/C++ | Function, Line | 264,919 | 4.5 | ✓ | 91 |
| D2A (Zheng et al., 2021) | C/C++ | Function, Line | 1,295,623 | 1.4 | ✓ | n/a |
| ReVeal (Chakraborty et al., 2022) | C/C++ | Function | 18,169 | 9.2 | ✗ | n/a |
| CVEfixes (v1.0.8) (Bhandari et al., 2021) | 27 languages | Function, Commit, File | 277,948 | 45.5 | ✓ | 272 |
| CrossVul (Nikitopoulos et al., 2021b) | 40+ languages | File | 27,476 | 50.0 | ✓ | 168 |
| SecurityEval (v2.1) (Siddiq and Santos, 2022) | Python | Program | 121 | 100.0 | ✓ | 69 |
| DiverseVul (Chen et al., 2023) | C/C++ | Function | 330,492 | 5.7 | ✓ | 150 |
| SVEN (He and Vechev, 2023) | C/C++, Python | Program | 1,606 | 50.0 | ✓ | 9 |
| FormAI* (Tihanyi et al., 2023) | C | Program, Line | 246,549 | 80.2 | (✓) | (41) |
| ReposVul* (Wang et al., 2024a) | C/C++, Java, Python | Project, File, Function, Line | 232,465 | 0.7 | ✓ | 236 |
| PrimeVul (Ding et al., 2025) | C/C++ | Function | 235,768 | 2.9 | ✓ | 140 |
| PairVul* (Du et al., 2025a) | C/C++ | Function | 8,628 | 50.0 | ✓ | 95 |
| MegaVul* (2024/04) (Ni et al., 2024) | C/C++ | Function | 353,873 | 5.1 | ✓ | 176 |
| CleanVul* (Li et al., 2025k) | C/C++, C#, Java, JavaScript, Python | Function | 11,632 | 50.0 | ✗ | n/a |
5.1.5. Dataset Programming Language
Commonly used vulnerability datasets are heavily biased towards C/C++, as shown in Table 5. While this bias reflects the prevalence of these languages in technological infrastructure, it creates a significant imbalance, as vulnerability patterns are intrinsically linked to the programming language. Qian et al. (Qian et al., 2025) highlight that unmanaged languages such as C/C++ are prone to memory safety violations (e.g., buffer overflows), whereas managed languages such as JavaScript or Python frequently suffer from logic bugs. In a comparative analysis of interpreted languages (Java, JavaScript, Python, and Ruby), Ruohonen (Ruohonen, 2021) observes that while certain weaknesses such as cross-site scripting and input validation errors are ubiquitous across all ecosystems, language-specific trends emerge. For instance, cryptographic issues were found to be prevalent in JavaScript but significantly less common in Java or Python (Ruohonen, 2021). Consequently, the predominance of C/C++ risks over-optimizing models for memory safety violations while potentially neglecting the distinct vulnerability landscapes of modern web languages.
Fewer datasets cover multiple programming languages, i.e., SARD, CVEfixes, CrossVul, and CleanVul; however, they are not represented in the same share. Recent studies have begun constructing datasets for less covered languages, e.g., Python (Ozturk et al., 2023) and Rust (Luo et al., 2025), contributing to broader language diversity in vulnerability research.
5.1.6. Dataset Size and Share of Vulnerable Samples
The size of a dataset is strongly influenced by the dataset source and its labeling method. Smaller datasets typically involve manual curation, enabling high label quality but limited scalability. SecurityEval, for example, focuses on evaluating secure code generation in Python. It consists of 121 manually curated prompt-based samples across 75 CWE types, resulting in a dataset with 100% vulnerable samples. Similarly, SVEN focuses on evaluating secure code generation, offering improved label reliability with a 50% share of vulnerable samples, but is limited in scope due to its focus on a fixed subset of CWEs. In contrast, larger datasets rely on automated pipelines for scalability. DiverseVul (Chen et al., 2023) scales up the volume and variety of functions compared to earlier VFC-based datasets such as Devign, ReVeal (Chakraborty et al., 2022), Big-Vul, CrossVul, and CVEfixes, with a share of 5% vulnerable samples.
The dataset type and the labeling method used in dataset construction influence both the quality and realism, cf. Figure 7. Synthetic datasets, such as Juliet or FormAI, offer high label accuracy due to controlled injection of known vulnerability patterns but often lack the structural complexity and semantic diversity found in real-world code. In contrast, datasets derived from real-world sources provide greater realism, better reflecting practical vulnerability scenarios. However, they often suffer from label noise due to assumptions inherent in VFC-based labeling. These datasets also vary in the proportion of vulnerable samples. Highly imbalanced datasets, i.e., where vulnerabilities are underrepresented, reflect the natural rarity of vulnerabilities in real code. While realistic, such an imbalance poses challenges for training models, which may become biased towards predicting the majority class. To mitigate this challenge, preprocessing techniques are often required, e.g., resampling: oversampling the minority (vulnerable) class or undersampling the majority (non-vulnerable) class to achieve a more balanced class distribution. Figure 7 visualizes this trade-off between labeling accuracy and realism, providing a practical framework for selecting datasets for different research objectives.
5.1.7. Dataset Multi-Class Labeling
Vulnerability detection is often framed as a binary classification task, i.e., distinguishing between vulnerable and non-vulnerable code. For example, the ReVeal dataset provides security-related patches collected from bug tracking systems, offering paired vulnerable and fixed code samples labeled in binary form.
Many datasets go beyond binary classification by providing labels for specific vulnerability types, enabling multi-class and multi-label classification. CVEfixes is a comprehensive and automatically curated dataset collected from CVE entries in the NVD. It identifies linked open-source projects and extracts function-level vulnerable and fixed code samples from VFCs. The latest release at the time of writing (v1.0.8) includes all CVEs published up to July 23, 2024, covering 272 distinct CWEs. The PairVul dataset (Du et al., 2025a) takes a more controlled approach by focusing on pairs of vulnerable and non-vulnerable code with high lexical similarity. These pairs are extracted from Linux kernel CVEs, labeled using VFCs, and filtered to ensure that the associated fixes are not reverted or modified in subsequent commits. While the dataset offers high-quality and balanced samples, its scope is currently limited to selected Linux kernel vulnerabilities.
A common limitation across vulnerability datasets is the lack of coverage of recently disclosed vulnerabilities. Static datasets may quickly become outdated, reducing their relevance for real-time or forward-looking research. In addition, this temporal lag introduces a critical validity threat for LLMs in the form of data leakage. Since widely used static datasets are likely contained within the pre-training corpora of LLMs and new models are continuously emerging, high performance on these benchmarks may indicate memorization rather than genuine generalization. This challenge highlights the need for continuously maintained and extensible datasets, e.g., CVEfixes, which is supported by an automated pipeline, to ensure evaluation against novel, unseen vulnerabilities.
Datasets such as Big-Vul, CrossVul, CVEfixes, and ReposVul further offer rich metadata, including CWE descriptions and severity scores. Such metadata supports diverse generative tasks, for example, root-cause reasoning and context-aware vulnerability analysis, broadening their use beyond classification.
5.2. CWE Coverage and Diversity
To assess the real-world applicability and limitations of LLM-based vulnerability detection approaches, it is crucial to use datasets that capture a diverse range of vulnerability types. Such diversity enables more realistic evaluations, as it reflects the variety and complexity of vulnerabilities encountered in practical software development. As shown in Table 5, the number and variety of vulnerability types (i.e., number of CWE classes) represented vary considerably across the selected datasets. Some datasets focus narrowly on specific vulnerability types and CWE classes, while others aim for broader CWE diversity. To assess the diversity and composition of CWEs, we analyze selected real-world and mixed datasets providing multi-class labels (cf. Table 4) in more depth, focusing on their CWE coverage and distribution. Specifically, we analyze VulDeePecker, Draper, Big-Vul, CVEfixes, CrossVul, SecurityEval, DiverseVul, SVEN, ReposVul, PrimeVul, PairVul, and MegaVul. D2A and FormAI are excluded from this analysis. D2A is not considered, as it provides only static analyzer outputs without CWE annotations. Similarly, FormAI, while performing manual CWE mapping, does not include these mappings as part of the dataset but only error types provided by the formal verification tool.
| CWE-ID | Pillar Name | Description | #CWEs |
| CWE-284 | Improper Access Control | Weaknesses related to protection mechanisms such as authentication, authorization, and accountability | 166 |
| CWE-435 | Improper Interaction Between Multiple Correctly-Behaving Entities | Weaknesses that arise due to unexpected consequences when multiple entities interact, even though each entity behaves correctly in isolation | 16 |
| CWE-664 | Improper Control of a Resource Through its Lifetime | Resource management weaknesses such as improper initialization, reuse, or cleanup; includes cases where explicit instructions for resource creation, usage, or destruction are not properly followed | 367 |
| CWE-682 | Incorrect Calculation | Calculations that generate incorrect or unintended results later used in security-critical decisions or resource management | 14 |
| CWE-691 | Insufficient Control Flow Management | Logic or flow vulnerabilities such as incorrect conditions or decision-making | 84 |
| CWE-693 | Protection Mechanism Failure | Weaknesses caused by bypassing, misconfiguring, or incorrectly using security mechanisms | 100 |
| CWE-697 | Incorrect Comparison | Weaknesses in comparison logic between variables, objects, or values | 22 |
| CWE-703 | Improper Check or Handling of Exceptional Conditions | Weaknesses where exceptional conditions that rarely occur during normal operation are not properly anticipated or handled | 59 |
| CWE-707 | Improper Neutralization | Inadequate sanitation or escaping of input/output data where the application fails to ensure that data conforms to expected formats and is safe | 144 |
| CWE-710 | Improper Adherence to Coding Standards | Violations of safe and established programming practices | 195 |
As discussed in prior works (Deng et al., 2024; Jiang et al., 2025b; Ni et al., 2024), real-world vulnerability datasets exhibit a characteristic long-tail distribution: A small number of CWE types are heavily represented, while the majority of CWE types appear only rarely, often with just one or two examples. This characteristic is visualized in Figure 8, which shows the count of CWEs for the 15 most and least represented CWE classes for selected datasets. For example, while CVEfixes covers 272 distinct CWE types, the majority are sparsely represented. A similar pattern is observed for DiverseVul, which covers 150 CWEs. Such skewed distributions lead to biased learning, where models disproportionately focus on a small subset of vulnerability types. MegaVul and SecurityEval represent efforts to encourage a more balanced representation. SecurityEval, in particular, provides one to six manually curated examples for 69 CWEs. These efforts represent a step towards reducing long-tail effects and developing more generalizable vulnerability detection models.
For a more structured comparison of the CWE distributions across the datasets, we adopt the CWE-1000 Research View (MITRE, 2023a). This view provides a hierarchical grouping of CWEs intended for research and analysis. It groups CWEs into broad research concepts to help identify inter-dependencies, shared characteristics, and root causes of vulnerabilities. The CWE-1000 view spans multiple levels of abstraction: At the highest level are pillars and categories, which provide high-level groupings but are not intended for direct vulnerability mapping. Beneath these are the class, base, and variant levels, with variant representing the most specific level, i.e., language- or technology-specific weaknesses. By design, the CWE-1000 view organizes every weakness in the CWE catalog into one of ten pillars, cf. Table 6.
We consider the mapping of CWE-1000 to cluster the CWEs represented in the individual datasets. This clustering enables a structured comparison beyond CWE counts by highlighting which vulnerability types are emphasized or underrepresented, assessing the thematic diversity and coverage of vulnerability datasets. Notably, many datasets were built using the NVD, which also includes deprecated or discouraged CWE entries not mapped by CWE-1000, as well as general categories such as CWE-OTHER or CWE-NOINFO, used when insufficient information is available. To account for these cases, we introduce an additional group labeled not mapped, capturing all CWE entries that fall outside the scope of the CWE-1000 view. For each dataset, we compute the relative share of each CWE based on the vulnerable functions (or files, depending on the granularity provided). A special case is the CVEfixes dataset, where VFCs are associated with CVE identifiers, and each CVE may be linked to one or more CWEs. Some commits address multiple CVEs, and the associated CWE mappings may span different levels of specificity. To normalize the distributions and ensure a consistent comparison, we compute the share of each CWE based on the total number of CWE labels assigned to all vulnerable samples, rather than the number of vulnerable functions.
Figure 9 visualizes the distribution of the Top 25 CWE types covered by vulnerable samples across the selected datasets. Among the mixed datasets and those with a CWE-specific focus, VulDeePecker, Draper, and SVEN include only a small number of CWEs, i.e., two, four, and nine, respectively. SecurityEval features the widest variety of CWEs among mixed datasets, with a similar distribution across its covered types. The CWEs represented in the Top 25 across all datasets account for only around 20% of its total CWE labels. In comparison, the real-world datasets exhibit a more pronounced long-tail distribution: the Top 25 CWEs typically account for around 80% of all vulnerability labels. This highlights a strong skew towards a small set of frequently occurring vulnerability types.
The most widely represented CWEs across datasets are those associated with the pillar Improper Control of a Resource Through its Lifetime, which includes memory and resource management vulnerabilities that are commonly found in C/C++, the most dominant language across the studied datasets (cf. discussion in Section 5.1.5). This pillar is also the largest in terms of CWE count (cf. Table 6), followed by Improper Neutralization, encompassing issues related to input validation and sanitization. Other pillars are underrepresented in comparison to their relative size in the CWE view. These pillars include Improper Adherence to Coding Standards and Improper Access Control. Notably, some pillars, such as Improper Interaction Between Multiple Correctly-Behaving Entities, Protection Mechanism Failure, and Incorrect Comparison, are not represented at all among the Top 25 CWEs, and are less represented in the overall CWEs covered, cf. Figure 10. These categories are inherently more difficult to capture in real-world datasets, particularly at the function level, as they often involve complex inter-component dependencies, system-level interactions, or contextual analysis that is not easily observable from isolated code snippets.
A considerable share of CWEs in the analyzed datasets fall into the not mapped category. This is often the result of outdated or abstract vulnerability labels sourced from the NVD. For instance, VulDeePecker covers two high-level CWEs, i.e., the class CWE-119 Improper Restriction of Operations within the Bounds of a Memory Buffer (MITRE, 2025b) and the category CWE-399 Resource Management Errors (MITRE, 2023b), flattening the existing CWE hierarchy and obscuring more specific vulnerability types. Importantly, the use of abstract categories such as classes for mapping has been discouraged since 2019, as they are frequently misapplied in low-information vulnerability reports where more specific child CWEs would be more appropriate (MITRE, 2025b). The use of categories for vulnerability mapping is explicitly prohibited (MITRE, 2023b). Several other CWEs commonly found in the selected datasets are also discouraged for use in vulnerability mappings, e.g., CWE-20 Improper Input Validation, CWE-200 Exposure of Sensitive Information to an Unauthorized Actor, and CWE-400 Uncontrolled Resource Consumption. In addition, pillars such as CWE-284 Improper Access Control and CWE-703 Improper Check or Handling of Exceptional Conditions are often used as placeholders in vulnerability reports lacking detailed analysis, leading to imprecise dataset annotations.
Some CWEs are marked as allowed with review, meaning they can be used in vulnerability mapping but should be applied carefully and only when more specific alternatives are not available. For example, CWE-120 Buffer Copy without Checking Size of Input is frequently selected based on keyword presence (’Classic Buffer Overflow’), but is only appropriate when the vulnerability involves unchecked buffer copy operations (MITRE, 2025c). Similar caution applies to CWEs such as CWE-362 Concurrent Execution using Shared Resource with Improper Synchronization (’Race Condition’) and CWE-94 Improper Control of Generation of Code (’Code Injection’).
The use of general or deprecated CWE labels presents challenges to dataset quality and limits the effectiveness of CWE-based classification and reasoning. Notably, there are 555 CWEs in the CWE-1000 view that are allowed for vulnerability mapping but are not represented as vulnerable samples in any of the analyzed datasets. This indicates that current datasets cover only a subset of CWE types, leaving many mappable software weaknesses absent from training and benchmarking. As a result, tools trained on these datasets may lack generalization capabilities when faced with previously unseen vulnerabilities. Further, evaluation results appear artificially strong, as they are often biased towards commonly occurring CWEs while excluding rare and long-tail vulnerability types. To address these limitations, more effort is needed to incorporate the CWE hierarchy into training and evaluation protocols, moving beyond binary labels to consider hierarchically related CWEs. For example, Tamberg and Bahsi (Tamberg and Bahsi, 2025) follow an evaluation method in which both parent and child CWEs are considered valid classifications using MITRE’s Research Concepts view.
5.3. Trends and Use Cases in Dataset Utilization
Understanding how vulnerability datasets are used over time provides insight into community practices, widely used benchmark datasets, and emerging research directions. Analyzing temporal trends also reveals how quickly new datasets are adopted, how long benchmark datasets remain influential, and where gaps or biases may persist in evaluation setups. Such an analysis is particularly important in this rapidly evolving field to guide future dataset development and advance performance through more representative and robust evaluations.
Figure 11 visualizes the temporal usage and usage frequency of the selected vulnerability datasets from Jan-2020 to Nov-2025 by the surveyed studies. Red markers indicate the publication dates of the datasets. The Figure shows that even today, early deep learning datasets such as Devign, which serves as a foundational dataset for graph-based and learning-based vulnerability detection models, and Big-Vul remain in widespread use. Classic test suites such as SARD and Juliet are also frequently used, cf. Figure 11, useful for controlled evaluations due to their synthetic nature. However, these earlier datasets often lack current CWE coverage, may reinforce outdated vulnerability mappings, and may already be part of training corpora.
Figure 11 also shows that since 2021, there has been a clear trend towards real-world vulnerability datasets, e.g., CVEfixes, CrossVul, and DiverseVul. Several more recent datasets, such as FormAI, PrimeVul, PairVul, and CleanVul, are explicitly designed for evaluating LLM-based vulnerability detection approaches. For these datasets, adoption lags of approximately one to six months can be observed between a dataset’s release and adoption in published research. This lag reflects the time needed for several practical steps, including dataset validation, integration into existing toolchains and model pipelines, and broader community uptake. In many cases, adoption may also depend on the availability of documentation and preprocessing scripts. Further, early adopters often play a key role in demonstrating a dataset’s utility, which can influence its acceptance and reuse.
Another notable trend is the increasing use of newly created or custom datasets, see row ”Custom” in Figure 11. A total of 87 studies either use or introduce custom datasets, often to address perceived limitations in existing datasets. Common motivations include correcting unrealistic data distributions, expanding project and vulnerability diversity, including more recent disclosures, or improving overall dataset quality (e.g., fixing incomplete or incorrectly merged functions) (Ni et al., 2024). This trend is also driven by domain-specific or task-aligned needs. While such efforts encourage innovation and enable tailored evaluations, they also complicate comparability and may introduce implicit biases tied to individual evaluation design choices and construction assumptions.
The choice of suitable vulnerability datasets reflects a combination of common benchmark datasets, task alignment, and system design. However, the use of various datasets across the surveyed studies illustrates a high degree of fragmentation. Most studies either use or introduce custom datasets. Only nine datasets are reused in more than ten studies, highlighting the lack of a standardized suite of benchmark datasets and consistency in dataset selection. Many studies use more than one dataset, often combining datasets to increase training data diversity or to evaluate model generalization, e.g., vulnerability types and programming languages. Figure 12 highlights the most frequently used dataset pairs across all surveyed studies. Among these, the most common combinations are Devign and ReVeal, Devign and Big-Vul, as well as Big-Vul and ReVeal. These pairings often complement each other in terms of data characteristics, such as vulnerable percentage or combining datasets with broad versus narrow CWE coverage.
The use of many different datasets and combinations, however, hampers direct comparability of evaluation results, particularly when studies apply different splits for training, validation, and testing, or pre-process the same datasets differently. Figure 13 offers insight into overlaps and divergences in dataset usage across specific use cases. When fine-tuning LLMs for vulnerability detection, e.g., using LoRA (Low-Rank Decomposition, 27 studies), researchers often resort to large and diverse datasets, most commonly DiverseVul, Big-Vul, PrimeVul, and Devign. However, many studies make use of additional, lesser-used datasets or introduce custom datasets (ten studies). This practice reduces comparability and makes it difficult to replicate results or isolate model performance. RAG studies (25 studies) most commonly use Big-Vul and Devign, appearing in six and five studies, respectively. Big-Vul, in particular, is often used for knowledge base construction of CWE entries due to the provided metadata, e.g., CWE descriptions. Eight RAG studies introduce custom datasets, which hinder comparison, especially concerning the semantics of the retrieved context. Agentic approaches (10 studies) show even greater fragmentation, with only four studies aligning on PrimeVul. GNN-based hybrid approaches (20 studies) most frequently rely on Devign, followed by custom datasets, Big-Vul, and ReVeal. Devign is particularly suited for GNNs due to its composite code representation, encoding program control and data dependencies as heterogeneous edges within a unified graph structure. The absence of shared benchmarks within the discussed use cases highlights the need for controlled and standardized evaluation protocols.
6. Current Limitations and the Road Ahead
Despite progress in leveraging LLMs for software vulnerability detection, we identified several remaining limitations in this SLR. In the following, we outline the key limitations derived from the surveyed studies and propose actionable research directions to address them. Table 7 summarizes the discussed limitations and future research opportunities.
6.1. Limitations
The limitations identified in this review emerge from recurring issues observed across the surveyed studies and span key aspects such as dataset realism and representativeness, vulnerability diversity, methodological rigor, and evaluation practices. We structure these findings into seven themes L1-L7 that outline the most pressing limitations.
| Limitations | Future Research Opportunities |
| L1 Dataset and Detection Granularity | |
| • Focus on function-level inputs | • Explore coarser detection granularity (project- and repo-level) |
| • Failure to capture complex vulnerabilities with dependencies | • Augment code context |
| • Focus on C/C++ code | • Expand to other languages |
| • Develop multilingual and cross-project benchmarks | |
| L2 Dataset Labeling and Quality | |
| • Synthetic datasets oversimplify real-world complexity | • Build new datasets focusing on vulnerability diversity |
| • Automated labeling introduces noise and redundancy | • Improve label quality (manual verification, improved VFC tools) |
| • Missing reasoning ground truth | • Reasoning-oriented datasets and benchmarks |
| • Class imbalance and long-tail distribution of CWE types | • Reevaluate approaches on diverse and representative datasets |
| • Data leakage | • Build new datasets considering knowledge cutoff |
| L3 Study Evaluation and Comparability | |
| • Inconsistent use of datasets | • Align dataset use across similar use cases for comparability |
| • Lack of predefined train/validate/test splits | • Use datasets with documented splits |
| • Diverse modifications to datasets | • Evaluate on multiple datasets for comparability |
| • Document all preprocessing and filtering steps | |
| • Encourage open science | |
| L4 Up-to-Date Vulnerability Knowledge | |
| • Static datasets quickly become outdated | • Use automatically updated datasets (e.g., CVEfixes) |
| • Retraining is expensive and often infeasible | • Further explore RAG and continual learning |
| L5 Code and Vulnerability Representation | |
| • Raw code often used without structure or abstraction | • Investigate structure-aware representations |
| • Reliance on superficial text patterns | • Combine LLMs with structural models (e.g., GNNs) |
| • Sensitivity to semantic-preserving transformations | • Reevaluate robustness on perturbed or transformed code |
| L6 Model Interpretability and Explainability | |
| • Limited insight into predictions | • Develop metrics for explanation trustworthiness |
| • Hallucinated or misleading justifications, self-contradictions | • Use structured explanation formats |
| • Leverage external knowledge to ground reasoning | |
| L7 Integration into Pipelines and Workflows | |
| • Few evaluations in realistic development settings | • Integrate tools into IDEs or pipelines with developer feedback |
| • Poor generalization to production codebases | • Evaluate models on proprietary or closed-source datasets |
| • High computational cost of tuning and inference | • Develop CWE-type classifiers for finer-grained prediction |
| • Explore agentic systems for reasoning or self-assessment | |
| • Explore model compression for lightweight deployment | |
6.1.1. L1 Dataset and Detection Granularity
In the surveyed studies, vulnerability detection is typically applied at the level of individual functions or small program slices. This focus on function-level granularity has been influenced by the dominance of function-level vulnerability datasets (cf. Table 4) and partly by architectural limitations, particularly the restricted input lengths of earlier LLMs (Steenhoek et al., 2025a). However, methods have been proposed to enhance context, such as the work by Chen et al. (Chen et al., 2024), and LLMs now support significantly longer input contexts, e.g., 128,000 tokens for GPT4 (OpenAI et al., 2024). Despite these capabilities, the continued focus on function-level datasets fails to capture the context necessary to generalize to complex vulnerabilities that span multiple functions or files. For example, Huynh et al. (Huynh et al., 2025) investigate how varying levels of code context influence detection performance, comparing snippets enriched with comments and docstrings against full-file inputs. The findings suggest that even with extended context windows, current models still perform poorly on distributed vulnerabilities.
Further, the surveyed studies focus predominantly on detecting vulnerabilities in C/C++code (cf. Table 5), with limited coverage of other widely used programming languages, such as Java or Python. This narrow programming language coverage limits the applicability of current models in diverse, multilingual software environments.
6.1.2. L2 Dataset Labeling and Quality
Synthetic datasets offer high label-precision, as vulnerabilities are injected and, thus, well-defined. However, synthetic datasets may oversimplify real-world scenarios and fail to capture the complexity and diversity of real-world vulnerabilities, limiting their effectiveness in preparing models for deployment in realistic environments. Similarly, the common formulation as a binary classification task (cf. Figure 4) lacks the granularity needed for practical application, where identifying the specific vulnerability type is essential for severity assessment, prioritization, and appropriate repair. Real-world datasets are typically labeled using automated methods such as VFCs or static analysis tools. While these approaches enable large-scale dataset creation, they are associated with label noise and redundancy, which reduce the uniqueness and representativeness of the dataset (Croft et al., 2023). Although manual labeling or verification can improve precision, it remains a resource-intensive process, especially when scaling to larger datasets.
Moreover, as research shifts toward generative tasks, a critical gap of missing reasoning ground truth emerges. While existing datasets typically provide the label and possibly a description of the vulnerability, they rarely provide a verified chain of thought reasoning the root cause. This limitation makes it difficult to quantitatively benchmark the quality of LLM-generated reasoning without expensive human evaluation.
Further, most datasets exhibit a strong class imbalance, where vulnerable samples are heavily outnumbered by non-vulnerable ones (cf. Table 5), and a long-tail distribution of vulnerability types, with few CWEs dominating and many others being severely underrepresented (cf. Figure 8 and Figure 9). These imbalances pose significant challenges for both training balance and evaluation reliability, causing models to overfit to frequent vulnerabilities while failing on rare ones (Huynh et al., 2025).
Beyond these traditional labeling and quality issues, LLMs introduce the unique challenge of data leakage. Unlike models trained on closed datasets, LLMs are pre-trained on vast open-source corpora. Consequently, popular benchmarks (e.g., Devign, Big-Vul) are likely contained within the pre-training corpus, rendering standard train/test splits unreliable. This challenge is compounded by the knowledge cut-off; without continuously updated datasets that post-date the model’s training window, it is difficult to evaluate whether an LLM is detecting a vulnerability via reasoning or simply recalling historical data it has already seen.
6.1.3. L3 Study Evaluation and Comparability
The lack of consistency and transparency in dataset use and evaluation protocols complicates meaningful comparison across studies. Even within the same use case, studies are often not comparable due to the use of different datasets (cf. Figure 13). Further, most datasets do not provide predefined train/validate/test splits, which are essential for ensuring a consistent and fair evaluation (Shereen et al., 2024). Only a few studied datasets offer standardized splits, e.g., Draper (Russell et al., 2018), CodeXGLUE (Lu et al., 2021) (a subset of Devign (Zhou et al., 2019)), Big-Vul (Fan et al., 2020), and PrimeVul (Ding et al., 2024a). Without fixed splits, results can vary significantly, even when using the same dataset. For example, if subsequent studies fine-tune on the same dataset but apply different splits, direct performance comparisons become unreliable. In addition, several studies further modify or refine existing datasets (Zhang et al., 2024c; Yin et al., 2024; Hort et al., 2025; Gonçalves et al., 2025; Wang et al., 2025c), or augment them with custom data (Qi et al., 2024). While such adaptations are often motivated by practical needs, e.g., improving data quality or evaluating generalization, they introduce additional variability and complicate comparability with prior and future works.
6.1.4. L4 Up-to-Date Vulnerability Knowledge
The number of disclosed software vulnerabilities continues to grow rapidly (CVE, 2025a). As a result, static datasets may quickly become outdated. Models trained on such outdated datasets may lack awareness of newly discovered vulnerabilities and emerging vulnerability patterns, limiting their practical relevance and effectiveness in real-time scenarios and for the detection of zero-day vulnerabilities (Nong et al., 2024). Another challenge lies in how to efficiently expose models to this growing volume of vulnerability information. Since retraining models from scratch is computationally expensive and impractical at high frequency, there is a critical need for scalable methods to update or augment model knowledge.
6.1.5. L5 Code and Vulnerability Representation
In most surveyed studies, the code to be analyzed is provided as raw code or embedded directly within the prompt during inference. However, models frequently rely on superficial textual patterns rather than capturing the underlying semantic structures of vulnerabilities (Ding et al., 2025; Zhao et al., 2023; Li et al., 2025a). As a result, model performance degrades under semantic-preserving code transformations, such as variable renaming, code reordering, or formatting changes, highlighting a lack of robustness and true semantic understanding (Ni et al., 2025; Risse and B”̈ohme, 2024; Qi et al., 2024; Hort et al., 2025; Panichella, 2025; Wang et al., 2025b). While contrastive learning is designed to produce more discriminative embeddings, models still often fail to draw the classification boundary correctly and identify the vulnerable patterns (Ding et al., 2025).
6.1.6. L6 Model Interpretability and Explainability
Current LLM-based vulnerability detection approaches offer limited insight into why a model classifies the input code as vulnerable or non-vulnerable. Even in studies that use CoT prompting or explicitly investigate reasoning, issues with hallucinated or misleading justifications are commonly reported (Espinha Gasiba et al., 2024; Noever, 2023). Similarly, studies implementing verification report instances of inaccurate corrections, self-contradictions, and hallucinated justifications (Yu et al., 2024). These issues undermine the trustworthiness of model outputs and hinder their integration in practical development workflows where interpretability and reliability are critical.
6.1.7. L7 Integration into Pipelines and Workflows
Despite growing interest, the integration of LLM-based vulnerability detection into practical development workflows remains underexplored. Only a few studies investigate real-world applicability through prototype tools or user studies (e.g., (Fu et al., 2024; Steenhoek et al., 2025b)), though such efforts are essential for assessing feasibility and trust in practical settings. Generalization to realistic codebases also remains challenging, with models often underperforming outside curated benchmarks (Chakraborty et al., 2024).
Resource efficiency is a further limiting factor for real-world deployment, as fine-tuning and inference with LLMs remain computationally expensive. While model compression techniques such as knowledge distillation reduce inference latency and memory usage (d’Aloisio et al., 2025; Shi et al., 2024; Ibanez-Lissen et al., 2025), their practical application for deployment on developer machines remains unexplored. This gap is particularly relevant when considering the privacy constraints of proprietary codebases, which may require on-premise or offline analysis (Bappy et al., 2025).
6.2. Actionable Insights and Future Research Opportunities
Building on the identified limitations, each theme reflects core challenges in current research and highlights open questions that future work must address to enable more robust, generalizable, and reproducible progress in LLM-based software vulnerability detection. In this section, we provide interpretations or promising research directions, offering actionable insights to guide future work towards addressing these challenges.
6.2.1. L1 Dataset and Detection Granularity
To move beyond the limitations of function-level detection, future work should focus on context-enhanced vulnerability detection, including project- and repository-level analysis (Zhou et al., 2024a; Wen et al., 2024b; Li et al., 2025f; Guo et al., 2025). Initial steps in this direction include augmenting function-level samples with additional context, such as surrounding code lines (Halder et al., 2025), function arguments, external functions, type definitions, global variables, environmental constraints (Risse et al., 2025), as well as dependency and execution flows (Zheng et al., 2025). For example, Ahmed et al. (Ahmed et al., 2025) use GPT-4 to identify the required contexts for a given vulnerability in a function, but still find limitations in pinpointing vulnerable statements and their root causes in complex real-world code. Li et al. (Li et al., 2025i) demonstrate that incorporating execution and data context improves model performance and reasoning quality, emphasizing that the challenge lies in precise, context-aware vulnerability reasoning. Similarly, Yang et al. (Yang et al., 2025b) propose a program analysis-based approach that abstracts complex function calls into primitive API representations to enrich contextual understanding. Further investigating such context-enhancements is essential for capturing dependencies that span multiple functions or files, enabling the detection of more complex vulnerability types that are otherwise missed in isolated, function-level analysis.
The current language focus on C/C++ should be expanded to include other widely used languages such as Python, Java, and Rust (Carbonnelle, 2025). Establishing multilingual and cross-project benchmarks will be essential for assessing robustness and model generalization.
6.2.2. L2 Dataset Labeling and Quality
To address current limitations in dataset construction, future work should focus on building real-world vulnerability datasets with improved label quality and balanced coverage across CWE types. Label quality can be improved through two main strategies: (1) refining existing datasets via targeted manual labeling to correct and verify automatically generated labels (He and Vechev, 2023), and (2) advancing automated labeling methods, such as VFCs, to reduce noise and improve labeling reliability (Ding et al., 2024a; Li et al., 2025k; Ahmed et al., 2025; Gao et al., 2025; Lu et al., 2025). Balanced coverage across frequent and less common vulnerability types can be achieved, e.g., by using LLMs to generate vulnerable samples for specific vulnerability types (Yong et al., 2025; Bappy et al., 2025; Shahzad et al., 2025).
Complementary to improved context, there is a critical need for further reasoning-oriented benchmarks (Sun et al., 2025b) to close the gap in ground truth. Future dataset creation should prioritize pairing vulnerable code with verified, step-by-step explanations of the root cause and exploitation logic. Establishing such reasoning ground truths is a prerequisite for advancing from simple pattern matching to genuine semantic understanding.
In addition, new datasets should be constructed with minimal overlap to existing corpora to prevent data leakage and inflated performance due to prior exposure during model training. These datasets should set a focus on programming language and vulnerability diversity, capturing variations in human-written and LLM-generated code, complex multi-line vulnerabilities, and less frequently represented CWE types (Steenhoek et al., 2023a). A practical direction for dataset expansion is to use the set of mappable-but-unused CWEs identified in Section 5.2, i.e., CWE classes that are currently not represented in the discussed datasets. Revisiting existing vulnerability detection approaches on more representative datasets may yield new insights into their robustness and effectiveness in real-world vulnerability detection.
6.2.3. L3 Study Evaluation and Comparability
To enhance reproducibility and comparability across studies, future research should prioritize standardized evaluation protocols. Specifically, we recommend the use of datasets with documented train/validation/test splits (or even closed evaluation of test splits) and careful alignment with datasets used in related studies to ensure meaningful comparisons within the same use case. We provide the mappings of datasets to surveyed use cases in the artifacts (Kaniewski et al., 2025).
To assess generalization and maximize comparability with prior works, models should be evaluated on multiple datasets, combining recent datasets such as PrimeVul (Ding et al., 2024a) or MegaVul (Ni et al., 2024) with established benchmark datasets such as Devign (Zhou et al., 2019) or Big-Vul (Fan et al., 2020). All pre-processing and filtering steps must be documented to enable replication. In support of open science (Nong et al., 2023), sharing of code, evaluation scripts, and model checkpoints is encouraged.
6.2.4. L4 Up-to-Date Vulnerability Knowledge
Given the broad and continuously growing number of CVEs, ensuring that models remain informed about newly discovered vulnerabilities is crucial for maintaining relevance in practical scenarios. To address the limitation of static datasets quickly becoming outdated, future work should focus on continuously maintained and automatically extensible datasets, such as CVEfixes (Bhandari et al., 2021), which employs an automated pipeline for updates. Advanced adaptation techniques also offer promising pathways to bridge this knowledge gap. RAG enables LLMs to access external, up-to-date knowledge at inference time. Continual learning can further support adaptability by integrating new vulnerability knowledge while preserving performance on previously seen vulnerability types. However, both techniques have seen limited adoption in the context of software vulnerability detection. Further research should explore their effectiveness, especially in handling large and heterogeneous vulnerability information. For RAG, this includes investigating how different representations, e.g., structured CWE hierarchy graphs or abstract function descriptions, affect retrieval quality and downstream detection accuracy.
6.2.5. L5 Code and Vulnerability Representation
To address the need for robust representations that capture vulnerability semantics, future work should further focus on structure-aware input representations. Directions include graph-based representations, such as control and data flow graphs, (LLM-generated) abstract descriptions that capture semantics and code behavior (Zhang et al., 2025a; Wu et al., 2025; Du et al., 2025a), and multi-modal inputs, e.g., combining API abstractions, data flow graphs, and natural language documentation (Yang et al., 2025b). Understanding how different representations complement each other can enhance vulnerability detection capabilities. Ideally, such representations should be language-agnostic to support generalization across diverse programming languages (cf. multi-language studies (Dozono et al., 2024; Zhang et al., 2025b)). In addition, hybrid architectures should be further explored. GNNs, in particular, have demonstrated potential for capturing code semantics and structural dependencies (Chang et al., 2024; Islam et al., 2023, 2024a; Jiang et al., 2024b; JianJie and Le, 2023; Kong et al., 2024; Liu et al., 2025b; Ni et al., 2023a; Yang et al., 2024a; Zheng et al., 2024). Future work should also evaluate the robustness of different representation strategies against syntactic perturbations and domain shifts, such as LLM-generated or refactored code, to better assess real-world generalization capabilities.
6.2.6. L6 Model Interpretability and Explainability
To build trust in LLM-based vulnerability detection systems, future research should go beyond standard evaluation metrics (e.g., accuracy, recall, precision, F1) and incorporate dedicated metrics that assess the trustworthiness and consistency of model-generated explanations. As LLMs increasingly take on roles in security workflows, their predictions must not only be correct but also justifiable and transparent to developers. Justification includes distinguishing between correct predictions made for the wrong reasons (spurious correlations) and those grounded in valid vulnerability semantics (Risse et al., 2025). In this matter, studies should increasingly look into Explainable AI (XAI) techniques (Takieldeen et al., 2025; Nguyen et al., 2025a; Alam et al., 2025).
Interpretability may further be enhanced through structured explanation formats, such as vulnerability propagation paths or flow annotations. These structured rationales offer greater reliability than natural language justifications and can be cross-checked against program logic or expert knowledge. Further, integrating RAG techniques with structured vulnerability knowledge, such as CWE hierarchies or exploit chains, could improve reasoning by anchoring the model’s output in verifiable context. This approach may not only strengthen the grounding of explanations but also align with the information needs of developers and auditors in practice.
6.2.7. L7 Integration into Pipelines and Workflows
To ensure the practical applicability of LLM-based vulnerability detection, future work should focus on the integration into real-world development workflows. Such works should, e.g., integrate solutions into IDEs and CI/CD pipelines, conduct field studies, and collect developer feedback. Closed-source or proprietary codebases can serve as valuable testbeds for evaluating real-world performance, robustness, and usability.
Agentic systems and reasoning-driven workflows offer promising pathways to autonomously assist developers (e.g., expert-in-the-loop system (Farr et al., 2025)) through tasks such as self-assessment, explanation generation, or iterative refinement.
To further improve generalization to practical vulnerability remediation workflows, LLM-based approaches should move towards CWE-type classification and report generation. Promising directions include the development of type-specific classifiers that align with the CWE hierarchy and research views, cf. (Atiiq et al., 2024b; Cao et al., 2024; Deng et al., 2024), enabling more actionable insights and effective prioritization.
A current pathway for industrial adoption lies in synergizing LLMs with traditional program analysis (Li et al., 2024). While LLMs excel at semantic reasoning, they lack the formal correctness guarantees of, e.g., formal verification tools. Pipelines where SAST tools detect candidates and LLMs filter false positives can significantly reduce alert fatigue by applying semantic understanding to noisy static reports. Conversely, workflows where LLMs propose vulnerabilities and static analyzers verify feasibility offer a mechanism to ground probabilistic outputs.
Finally, trade-offs between performance and resource efficiency must be considered for real-world deployment. Fine-tuning smaller models for on-premise use is both practical and privacy-preserving (Bappy et al., 2025). Alternatively, model compression techniques such as knowledge distillation and quantization offer promising pathways for enabling lightweight deployment on developer machines.
7. Threats to Validity
To ensure transparency of our findings, we acknowledge potential threats to the validity of this SLR along with the mitigation strategies employed.
Paper retrieval omissions. A primary threat to any SLR is the omission of relevant studies due to studies with incomplete titles, abstracts, or varied keyword terminologies. To address this threat, we constructed a broad search string based on manually selected keywords from primary studies as well as keywords from prior SLRs. We combined manual search, automated search, and snowballing to minimize overlooking relevant papers. Despite these efforts, due to the living nature of the field, a static snapshot may miss very recent developments. To address this threat, we employ a continuous search strategy using alerts and maintain a repository to update the study corpus.
Study selection bias. The selection of studies involves subjective judgment, which introduces potential bias. To minimize this threat, we conducted a revision of all crawled papers, strictly applying predefined inclusion and exclusion criteria (cf. Table 2). A specific threat regarding quality arises from our decision not to filter exclusively for high-ranked conferences or journals. We included a wide range of studies to provide a holistic picture of the landscape and identify gaps that might be missing in a narrower scope. To address the quality implications of this decision, we provide an analysis of the distribution of studies across conference and journal ranks as a quality indicator. Additionally, we include preprints from arXiv. While necessary to reflect the latest advancements, preprints have not undergone rigorous peer review. To mitigate this threat, we conducted a manual quality assessment, excluding studies that lacked clear evaluation setups or coherent contributions.
Data extraction and classification. The manual extraction of information and the construction of the taxonomy are subject to researcher bias and interpretation errors. For instance, determining whether a study performs ”reasoning” versus ”description” can be subjective. To mitigate this risk, we aligned our taxonomy with existing definitions where possible (e.g., dataset granularity, PEFT techniques) and refined our categories iteratively during the full-text review. We further support the reliability of our extraction by publicly releasing the full list of categorized studies and their attributes in our repository (Kaniewski et al., 2025).
8. Conclusion
In this SLR, we analyzed 263 studies on LLM-based software vulnerability detection. To structure this rapidly evolving field, we introduced a comprehensive taxonomy that covers detection task formulation, i.e., the classification and generation tasks presented to the LLM; input representation, i.e., how code and auxiliary information are provided to the LLM; system architecture, differentiating between LLM-centric and hybrid approaches; as well as model adaptation (prompt engineering, fine-tuning, and learning paradigms) and orchestration techniques, e.g., agentic and ensemble systems. The analysis shows that most studies perform binary code classification (vulnerable/non-vulnerable). Common adaptation techniques include full-parameter fine-tuning and zero-shot prompting. More advanced adaptation techniques, learning paradigms, and orchestration designs, such as parameter-efficient fine-tuning, retrieval-augmented generation, continual learning, and agentic workflows, are only beginning to emerge. We further analyzed common datasets used, investigating their realistic nature, diversity in vulnerability types with respect to CWE-1000 pillars, and usage trends. Despite notable progress in the field, we identified several key limitations that hinder practical adoption of LLMs in software vulnerability detection, including limited robustness in code representations, vulnerability data leakage, and limited interpretability and explainability. To address the identified limitations, we outlined actionable research directions, such as advancing structure-aware and language-agnostic input representations, aligning datasets across use cases to enable cross-study comparison and standardized evaluation protocols, adopting retrieval and continual learning techniques to improve adaptability, and fitting approaches into practical development workflow requirements. By mapping existing studies, identifying open challenges, and proposing future research directions, this review aims to guide researchers and practitioners in advancing the development of reliable, generalizable, and practically applicable LLM-based vulnerability detection systems.
Acknowledgements.
This work has been supported in part by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) Project-ID 528745080 - FIP 68. The authors alone are responsible for the content of the paper.References
- (1)
- Abdin et al. (2024) Marah Abdin, Jyoti Aneja, Hany Awadalla, Ahmed Awadallah, Ammar Ahmad Awan, Nguyen Bach, Amit Bahree, Arash Bakhtiari, Jianmin Bao, Harkirat Behl, et al. 2024. Phi-3 Technical Report: A Highly Capable Language Model Locally on Your Phone. arXiv:2404.14219
- Ahmad et al. (2021) WU Ahmad, S Chakraborty, B Ray, and KW Chang. 2021. Unified Pre-training for Program Understanding and Generation. In Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies (NAACL HLT). ACL, Mexico City, Mexico, 2655–2668. doi:10.18653/v1/2021.naacl-main.211
- Ahmed et al. (2025) Md Basim Uddin Ahmed, Nima Shiri Harzevili, Jiho Shin, Hung Viet Pham, and Song Wang. 2025. SecVulEval: Benchmarking LLMs for Real-World C/C++ Vulnerability Detection. arXiv:2505.19828
- Akli et al. (2025) Amal Akli, Maxime Cordy, Mike Papadakis, and Yves Le Traon. 2025. AutoAdapt: On the Application of AutoML for Parameter-Efficient Fine-Tuning of Pre-Trained Code Models. ACM Transactions on Software Engineering and Methodology (TOSEM) (2025). doi:10.1145/3734867 Just Accepted.
- Alam et al. (2025) Syeda Sadia Alam, Mst Shapna Akter, and Alfredo Cuzzocrea. 2025. Improving Software Security Through a LLM-Based Vulnerability Detection Model. In International Conference on Database and Expert Systems Applications (DEXA). Springer, Bangkok, Thailand, 122–129. doi:10.1007/978-3-032-02049-9_9
- Allamanis et al. (2018) Miltiadis Allamanis, Earl T. Barr, Premkumar Devanbu, and Charles Sutton. 2018. A Survey of Machine Learning for Big Code and Naturalness. Comput. Surveys 51, 4, Article 81 (2018), 37 pages. doi:10.1145/3212695
- Alqarni and Azim (2025) Mansour Alqarni and Akramul Azim. 2025. An Advanced Detection Framework for Embedded System Vulnerabilities. IEEE Access 13 (2025), 159207–159216. doi:10.1109/ACCESS.2025.3607595
- Alrashedy et al. (2024) Kamel Alrashedy, Vincent J. Hellendoorn, and Alessandro Orso. 2024. Learning Defect Prediction from Unrealistic Data. In International Conference on Software Analysis, Evolution and Reengineering (SANER). IEEE, Rovaniemi, Finland, 556–567.
- Ambati et al. (2024) Sri Haritha Ambati, Norah Ridley, Enrico Branca, and Natalia Stakhanova. 2024. Navigating (In)Security of AI-Generated Code. In International Conference on Cyber Security and Resilience (CSR). IEEE, London, UK, 1–8. doi:10.1109/CSR61664.2024.10679468
- Amodei (2025) Dario Amodei. 2025. CEO Speaker Series With Dario Amodei of Anthropic. https://www.cfr.org/event/ceo-speaker-series-dario-amodei-anthropic. Interview by Michael Froman.
- Anthropic (2024) AI Anthropic. 2024. The Claude 3 Model Family: Opus, Sonnet, Haiku. Claude-3 Model Card 1 (2024), 1.
- Aslan et al. (2025) Mert Kaan Aslan, Yunus Emre Alkan, Muhammed Burak Alican, and Özgür Özdemir. 2025. Utilizing Large Programming Language Models on Software Vulnerability Detection. In Innovations in Intelligent Systems and Applications Conference (ASYU). IEEE, Bursa, Turkey, 1–6. doi:10.1109/ASYU67174.2025.11208282
- Atiiq et al. (2024a) Syafiq Al Atiiq, Christian Gehrmann, and Kevin Dahlén. 2024a. Vulnerability Detection in Popular Programming Languages with Language Models. arXiv:2412.15905
- Atiiq et al. (2024b) Syafiq Al Atiiq, Christian Gehrmann, Kevin Dahlén, and Karim Khalil. 2024b. From Generalist to Specialist: Exploring CWE-Specific Vulnerability Detection. arXiv:2408.02329
- Bae et al. (2024) Jaehyeon Bae, Seoryeong Kwon, and Seunghwan Myeong. 2024. Enhancing Software Code Vulnerability Detection Using GPT-4o and Claude-3.5 Sonnet: A Study on Prompt Engineering Techniques. Electronics 13, 13 (2024), 2657. doi:10.3390/electronics13132657
- Bahaa et al. (2024) Ahmed Bahaa, Aya El-Rahman Kamal, Hanan Fahmy, and Amr S. Ghoneim. 2024. DB-CBIL: A DistilBert-Based Transformer Hybrid Model Using CNN and BiLSTM for Software Vulnerability Detection. IEEE Access 12 (2024), 64446–64460. doi:10.1109/ACCESS.2024.3396410
- Bai et al. (2023) Jinze Bai, Shuai Bai, Yunfei Chu, Zeyu Cui, Kai Dang, Xiaodong Deng, Yang Fan, Wenbin Ge, Yu Han, Fei Huang, et al. 2023. Qwen Technical Report. arXiv:2309.16609
- Bakhshandeh et al. (2023) Atieh Bakhshandeh, Abdalsamad Keramatfar, Amir Norouzi, and Mohammad Mahdi Chekidehkhoun. 2023. Using ChatGPT as a Static Application Security Testing Tool. arXiv:2308.14434
- Bappy et al. (2025) Md. Azizul Hakim Bappy, Hossen A Mustafa, Prottoy Saha, and Rajinus Salehat. 2025. Case Study: Fine-tuning Small Language Models for Accurate and Private CWE Detection in Python Code. arXiv:2504.16584
- Basic and Giaretta (2025) Enna Basic and Alberto Giaretta. 2025. From Vulnerabilities to Remediation: A Systematic Literature Review of LLMs in Code Security. arXiv:2412.15004
- Bhandari et al. (2021) Guru Bhandari, Amara Naseer, and Leon Moonen. 2021. CVEfixes: Automated Collection of Vulnerabilities and Their Fixes from Open-Source Software. In International Conference on Predictive Models and Data Analytics in Software Engineering (PROMISE). ACM, Athens Greece, 30–39. doi:10.1145/3475960.3475985
- Brown et al. (2020) Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared D Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, et al. 2020. Language Models are Few-shot Learners. In Advances in Neural Information Processing Systems (NeurIPS), Vol. 33. Curran Associates, Inc., Virtual-only, 1877–1901.
- Buratti et al. (2020) Luca Buratti, Saurabh Pujar, Mihaela Bornea, Scott McCarley, Yunhui Zheng, Gaetano Rossiello, Alessandro Morari, Jim Laredo, Veronika Thost, Yufan Zhuang, and Giacomo Domeniconi. 2020. Exploring Software Naturalness through Neural Language Models. arXiv:2006.12641
- Cao and Jun (2024) Daipeng Cao and W. Jun. 2024. LLM-CloudSec: Large Language Model Empowered Automatic and Deep Vulnerability Analysis for Intelligent Clouds. In International Conference on Computer Communications Workshops (INFOCOM WKSHPS). IEEE, Vancouver, Canada, 1–6. doi:10.1109/INFOCOMWKSHPS61880.2024.10620804
- Cao et al. (2024) Di Cao, Yong Liao, and Xiuwei Shang. 2024. RealVul: Can We Detect Vulnerabilities in Web Applications with LLM?. In Conference on Empirical Methods in Natural Language Processing (EMNLP). ACL, Miami, FL, USA, 8268–8282. doi:10.18653/v1/2024.emnlp-main.472
- Carbonnelle (2025) Pierre Carbonnelle. 2025. PYPL PopularitY of Programming Language Index. https://pypl.github.io/PYPL.html.
- Ceka et al. (2025) Ira Ceka, Feitong Qiao, Anik Dey, Aastha Valecha, Gail Kaiser, and Baishakhi Ray. 2025. Can LLM Prompting Serve as a Proxy for Static Analysis in Vulnerability Detection. arXiv:2412.12039
- CGCL Codes (2017) CGCL Codes. 2017. VulDeePecker: A Deep Learning-Based System for Vulnerability Detection. https://github.com/CGCL-codes/VulDeePecker.
- Chakraborty et al. (2024) Partha Chakraborty, Krishna Kanth Arumugam, Mahmoud Alfadel, Meiyappan Nagappan, and Shane McIntosh. 2024. Revisiting the Performance of Deep Learning-Based Vulnerability Detection on Realistic Datasets. IEEE Transactions on Software Engineering (TSE) 50, 8 (2024), 2163–2177. doi:10.1109/TSE.2024.3423712
- Chakraborty et al. (2022) Saikat Chakraborty, Rahul Krishna, Yangruibo Ding, and Baishakhi Ray. 2022. Deep Learning Based Vulnerability Detection: Are We There Yet? IEEE Transactions on Software Engineering (TSE) 48, 9 (2022), 3280–3296. doi:10.1109/TSE.2021.3087402
- Chan et al. (2023) Aaron Chan, Anant Kharkar, Roshanak Zilouchian Moghaddam, Yevhen Mohylevskyy, Alec Helyar, Eslam Kamal, Mohamed Elkamhawy, and Neel Sundaresan. 2023. Transformer-based Vulnerability Detection in Code at EditTime: Zero-shot, Few-shot, or Fine-tuning? arXiv:2306.01754
- Chang et al. (2024) Wei Chang, Chunyang Ye, and Hui Zhou. 2024. Fine-Tuning Pre-trained Model with Optimizable Prompt Learning for Code Vulnerability Detection. In International Symposium on Software Reliability Engineering (ISSRE). IEEE, Tsukuba, Japan, 108–119. doi:10.1109/ISSRE62328.2024.00021
- Chen et al. (2021) Mark Chen, Jerry Tworek, Heewoo Jun, Qiming Yuan, Henrique Ponde de Oliveira Pinto, Jared Kaplan, Harri Edwards, Yuri Burda, Nicholas Joseph, Greg Brockman, et al. 2021. Evaluating Large Language Models Trained on Code. arXiv:2107.03374
- Chen et al. (2025b) Siyu Chen, Jiongyi Yang, Xiang Chen, Menglin Zheng, Minnan Wei, and Xiaolin Ju. 2025b. Improving Vulnerability Type Prediction and Line-Level Detection via Adversarial Training-based Data Augmentation and Multi-Task Learning. arXiv:2506.23534
- Chen et al. (2023) Yizheng Chen, Zhoujie Ding, Lamya Alowain, Xinyun Chen, and David Wagner. 2023. DiverseVul: A New Vulnerable Source Code Dataset for Deep Learning Based Vulnerability Detection. In International Symposium on Research in Attacks, Intrusions and Defenses (RAID). IEEE, Hong Kong, China, 654–668. doi:10.1145/3607199.3607242
- Chen et al. (2024) Yujia Chen, Cuiyun Gao, Zezhou Yang, Hongyu Zhang, and Qing Liao. 2024. Bridge and Hint: Extending Pre-trained Language Models for Long-Range Code. In International Symposium on Software Testing and Analysis (ISSTA). ACM, Vienna, Austria, 274–286. doi:10.1145/3650212.3652127
- Chen et al. (2025a) Yinan Chen, Yuan Huang, Xiangping Chen, Pengfei Shen, and Lei Yun. 2025a. GPTVD: vulnerability detection and analysis method based on LLM’s chain of thoughts. Automated Software Engineering (ASE) 33, 1 (2025), 1–23. doi:10.1007/s10515-025-00550-4
- Chen and Zhang (2025) Zhuo Chen and Xuejun Zhang. 2025. VulPr: A Prompt Learning-based Method for Vulnerability Detection. In International Conference on Electronic Information Technology (EIT). IEEE, Valparaiso, IN, USA, 594–597. doi:10.1109/EIT67313.2025.11231886
- Cheng et al. (2025) Zeru Cheng, Yanjing Yang, He Zhang, Lanxin Yang, Jinghao Hu, Jinwei Xu, Bohan Liu, and Haifeng Shen. 2025. Leveraging Self-Paced Learning for Software Vulnerability Detection. arXiv:2511.09212
- Cheshkov et al. (2023) Anton Cheshkov, Pavel Zadorozhny, and Rodion Levichev. 2023. Evaluation of ChatGPT Model for Vulnerability Detection. arXiv:2304.07232
- Cipollone et al. (2025) Daniele Cipollone, Changjie Wang, Mariano Scazzariello, Simone Ferlin, Maliheh Izadi, Dejan Kostić, and Marco Chiesa. 2025. Automating the Detection of Code Vulnerabilities by Analyzing GitHub Issues. In International Workshop on Large Language Models for Code (LLM4Code). IEEE/ACM, Ottawa, Ontario, Canada, 41–48. doi:10.1109/LLM4Code66737.2025.00010
- Croft et al. (2023) Roland Croft, M. Ali Babar, and M. Mehdi Kholoosi. 2023. Data Quality for Software Vulnerability Datasets. In International Conference on Software Engineering (ICSE). IEEE/ACM, Melbourne, Australia, 121–133. doi:10.1109/ICSE48619.2023.00022
- Curto et al. (2024a) Claudio Curto, Daniela Giordano, Daniel Gustav Indelicato, and Vladimiro Patatu. 2024a. Can a Llama Be a Watchdog? Exploring Llama 3 and Code Llama for Static Application Security Testing. In International Conference on Cyber Security and Resilience (CSR). IEEE, London, UK, 395–400. doi:10.1109/CSR61664.2024.10679444
- Curto et al. (2024b) Claudio Curto, Daniela Giordano, Simone Palazzo, and D Indelicato. 2024b. MultiVD: A Transformer-based Multitask Approach for Software Vulnerability Detection. In International Conference on Security and Cryptography (SECRYPT). SCITEPRESS, Dijon, France, 416–423.
- CVE (2025a) CVE. 2025a. CVE Metrics. https://www.cve.org/About/Metrics.
- CVE (2025b) CVE. 2025b. CVE – Common Vulnerabilities and Exposures. https://www.cve.org/.
- d’Aloisio et al. (2025) Giordano d’Aloisio, Luca Traini, Federica Sarro, and Antinisca Di Marco. 2025. On the Compression of Language Models for Code: An Empirical Study on CodeBERT. In International Conference on Software Analysis, Evolution and Reengineering (SANER). IEEE, Montréal, Québec, Canada, 12–23. doi:10.1109/SANER64311.2025.00010
- DeepSeek-AI et al. (2025) DeepSeek-AI, Daya Guo, Dejian Yang, Haowei Zhang, Junxiao Song, Ruoyu Zhang, Runxin Xu, Qihao Zhu, Shirong Ma, Peiyi Wang, et al. 2025. DeepSeek-R1: Incentivizing Reasoning Capability in LLMs via Reinforcement Learning. arXiv:2501.12948
- DeepSeek-AI et al. (2024a) DeepSeek-AI, Aixin Liu, Bei Feng, Bin Wang, Bingxuan Wang, Bo Liu, Chenggang Zhao, Chengqi Dengr, Chong Ruan, Damai Dai, et al. 2024a. DeepSeek-V2: A Strong, Economical, and Efficient Mixture-of-Experts Language Model. arXiv:2405.04434
- DeepSeek-AI et al. (2024b) DeepSeek-AI, Qihao Zhu, Daya Guo, Zhihong Shao, Dejian Yang, Peiyi Wang, Runxin Xu, Y. Wu, Yukun Li, Huazuo Gao, et al. 2024b. DeepSeek-Coder-V2: Breaking the Barrier of Closed-Source Models in Code Intelligence. arXiv:2406.11931
- Deng et al. (2024) Xiao Deng, Fuyao Duan, Rui Xie, Wei Ye, and Shikun Zhang. 2024. Improving Long-Tail Vulnerability Detection Through Data Augmentation Based on Large Language Models. In International Conference on Software Maintenance and Evolution (ICSME). IEEE, Flagstaff, AZ, USA, 262–274. doi:10.1109/ICSME58944.2024.00033
- Dettmers et al. (2023) Tim Dettmers, Artidoro Pagnoni, Ari Holtzman, and Luke Zettlemoyer. 2023. QLoRA: Efficient Finetuning of Quantized LLMs. In Advances in Neural Information Processing Systems (NeurIPS), Vol. 36. Curran Associates, Inc., New Orleans, LA, USA, 10088–10115.
- Devlin et al. (2019) Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2019. Bert: Pre-training of Deep Bidirectional Transformers for Language Understanding. In Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies (NAACL HLT). ACL, Minneapolis, Minnesota, USA, 4171–4186.
- Ding et al. (2024a) Yangruibo Ding, Yanjun Fu, Omniyyah Ibrahim, Chawin Sitawarin, Xinyun Chen, Basel Alomair, David Wagner, Baishakhi Ray, and Yizheng Chen. 2024a. Vulnerability Detection with Code Language Models: How Far Are We? arXiv:2403.18624
- Ding et al. (2025) Yangruibo Ding, Yanjun Fu, Omniyyah Ibrahim, Chawin Sitawarin, Xinyun Chen, Basel Alomair, David Wagner, Baishakhi Ray, and Yizheng Chen. 2025. Vulnerability Detection with Code Language Models: How Far are We?. In International Conference on Software Engineering (ICSE). IEEE/ACM, Ottawa, Ontario, Canada, 1729–1741. doi:10.1109/ICSE55347.2025.00038
- Ding et al. (2024b) Yangruibo Ding, Benjamin Steenhoek, Kexin Pei, Gail Kaiser, Wei Le, and Baishakhi Ray. 2024b. TRACED: Execution-aware Pre-training for Source Code. In International Conference on Software Engineering (ICSE). IEEE/ACM, Lisbon, Portugal, Article 36, 12 pages. doi:10.1145/3597503.3608140
- Ding et al. (2023) Yue Ding, Qian Wu, Yinzhu Li, Dongdong Wang, and Jiaxin Huang. 2023. Leveraging Deep Learning Models for Cross-function Null Pointer Risks Detection. In International Conference On Artificial Intelligence Testing (AITest). IEEE, Athens, Greece, 107–113. doi:10.1109/AITest58265.2023.00025
- DLVulDet (2024) DLVulDet. 2024. PrimeVul: Vulnerability Detection with Code Language Models: How Far Are We? https://github.com/DLVulDet/PrimeVul.
- Dolcetti et al. (2025) Greta Dolcetti, Vincenzo Arceri, Eleonora Iotti, Sergio Maffeis, Agostino Cortesi, and Enea Zaffanella. 2025. Helping LLMs Improve Code Generation Using Feedback from Testing and Static Analysis. arXiv:2412.14841
- Dozono et al. (2024) Kohei Dozono, Tiago Espinha Gasiba, and Andrea Stocco. 2024. Large Language Models for Secure Code Assessment: A Multi-Language Empirical Study. arXiv:2408.06428
- Du et al. (2023) Qianjin Du, Shiji Zhou, Xiaohui Kuang, Gang Zhao, and Jidong Zhai. 2023. Joint Geometrical and Statistical Domain Adaptation for Cross-domain Code Vulnerability Detection. In Conference on Empirical Methods in Natural Language Processing (EMNLP). ACL, Singapore, 12791–12800.
- Du et al. (2024) Xiaohu Du, Ming Wen, Jiahao Zhu, Zifan Xie, Bin Ji, Huijun Liu, Xuanhua Shi, and Hai Jin. 2024. Generalization-Enhanced Code Vulnerability Detection via Multi-Task Instruction Fine-Tuning. In Findings of the Association for Computational Linguistics (ACL). ACL, Bangkok, Thailand, 10507–10521. doi:10.18653/v1/2024.findings-acl.625
- Du et al. (2025a) Xueying Du, Geng Zheng, Kaixin Wang, Yi Zou, Yujia Wang, Wentai Deng, Jiayi Feng, Mingwei Liu, Bihuan Chen, Xin Peng, Tao Ma, and Yiling Lou. 2025a. Vul-RAG: Enhancing LLM-based Vulnerability Detection via Knowledge-level RAG. arXiv:2406.11147
- Du et al. (2025b) Xiaozhi Du, Yanrong Zhou, and Hongyuan Du. 2025b. DMVL4AVD: A Deep Multi-View Learning Model for Automated Vulnerability Detection. Neural Computing and Applications (Neural Comput. Appl.) 37 (2025), 5873–5889. doi:10.1007/s00521-024-10892-x
- Dubniczky et al. (2025) Richard A Dubniczky, Krisztofer Zoltan Horvát, Tamás Bisztray, Mohamed Amine Ferrag, Lucas C Cordeiro, and Norbert Tihanyi. 2025. CASTLE: Benchmarking Dataset for Static Code Analyzers and LLMs towards CWE Detection. In International Symposium on Theoretical Aspects of Software Engineering (TASE). Springer, Limassol, Cyprus, 253–272.
- Eberhardt and Milánkovich (2024) Gergely Eberhardt and Ákos Milánkovich. 2024. VulnGPT: Enhancing Source Code Vulnerability Detection Using AutoGPT and Adaptive Supervision Strategies. In International Conference on Distributed Computing in Smart Systems and the Internet of Things (DCOSS-IoT). IEEE, Abu Dhabi, United Arab Emirates, 450–454. doi:10.1109/DCOSS-IoT61029.2024.00072
- Ehrenberg et al. (2024) Melanie Ehrenberg, Shahram Sarkani, and Thomas A Mazzuchi. 2024. Python Source Code Vulnerability Detection with Named Entity Recognition. Computers & Security (COSE) 140 (2024), 103802. doi:10.1016/j.cose.2024.103802
- Espinha Gasiba et al. (2024) Tiago Espinha Gasiba, Andrei-Cristian Iosif, Ibrahim Kessba, Sathwik Amburi, Ulrike Lechner, and Maria Pinto-Albuquerque. 2024. May the Source Be with You: On ChatGPT, Cybersecurity, and Secure Coding. Information 15, 9 (2024), 572. doi:10.3390/info15090572
- Ezenwoye et al. (2024) Onyeka Ezenwoye, Eduard Pinconschi, and Earl Roberts. 2024. Exploring AI for Vulnerability Detection and Repair. In Cyber Awareness and Research Symposium (CARS). IEEE, Grand Forks, ND, USA, 1–9. doi:10.1109/CARS61786.2024.10778769
- Fan (2020) Jiahao Fan. 2020. A C/C++ Code Vulnerability Dataset with Code Changes and CVE Summaries. https://github.com/ZeoVan/MSR_20_Code_vulnerability_CSV_Dataset.
- Fan et al. (2020) Jiahao Fan, Yi Li, Shaohua Wang, and Tien N. Nguyen. 2020. A C/C++ Code Vulnerability Dataset with Code Changes and CVE Summaries. In International Conference on Mining Software Repositories (MSR). ACM, Seoul, South Korea, 508–512. doi:10.1145/3379597.3387501
- Farr et al. (2025) David Farr, Kevin Talty, Alexandra Farr, John Stockdale, Iain Cruickshank, and Jevin West. 2025. Expert-in-the-Loop Systems with Cross-Domain and In-Domain Few-Shot Learning for Software Vulnerability Detection. arXiv:2506.10104
- Feng et al. (2025) Ruijun Feng, Hammond Pearce, Pietro Liguori, and Yulei Sui. 2025. CGP-Tuning: Structure-Aware Soft Prompt Tuning for Code Vulnerability Detection. arXiv:2501.04510
- Feng et al. (2020) Zhangyin Feng, Daya Guo, Duyu Tang, Nan Duan, Xiaocheng Feng, Ming Gong, Linjun Shou, Bing Qin, Ting Liu, Daxin Jiang, and Ming Zhou. 2020. CodeBERT: A Pre-Trained Model for Programming and Natural Languages. arXiv:2002.08155
- Ferrag et al. (2025a) Mohamed Amine Ferrag, Fatima Alwahedi, Ammar Battah, Bilel Cherif, Abdechakour Mechri, Norbert Tihanyi, Tamas Bisztray, and Merouane Debbah. 2025a. Generative AI in Cybersecurity: A Comprehensive Review of LLM Applications and Vulnerabilities. Internet of Things and Cyber-Physical Systems (IOTCPS) 5 (2025), 1–46. doi:10.1016/j.iotcps.2025.01.001
- Ferrag et al. (2025b) Mohamed Amine Ferrag, Ammar Battah, Norbert Tihanyi, Ridhi Jain, Diana Maimuţ, Fatima Alwahedi, Thierry Lestable, Narinderjit Singh Thandi, Abdechakour Mechri, Merouane Debbah, and Lucas C. Cordeiro. 2025b. SecureFalcon: Are We There Yet in Automated Software Vulnerability Detection With LLMs? Transactions on Software Engineering (TSE) 51, 4 (2025), 1248–1265. doi:10.1109/TSE.2025.3548168
- Flicke et al. (2025) Markus Flicke, Glenn Angrabeit, Madhav Iyengar, Vitalii Protsenko, Illia Shakun, Jovan Cicvaric, Bora Kargi, Haoyu He, Lukas Schuler, Lewin Scholz, Kavyanjali Agnihotri, Yong Cao, and Andreas Geiger. 2025. Scholar Inbox: Personalized Paper Recommendations for Scientists. arXiv:2504.08385
- FormAI Dataset (2023) FormAI Dataset. 2023. FormAI Dataset. https://github.com/FormAI-Dataset/FormAI-dataset.
- Fu et al. (2023a) Michael Fu, Van Nguyen, Chakkrit Kla Tantithamthavorn, Trung Le, and Dinh Phung. 2023a. VulExplainer: A Transformer-Based Hierarchical Distillation for Explaining Vulnerability Types. IEEE Transactions on Software Engineering (TSE) 49, 10 (2023), 4550–4565.
- Fu et al. (2025b) Michael Fu, Jirat Pasuksmit, and Chakkrit Tantithamthavorn. 2025b. AI for DevSecOps: A Landscape and Future Opportunities. ACM Transactions on Software Engineering and Methodology (TOSEM) 34, 4, Article 115 (2025), 61 pages. doi:10.1145/3712190
- Fu and Tantithamthavorn (2022) Michael Fu and Chakkrit Tantithamthavorn. 2022. LineVul: A Transformer-based Line-level Vulnerability Prediction. In International Conference on Mining Software Repositories (MSR). ACM, Pittsburgh, PA, USA, 608–620. doi:10.1145/3524842.3528452
- Fu et al. (2024) Michael Fu, Chakkrit Tantithamthavorn, Trung Le, Yuki Kume, Van Nguyen, Dinh Phung, and John Grundy. 2024. AIBugHunter: A Practical Tool for Predicting, Classifying and Repairing Software Vulnerabilities. Empirical Software Engineering (EMSE) 29, 1 (2024), 4.
- Fu et al. (2023b) Michael Fu, Chakkrit Kla Tantithamthavorn, Van Nguyen, and Trung Le. 2023b. ChatGPT for Vulnerability Detection, Classification, and Repair: How Far Are We?. In Asia-Pacific Software Engineering Conference (APSEC). IEEE, Seoul, South Korea, 632–636. doi:10.1109/APSEC60848.2023.00085
- Fu et al. (2025a) Yujia Fu, Peng Liang, Amjed Tahir, Zengyang Li, Mojtaba Shahin, Jiaxin Yu, and Jinfu Chen. 2025a. Security Weaknesses of Copilot-Generated Code in GitHub Projects: An Empirical Study. doi:10.1145/3716848
- Gajjar et al. (2025a) Jugal Gajjar, Kaustik Ranaware, and Kamalasankari Subramaniakuppusamy. 2025a. Bridging Semantics & Structure for Software Vulnerability Detection using Hybrid Network Models. arXiv:2510.10321
- Gajjar et al. (2025b) Jugal Gajjar, Kamalasankari Subramaniakuppusamy, and Noha El Kachach. 2025b. MalCodeAI: Autonomous Vulnerability Detection and Remediation via Language Agnostic Code Reasoning. In IEEE International Conference on Information Reuse and Integration and Data Science (IRI). IEEE, San Jose, CA, USA, 31–36. doi:10.1109/IRI66576.2025.00014
- Gao et al. (2023) Shuzheng Gao, Hongyu Zhang, Cuiyun Gao, and Chaozheng Wang. 2023. Keeping Pace with Ever-Increasing Data: Towards Continual Learning of Code Intelligence Models. In International Conference on Software Engineering (ICSE). IEEE/ACM, Melbourne, Australia, 30–42. doi:10.1109/ICSE48619.2023.00015
- Gao et al. (2025) Zeyu Gao, Junlin Zhou, Bolun Zhang, Yi He, Chao Zhang, Yuxin Cui, and Hao Wang. 2025. Mono: Is Your ”Clean” Vulnerability Dataset Really Solvable? Exposing and Trapping Undecidable Patches and Beyond. arXiv:2506.03651
- Germano et al. (2025) Lucas Bastos Germano, Ronaldo Ribeiro Goldschmidt, Ricardo Choren Noya, and Julio Cesar Duarte. 2025. A Systematic Review on Detection, Repair, and Explanation of Vulnerabilities in Source Code Using Large Language Models. IEEE Access 13 (2025), 192263–192293. doi:10.1109/ACCESS.2025.3631363
- Gnieciak and Szandala (2025) Damian Gnieciak and Tomasz Szandala. 2025. Large Language Models Versus Static Code Analysis Tools: A Systematic Benchmark for Vulnerability Detection. arXiv:2508.04448
- Gonçalves et al. (2025) José Gonçalves, Miguel Silva, Bernardo Cabral, Tiago Dias, Eva Maia, Isabel Praça, Ricardo Severino, and Luís Lino Ferreira. 2025. Evaluating LLaMA 3.2 for Software Vulnerability Detection. In European Interdisciplinary Cybersecurity Conference (EICC). Springer, Rennes, France, 38–51.
- Gong et al. (2025) Jianian Gong, Nachuan Duan, Ziheng Tao, Zhaohui Gong, Yuan Yuan, and Minlie Huang. 2025. How Well Do Large Language Models Serve as End-to-End Secure Code Agents for Python? arXiv:2408.10495
- Grattafiori et al. (2024) Aaron Grattafiori, Abhimanyu Dubey, Abhinav Jauhri, Abhinav Pandey, Abhishek Kadian, Ahmad Al-Dahle, Aiesha Letman, Akhil Mathur, Alan Schelten, Alex Vaughan, et al. 2024. The Llama 3 Herd of Models. arXiv:2407.21783
- Grishina et al. (2023) Anastasiia Grishina, Max Hort, and Leon Moonen. 2023. The EarlyBIRD Catches the Bug: On Exploiting Early Layers of Encoder Models for More Efficient Code Classification. In Joint European Software Engineering Conference and Symposium on the Foundations of Software Engineering (ESEC/FSE). ACM, San Francisco, CA, USA, 895–907. doi:10.1145/3611643.3616304
- Gujar (2024) Shantanu Sudhir Gujar. 2024. DetectBERT: Code Vulnerability Detection. In Global Conference on Communications and Information Technologies (GCCIT). IEEE, Bangalore, India, 1–21. doi:10.1109/GCCIT63234.2024.10862235
- Guo et al. (2022) Daya Guo, Shuai Lu, Nan Duan, Yanlin Wang, Ming Zhou, and Jian Yin. 2022. UniXcoder: Unified Cross-Modal Pre-training for Code Representation. arXiv:2203.03850
- Guo et al. (2021) Daya Guo, Shuo Ren, Shuai Lu, Zhangyin Feng, Duyu Tang, Shujie Liu, Long Zhou, Nan Duan, Alexey Svyatkovskiy, Shengyu Fu, et al. 2021. GraphCodeBERT: Pre-training Code Representations with Data Flow. arXiv:2009.08366
- Guo et al. (2024c) Daya Guo, Qihao Zhu, Dejian Yang, Zhenda Xie, Kai Dong, Wentao Zhang, Guanting Chen, Xiao Bi, Y. Wu, Y. K. Li, et al. 2024c. DeepSeek-Coder: When the Large Language Model Meets Programming – The Rise of Code Intelligence. arXiv:2401.14196
- Guo et al. (2025) Jinyao Guo, Chengpeng Wang, Xiangzhe Xu, Zian Su, and Xiangyu Zhang. 2025. RepoAudit: An Autonomous LLM-Agent for Repository-Level Code Auditing. arXiv:2501.18160
- Guo and Bettaieb (2023) Yuejun Guo and Seifeddine Bettaieb. 2023. An Investigation of Quality Issues in Vulnerability Detection Datasets. In European Symposium on Security and Privacy Workshops (EuroS&PW). IEEE, Delft, Netherlands, 29–33. doi:10.1109/EuroSPW59978.2023.00008
- Guo et al. (2024a) Yuejun Guo, Seifeddine Bettaieb, and Fran Casino. 2024a. A Comprehensive Analysis on Software Vulnerability Detection Datasets: Trends, Challenges, and Road Ahead. International Journal of Information Security 23, 5 (2024), 3311–3327. doi:10.1007/s10207-024-00888-y
- Guo et al. (2024b) Yuejun Guo, Constantinos Patsakis, Qiang Hu, Qiang Tang, and Fran Casino. 2024b. Outside the Comfort Zone: Analysing LLM Capabilities in Software Vulnerability Detection. In European Symposium on Research in Computer Security (ESORICS). Springer, Bydgoszcz, Poland, 271–289. doi:10.1007/978-3-031-70879-4_14
- G”̈urfidan (2024) Remzi G”̈urfidan. 2024. VULREM: Fine-Tuned BERT-Based Source-Code Potential Vulnerability Scanning System to Mitigate Attacks in Web Applications. Applied Sciences 14, 21 (2024), 9697. doi:10.3390/app14219697
- Hairan and Şahman (2025) Borhanullah Hairan and Mehmet Akif Şahman. 2025. Compressing Large Language Models for SQL Injection Detection: A Case Study on Deep Seek-Coder and Meta-llama-3-70b-instruct. In Conference of Open Innovations Association (FRUCT). FRUCT Oy, Helsinki, Finland, 90–96. doi:10.23919/FRUCT67853.2025.11239157
- Halder et al. (2025) Sajal Halder, Muhammad Ejaz Ahmed, and Seyit Camtepe. 2025. FuncVul: An Effective Function Level Vulnerability Detection Model using LLM and Code Chunk. In European Symposium on Research in Computer Security (ESORICS). Springer, Toulouse, France, 166–185. doi:10.1007/978-3-032-07884-1_9
- Han and Zheng (2025) Yaolin Han and Wei Zheng. 2025. An Automatic Classification Model for Long Code Vulnerabilities Based on the Teacher-Student Framework. In International Conference on Software Quality, Reliability, and Security (QRS) Companion. IEEE, Hangzhou, China, 472–482. doi:10.1109/QRS-C65679.2025.00065
- Han et al. (2024) Zeyu Han, Chao Gao, Jinyang Liu, Jeff Zhang, and Sai Qian Zhang. 2024. Parameter-Efficient Fine-Tuning for Large Models: A Comprehensive Survey. arXiv:2403.14608
- Hanif and Maffeis (2022) Hazim Hanif and Sergio Maffeis. 2022. VulBERTa: Simplified Source Code Pre-Training for Vulnerability Detection. In International Joint Conference on Neural Networks (IJCNN). IEEE, Padua, Italy, 1–8. doi:10.1109/IJCNN55064.2022.9892280
- Hannan et al. (2025) Md Abdul Hannan, Ronghao Ni, Chi Zhang, Limin Jia, Ravi Mangal, and Corina S. Pasareanu. 2025. On Selecting Few-Shot Examples for LLM-based Code Vulnerability Detection. arXiv:2510.27675
- Haque et al. (2025) Radowanul Haque, Aftab Ali, Sally McClean, and Naveed Khan. 2025. A Zero-Shot Framework for Cross-Project Vulnerability Detection in Source Code. Empirical Software Engineering (EMSE) 31, 1 (2025), 3. doi:10.1007/s10664-025-10749-4
- He and Vechev (2023) Jingxuan He and Martin Vechev. 2023. Large Language Models for Code: Security Hardening and Adversarial Testing. In Conference on Computer and Communications Security (CCS). ACM, Copenhagen, Denmark, 1865–1879. doi:10.1145/3576915.3623175
- Hin et al. (2022) David Hin, Andrey Kan, Huaming Chen, and M. Ali Babar. 2022. LineVD: Statement-level Vulnerability Detection using Graph Neural Networks. In International Conference on Mining Software Repositories (MSR). ACM, Pittsburgh, PA, USA, 596–607. doi:10.1145/3524842.3527949
- Hort et al. (2025) Max Hort, Linas Vidziunas, and Leon Moonen. 2025. Semantic-Preserving Transformations as Mutation Operators: A Study on Their Effectiveness in Defect Detection. In International Conference on Software Testing, Verification and Validation Workshops (ICSTW). IEEE, Naples, Italy, 337–346. doi:10.1109/ICSTW64639.2025.10962512
- Hou et al. (2024) Xinyi Hou, Yanjie Zhao, Yue Liu, Zhou Yang, Kailong Wang, Li Li, Xiapu Luo, David Lo, John Grundy, and Haoyu Wang. 2024. Large Language Models for Software Engineering: A Systematic Literature Review. ACM Transactions on Software Engineering and Methodology (TOSEM) 33, 8, Article 220 (2024), 79 pages. doi:10.1145/3695988
- Hu et al. (2022) Edward J. Hu, Yelong Shen, Phillip Wallis, Zeyuan Allen-Zhu, Yuanzhi Li, Shean Wang, Lu Wang, and Weizhu Chen. 2022. LoRA: Low-Rank Adaptation of Large Language Models. International Conference on Learning Representations (ICLR) 1, 2 (2022), 3.
- Hu et al. (2025) Xing Hu, Feifei Niu, Junkai Chen, Xin Zhou, Junwei Zhang, Junda He, Xin Xia, and David Lo. 2025. Assessing and Advancing Benchmarks for Evaluating Large Language Models in Software Engineering Tasks. arXiv:2505.08903
- Huang et al. (2022) Weichang Huang, Shuyuan Lin, and Li Chen. 2022. BBVD: A BERT-based Method for Vulnerability Detection. International Journal of Advanced Computer Science and Applications (IJACSA) 13, 12 (2022). doi:10.14569/IJACSA.2022.01312103
- Hui et al. (2024) Binyuan Hui, Jian Yang, Zeyu Cui, Jiaxi Yang, Dayiheng Liu, Lei Zhang, Tianyu Liu, Jiajun Zhang, Bowen Yu, Keming Lu, et al. 2024. Qwen2.5-Coder Technical Report. arXiv:2409.12186
- Huynh et al. (2025) Larry Huynh, Yinghao Zhang, Djimon Jayasundera, Woojin Jeon, Hyoungshick Kim, Tingting Bi, and Jin B. Hong. 2025. Detecting Code Vulnerabilities using LLMs. In International Conference on Dependable Systems and Networks (DSN). IEEE/IFIP, Naples, Italy, 401–414. doi:10.1109/DSN64029.2025.00047
- Ibanez-Lissen et al. (2025) Luis Ibanez-Lissen, Lorena Gonzalez-Manzano, Jose Maria de Fuentes, and Nicolas Anciaux. 2025. LPASS: Linear Probes as Stepping Stones for Vulnerability Detection using Compressed LLMs. Journal of Information Security and Applications (JISA) 93 (2025), 104125. doi:10.1016/j.jisa.2025.104125
- IBM (2021) IBM. 2021. D2A Dataset and Generation Pipeline. https://github.com/IBM/D2A.
- Icyrockton (2023) Icyrockton. 2023. MegaVul. https://github.com/Icyrockton/MegaVul.
- Ingemann Tuffveson Jensen et al. (2024) Rasmus Ingemann Tuffveson Jensen, Vali Tawosi, and Salwa Alamir. 2024. Software Vulnerability and Functionality Assessment using Large Language Models. In International Workshop on NL-Based Software Engineering (NLBSE). ACM/IEEE, Lisbon, Portugal, 25–28. doi:10.1145/3643787.3648036
- Islam et al. (2024a) Nafis Tanveer Islam, Mazal Bethany, Dylan Manuel, Murtuza Jadliwala, and Peyman Najafirad. 2024a. Unintentional Security Flaws in Code: Automated Defense via Root Cause Analysis. arXiv:2409.00199
- Islam et al. (2023) Nafis Tanveer Islam, Gonzalo De La Torre Parra, Dylan Manuel, Elias Bou-Harb, and Peyman Najafirad. 2023. An Unbiased Transformer Source Code Learning with Semantic Vulnerability Graph. In European Symposium on Security and Privacy (EuroS&P). IEEE, Delft, Netherlands, 144–159. doi:10.1109/EuroSP57164.2023.00018
- Islam et al. (2024b) Nafis Tanveer Islam, Joseph Khoury, Andrew Seong, Elias Bou-Harb, and Peyman Najafirad. 2024b. Enhancing Source Code Security with LLMs: Demystifying The Challenges and Generating Reliable Repairs. arXiv:2409.00571
- Jain et al. (2023) Ridhi Jain, Nicole Gervasoni, Mthandazo Ndhlovu, and Sanjay Rawat. 2023. A Code Centric Evaluation of C/C++ Vulnerability Datasets for Deep Learning Based Vulnerability Detection Techniques. In Innovations in Software Engineering Conference (ISEC). ACM, Allahabad, India, Article 6, 10 pages. doi:10.1145/3578527.3578530
- Jararweh et al. (2025) Ala Jararweh, Michael Adams, Avinash Sahu, Abdullah Mueen, and Afsah Anwar. 2025. LLaVul: A Multimodal LLM for Interpretable Vulnerability Reasoning about Source Code. In Intelligent Cybersecurity Conference (ICSC). IEEE, Tampa, FL, USA, 232–241. doi:10.1109/ICSC65596.2025.11140501
- Javaheripi et al. (2023) Mojan Javaheripi, Sébastien Bubeck, Marah Abdin, Jyoti Aneja, Sebastien Bubeck, Caio César Teodoro Mendes, Weizhu Chen, Allie Del Giorno, Ronen Eldan, Sivakanth Gopi, et al. 2023. Phi-2: The Surprising Power of Small Language Models. Microsoft Research Blog 1, 3 (2023), 3.
- Ji et al. (2024) Chen Ji, Su Yang, Hongyu Sun, and Yuqing Zhang. 2024. Applying Contrastive Learning to Code Vulnerability Type Classification. In Conference on Empirical Methods in Natural Language Processing (EMNLP). ACL, Miami, FL, USA, 11942–11952.
- Jiang et al. (2023) Albert Q. Jiang, Alexandre Sablayrolles, Arthur Mensch, Chris Bamford, Devendra Singh Chaplot, Diego de las Casas, Florian Bressand, Gianna Lengyel, Guillaume Lample, Lucile Saulnier, et al. 2023. Mistral 7B. arXiv:2310.06825
- Jiang et al. (2024a) Albert Q. Jiang, Alexandre Sablayrolles, Antoine Roux, Arthur Mensch, Blanche Savary, Chris Bamford, Devendra Singh Chaplot, Diego de las Casas, Emma Bou Hanna, Florian Bressand, et al. 2024a. Mixtral of Experts. arXiv:2401.04088
- Jiang et al. (2025b) Xuefeng Jiang, Lvhua Wu, Sheng Sun, Jia Li, Jingjing Xue, Yuwei Wang, Tingting Wu, and Min Liu. 2025b. Investigating Large Language Models for Code Vulnerability Detection: An Experimental Study. arXiv:2412.18260
- Jiang et al. (2022) Xuxiang Jiang, Yinhao Xiao, Jun Wang, and Wei Zhang. 2022. Multi-view Pre-trained Model for Code Vulnerability Identification. In International Conference on Wireless Algorithms, Systems, and Applications (WASA). Springer, Dalian, China, 127–135. doi:10.1007/978-3-031-19211-1_11
- Jiang et al. (2025a) Yuan Jiang, Zhichen Qu, Christoph Treude, Xiaohong Su, and Tiantian Wang. 2025a. Enhancing Fine-Grained Vulnerability Detection With Reinforcement Learning. IEEE Transactions on Software Engineering (TSE) 51, 10 (2025), 2900–2920. doi:10.1109/TSE.2025.3603400
- Jiang et al. (2024c) Yuan Jiang, Yujian Zhang, Xiaohong Su, Christoph Treude, and Tiantian Wang. 2024c. StagedVulBERT: Multigranular Vulnerability Detection With a Novel Pretrained Code Model. IEEE Transactions on Software Engineering (TSE) 50, 12 (2024), 3454–3471. doi:10.1109/TSE.2024.3493245
- Jiang et al. (2024b) Zhonghao Jiang, Weifeng Sun, Xiaoyan Gu, Jiaxin Wu, Tao Wen, Haibo Hu, and Meng Yan. 2024b. DFEPT: Data Flow Embedding for Enhancing Pre-Trained Model Based Vulnerability Detection. In Asia-Pacific Symposium on Internetware (Internetware). ACM, Macau, China, 95–104. doi:10.1145/3671016.3671388
- JianJie and Le (2023) Ye JianJie and Wei Le. 2023. Code Defect Detection Method Based on BERT and Ensemble. In International Conference on Computer and Communications (ICCC). IEEE, Chengdu, China, 2130–2138. doi:10.1109/ICCC59590.2023.10507306
- Jiao et al. (2025) Yutong Jiao, Jiaxuan Han, and Cheng Huang. 2025. DeepVulHunter: Enhancing the Code Vulnerability Detection Capability of LLMs through Multi-Round Analysis. Journal of Intelligent Information Systems (JIIS) 63, 6 (2025), 2237–2264. doi:10.1007/s10844-025-00982-0
- Kaanan et al. (2024) Emran Kaanan, Tasmin Karim, Md. Shazzad Hossain Shaon, Md. Fahim Sultan, Alfredo Cuzzocrea, and Mst Shapna Akter. 2024. LLM-Based Approach for Buffer Overflow Detection in Source Code. In International Conference on Computer and Information Technology (ICCIT). IEEE, Cox’s Bazar, Bangladesh, 1898–1902. doi:10.1109/ICCIT64611.2024.11021816
- Kalouptsoglou et al. (2024a) Ilias Kalouptsoglou, Miltiadis Siavvas, Apostolos Ampatzoglou, Dionysios Kehagias, and Alexander Chatzigeorgiou. 2024a. Vulnerability Classification on Source Code Using Text Mining and Deep Learning Techniques. In International Conference on Software Quality, Reliability, and Security Companion (QRS-C). IEEE, Cambridge, UK, 47–56. doi:10.1109/QRS-C63300.2024.00017
- Kalouptsoglou et al. (2024b) Ilias Kalouptsoglou, Miltiadis Siavvas, Apostolos Ampatzoglou, Dionysios Kehagias, and Alexander Chatzigeorgiou. 2024b. Vulnerability Prediction using Pre-trained Models: An Empirical Evaluation. In International Conference on Modeling, Analysis and Simulation of Computer and Telecommunication Systems (MASCOTS). IEEE, Paris, France, 1–6. doi:10.1109/MASCOTS64422.2024.10786510
- Kaniewski et al. (2024) Sabrina Kaniewski, Dieter Holstein, Fabian Schmidt, and Tobias Heer. 2024. Vulnerability Handling of AI-Generated Code - Existing Solutions and Open Challenges. In Conference on AI, Science, Engineering, and Technology (AIxSET). IEEE, Laguna Hills, CA, USA, 145–148. doi:10.1109/AIxSET62544.2024.00026
- Kaniewski et al. (2025) Sabrina Kaniewski, Fabian Schmidt, Markus Enzweiler, Michael Menth, and Tobias Heer. 2025. Artifacts - Awesome-LLM4SVD. https://github.com/hs-esslingen-it-security/Awesome-LLM4SVD.
- Keltek et al. (2024) Mete Keltek, Rong Hu, Mohammadreza Fani Sani, and Ziyue Li. 2024. Boosting Cybersecurity Vulnerability Scanning based on LLM-supported Static Application Security Testing. arXiv:2409.15735
- Khare et al. (2025) Avishree Khare, Saikat Dutta, Ziyang Li, Alaia Solko-Breslin, Rajeev Alur, and Mayur Naik. 2025. Understanding the Effectiveness of Large Language Models in Detecting Security Vulnerabilities. In Conference on Software Testing, Verification and Validation (ICST). IEEE, Naples, Italy, 103–114. doi:10.1109/ICST62969.2025.10988968
- Kholoosi et al. (2024) M. Mehdi Kholoosi, M. Ali Babar, and Roland Croft. 2024. A Qualitative Study on Using ChatGPT for Software Security: Perception vs. Practicality. In International Conference on Trust, Privacy and Security in Intelligent Systems, and Applications (TPS-ISA). IEEE, Washington, DC, USA, 107–117. doi:10.1109/TPS-ISA62245.2024.00022
- Kim and Russell (2018) Louis Kim and Rebecca Russell. 2018. Draper VDISC Dataset - Vulnerability Detection in Source Code. https://osf.io/d45bw/.
- Kim et al. (2022) Soolin Kim, Jusop Choi, Muhammad Ejaz Ahmed, Surya Nepal, and Hyoungshick Kim. 2022. VulDeBERT: A Vulnerability Detection System Using BERT. In International Symposium on Software Reliability Engineering Workshops (ISSREW). IEEE, Charlotte, NC, USA, 69–74. doi:10.1109/ISSREW55968.2022.00042
- Kitchenham and Charters (2007) Barbara Kitchenham and Stuart Charters. 2007. Guidelines for Performing Systematic Literature Reviews in Software Engineering. Technical Report. Keele University.
- KnowledgeRAG4LLMVulD (2024) KnowledgeRAG4LLMVulD. 2024. KnowledgeRAG4LLMVulD. https://github.com/KnowledgeRAG4LLMVulD/KnowledgeRAG4LLMVulD/tree/main/dataset.
- Kong et al. (2024) Zihan Kong, Yong Piao, and Tianran Li. 2024. A Source Code Vulnerability Detection Method Based on Positive-Unlabeled Learning. In International Conference on Robotics, Intelligent Control and Artificial Intelligence (RICAI). IEEE, Nanjing, China, 551–556. doi:10.1109/RICAI64321.2024.10911761
- Kouliaridis et al. (2024) Vasileios Kouliaridis, Georgios Karopoulos, and Georgios Kambourakis. 2024. Assessing the Effectiveness of LLMs in Android Application Vulnerability Analysis. In International Conference on Attacks and Defenses for Internet-of-Things (ADIoT). Springer, Hangzhou, China, 139–154. doi:10.1007/978-3-031-85593-1_9
- Lan et al. (2025) Xiang Lan, Tim Menzies, and Bowen Xu. 2025. Smart Cuts: Enhance Active Learning for Vulnerability Detection by Pruning Bad Seeds. arXiv:2506.20444
- Le et al. (2024) Triet Huynh Minh Le, M. Ali Babar, and Tung Hoang Thai. 2024. Software Vulnerability Prediction in Low-Resource Languages: An Empirical Study of CodeBERT and ChatGPT. In International Conference on Evaluation and Assessment in Software Engineering (EASE). ACM, Istanbul, Turkey, 679–685. doi:10.1145/3661167.3661281
- Le et al. (2025) Thi-Thu-Huong Le, Andro Aprila Adiputra, YeonJeong Hwang, JunYoung Son, and Howon Kim. 2025. Fine-Tuning Transformer LLMs for Detecting SQL Injection and XSS Vulnerabilities. In International Conference on Artificial Intelligence in Information and Communication (ICAIIC). IEEE, Fukuoka, Japan, 0946–0951. doi:10.1109/ICAIIC64266.2025.10920868
- Lekssays et al. (2025) Ahmed Lekssays, Hamza Mouhcine, Khang Tran, Ting Yu, and Issa Khalil. 2025. LLMxCPG: Context-Aware Vulnerability Detection Through Code Property Graph-Guided Large Language Models. In 34th USENIX Security Symposium (USENIX Security). USENIX, Seattle, WA, USA, 489–507.
- Lewis et al. (2020) Patrick Lewis, Ethan Perez, Aleksandra Piktus, Fabio Petroni, Vladimir Karpukhin, Naman Goyal, Heinrich Küttler, Mike Lewis, Wen-tau Yih, Tim Rocktäschel, et al. 2020. Retrieval-Augmented Generation for Knowledge-Intensive NLP Tasks. In Advances in Neural Information Processing Systems (NeurIPS), Vol. 33. Curran Associates, Inc., Virtual-only, 9459–9474.
- Li et al. (2024) Haonan Li, Yu Hao, Yizhuo Zhai, and Zhiyun Qian. 2024. Enhancing Static Analysis for Practical Bug Detection: An LLM-Integrated Approach. Proceedings of the ACM on Programming Languages (PACMPL) 8, OOPSLA1 (2024), 26 pages. doi:10.1145/3649828
- Li et al. (2025c) Jiayuan Li, Lei Cui, Wenyan Yu, Haiqiang Fei, Feng Cheng, and Hongsong Zhu. 2025c. VulnTeam: A Team Collaboration Framework for LLM-based Vulnerability Detection. In International Joint Conference on Neural Networks (IJCNN). IEEE, Rome, Italy, 1–8. doi:10.1109/IJCNN64981.2025.11229292
- Li et al. (2025d) Jiayuan Li, Lei Cui, Jie Zhang, Haiqiang Fei, Yu Chen, and Hongsong Zhu. 2025d. Steering Large Language Models for Vulnerability Detection. In International Conference on Acoustics, Speech and Signal Processing (ICASSP). IEEE, Hyderabad, India, 1–5. doi:10.1109/ICASSP49660.2025.10887736
- Li et al. (2025e) Jiayuan Li, Lei Cui, Sen Zhao, Yun Yang, Lun Li, and Hongsong Zhu. 2025e. CLeVeR: Multi-modal Contrastive Learning for Vulnerability Code Representation. In Findings of the Association for Computational Linguistics (ACL). ACL, Vienna, Austria, 7940–7951. doi:10.18653/v1/2025.findings-acl.414
- Li et al. (2021) Na Li, Haoyu Zhang, Zhihui Hu, Guang Kou, and Huadong Dai. 2021. Automated Software Vulnerability Detection via Pre-trained Context Encoder and Self Attention. In International Conference on Digital Forensics and Cyber Crime (ICDF2C). Springer, Singapore, 248–264.
- Li et al. (2023) Raymond Li, Loubna Ben Allal, Yangtian Zi, Niklas Muennighoff, Denis Kocetkov, Chenghao Mou, Marc Marone, Christopher Akiki, Jia Li, Jenny Chim, et al. 2023. StarCoder: may the source be with you! arXiv:2305.06161
- Li (2024) Yikun Li. 2024. CleanVul. https://github.com/yikun-li/CleanVul.
- Li et al. (2025a) Yansong Li, Paula Branco, Alexander M. Hoole, Manish Marwah, Hari Manassery Koduvely, Guy-Vincent Jourdan, and Stephan Jou. 2025a. SV-TrustEval-C: Evaluating Structure and Semantic Reasoning in Large Language Models for Source Code Vulnerability Analysis. In Symposium on Security and Privacy (SP). IEEE, San Francisco, CA, USA, 3014–3032. doi:10.1109/SP61157.2025.00191
- Li et al. (2025b) Yikun Li, Ngoc Tan Bui, Ting Zhang, Martin Weyssow, Chengran Yang, Xin Zhou, Jinfeng Jiang, Junkai Chen, Huihui Huang, Huu Hung Nguyen, Chiok Yew Ho, Jie Tan, Ruiyin Li, Yide Yin, Han Wei Ang, Frank Liauw, Eng Lieh Ouh, Lwin Khin Shar, and David Lo. 2025b. Out of Distribution, Out of Luck: How Well Can LLMs Trained on Vulnerability Datasets Detect Top 25 CWE Weaknesses? arXiv:2507.21817
- Li et al. (2025h) Youpeng Li, Kartik Joshi, Xinda Wang, and Eric Wong. 2025h. MAVUL: Multi-Agent Vulnerability Detection via Contextual Reasoning and Interactive Refinement. arXiv:2510.00317
- Li et al. (2025i) Yue Li, Xiao Li, Hao Wu, Minghui Xu, Yue Zhang, Xiuzhen Cheng, Fengyuan Xu, and Sheng Zhong. 2025i. Everything You Wanted to Know About LLM-based Vulnerability Detection But Were Afraid to Ask. arXiv:2504.13474
- Li et al. (2025j) Youpeng Li, Weiliang Qi, Xuyu Wang, Fuxun Yu, and Xinda Wang. 2025j. Revisiting Pre-trained Language Models for Vulnerability Detection. arXiv:2507.16887
- Li et al. (2025k) Yikun Li, Ting Zhang, Ratnadira Widyasari, Yan Naing Tun, Huu Hung Nguyen, Tan Bui, Ivana Clairine Irsan, Yiran Cheng, Xiang Lan, Han Wei Ang, Frank Liauw, Martin Weyssow, Hong Jin Kang, Eng Lieh Ouh, Lwin Khin Shar, and David Lo. 2025k. CleanVul: Automatic Function-Level Vulnerability Detection in Code Commits Using LLM Heuristics. arXiv:2411.17274
- Li et al. (2025f) Ziyang Li, Saikat Dutta, and Mayur Naik. 2025f. IRIS: LLM-Assisted Static Analysis for Detecting Security Vulnerabilities. In International Conference on Learning Representations (ICLR). ICLR, Singapore, Singapore.
- Li et al. (2025g) Zhihao Li, Zimo Ji, Tao Zheng, Hao Ren, and Xiao Lan. 2025g. CryptoScope: Utilizing Large Language Models for Automated Cryptographic Logic Vulnerability Detection. arXiv:2508.11599
- Li et al. (2018) Zhen Li, Deqing Zou, Shouhuai Xu, Xinyu Ou, Hai Jin, Sujuan Wang, Zhijun Deng, and Yuyi Zhong. 2018. VulDeePecker: A Deep Learning-Based System for Vulnerability Detection. In Network and Distributed Systems Security Symposium (NDSS). The Internet Society, San Diego, CA, USA, 1–15. doi:10.14722/ndss.2018.23158
- Liang et al. (2024) Chen Liang, Qiang Wei, Zirui Jiang, Yisen Wang, and Jiang Du. 2024. A Source Code Vulnerability Detection Method Based on Adaptive Graph Neural Networks. In IEEE/ACM International Conference on Automated Software Engineering (ASE) Workshops. IEEE, Sacramento, CA, USA, 187–196.
- Lin et al. (2022a) Guanjun Lin, Heming Jia, and Di Wu. 2022a. Distilled and Contextualized Neural Models Benchmarked for Vulnerable Function Detection. Mathematics 10, 23 (2022), 4482. doi:10.3390/math10234482
- Lin et al. (2020) Guanjun Lin, Sheng Wen, Qing-Long Han, Jun Zhang, and Yang Xiang. 2020. Software Vulnerability Detection Using Deep Neural Networks: A Survey. Proc. IEEE 108, 10 (2020), 1825–1848. doi:10.1109/JPROC.2020.2993293
- Lin and Mohaisen (2024) Jie Lin and David Mohaisen. 2024. Evaluating Large Language Models in Vulnerability Detection Under Variable Context Windows. In International Conference on Machine Learning and Applications (ICMLA). IEEE, Miami, FL, USA, 1131–1134. doi:10.1109/ICMLA61862.2024.00173
- Lin et al. (2022b) Yuhao Lin, Ying Li, Mianxue Gu, Hongyu Sun, Qiuling Yue, Jinglu Hu, Chunjie Cao, and Yuqing Zhang. 2022b. Vulnerability Dataset Construction Methods Applied To Vulnerability Detection: A Survey. In International Conference on Dependable Systems and Networks Workshops (DSN-W). IEEE, Baltimore, MD, USA, 141–146. doi:10.1109/DSN-W54100.2022.00032
- Liu et al. (2024a) Chongyang Liu, Xiang Chen, Xiangwei Li, and Yinxing Xue. 2024a. Making Vulnerability Prediction more Practical: Prediction, Categorization, and Localization. Information and Software Technology (IST) 171 (2024), 107458. doi:10.1016/j.infsof.2024.107458
- Liu et al. (2024c) Fang Liu, Yang Liu, Lin Shi, Houkun Huang, Ruifeng Wang, Zhen Yang, Li Zhang, Zhongqi Li, and Yuchi Ma. 2024c. Exploring and Evaluating Hallucinations in LLM-Powered Code Generation. arXiv:2404.00971
- Liu et al. (2025a) Jianing Liu, Guanjun Lin, Huan Mei, Fan Yang, and Yonghang Tai. 2025a. Enhancing Vulnerability Detection Efficiency: An Exploration of Light-weight LLMs with Hybrid Code Features. Journal of Information Security and Applications (JISA) 88 (2025), 103925. doi:10.1016/j.jisa.2024.103925
- Liu et al. (2025b) Ruitong Liu, Yanbin Wang, Haitao Xu, Jianguo Sun, Fan Zhang, Peiyue Li, and Zhenhao Guo. 2025b. Vul-LMGNNs: Fusing Language Models and Online-distilled Graph Neural Networks for Code Vulnerability Detection. Information Fusion 115 (2025), 102748. doi:10.1016/j.inffus.2024.102748
- Liu et al. (2026) Xiangyue Liu, Xinwei Liu, Lili Bo, Xiaoxue Wu, Yun Yang, Xiaobing Sun, and Feng Zhou. 2026. PIONEER: Improving the Robustness of Student Models when Compressing Pre-Trained Models of Code. Automated Software Engineering (ASE) 33, 1 (2026), 1–30. doi:10.1007/s10515-025-00560-2
- Liu et al. (2024b) Yu Liu, Lang Gao, Mingxin Yang, Yu Xie, Ping Chen, Xiaojin Zhang, and Wei Chen. 2024b. VulDetectBench: Evaluating the Deep Capability of Vulnerability Detection with Large Language Models. arXiv:2406.07595
- Liu et al. (2019) Yinhan Liu, Myle Ott, Naman Goyal, Jingfei Du, Mandar Joshi, Danqi Chen, Omer Levy, Mike Lewis, Luke Zettlemoyer, and Veselin Stoyanov. 2019. RoBERTa: A Robustly Optimized BERT Pretraining Approach. arXiv:1907.11692
- Liu et al. (2023) Zhihong Liu, Qing Liao, Wenchao Gu, and Cuiyun Gao. 2023. Software Vulnerability Detection with GPT and In-Context Learning. In International Conference on Data Science in Cyberspace (DSC). IEEE, Hefei, China, 229–236. doi:10.1109/DSC59305.2023.00041
- Liu et al. (2024d) Zhijie Liu, Yutian Tang, Xiapu Luo, Yuming Zhou, and Liang Feng Zhang. 2024d. No Need to Lift a Finger Anymore? Assessing the Quality of Code Generation by ChatGPT. IEEE Transactions on Software Engineering (TSE) 50, 6 (2024), 1548–1584. doi:10.1109/TSE.2024.3392499
- Liu et al. (2024e) Zhongxin Liu, Zhijie Tang, Junwei Zhang, Xin Xia, and Xiaohu Yang. 2024e. Pre-training by Predicting Program Dependencies for Vulnerability Analysis Tasks. In International Conference on Software Engineering (ICSE). IEEE/ACM, Lisbon, Portugal, 1863–1875. doi:10.1145/3597503.3639142
- Liu et al. (2024f) Zhihong Liu, Zezhou Yang, and Qing Liao. 2024f. Exploration On Prompting LLM With Code-Specific Information For Vulnerability Detection. In International Conference on Software Services Engineering (SSE). IEEE, Shenzhen, China, 273–281. doi:10.1109/SSE62657.2024.00049
- Lozhkov et al. (2024) Anton Lozhkov, Raymond Li, Loubna Ben Allal, Federico Cassano, Joel Lamy-Poirier, Nouamane Tazi, Ao Tang, Dmytro Pykhtar, Jiawei Liu, Yuxiang Wei, et al. 2024. StarCoder 2 and The Stack v2: The Next Generation. arXiv:2402.19173
- Lu et al. (2025) Chaomeng Lu, Tianyu Li, Toon Dehaene, and Bert Lagaisse. 2025. ICVul: A Well-labeled C/C++ Vulnerability Dataset with Comprehensive Metadata and VCCs. In International Conference on Mining Software Repositories (MSR). IEEE/ACM, Ottawa, ON, Canada, 154–158. doi:10.1109/MSR66628.2025.00034
- Lu et al. (2024) Guilong Lu, Xiaolin Ju, Xiang Chen, Wenlong Pei, and Zhilong Cai. 2024. GRACE: Empowering LLM-based Software Vulnerability Detection with Graph Structure and In-Context Learning. Journal of Systems and Software (JSS) 212 (2024), 112031.
- Lu et al. (2023) Guilong Lu, Xiaolin Ju, Xiang Chen, Shaoyu Yang, Liang Chen, and Hao Shen. 2023. Assessing the Effectiveness of Vulnerability Detection via Prompt Tuning: An Empirical Study. In Asia-Pacific Software Engineering Conference (APSEC). IEEE, Seoul, South Korea, 415–424. doi:10.1109/APSEC60848.2023.00052
- Lu et al. (2021) Shuai Lu, Daya Guo, Shuo Ren, Junjie Huang, Alexey Svyatkovskiy, Ambrosio Blanco, Colin Clement, Dawn Drain, Daxin Jiang, Duyu Tang, Ge Li, Lidong Zhou, Linjun Shou, Long Zhou, Michele Tufano, Ming Gong, Ming Zhou, Nan Duan, Neel Sundaresan, Shao Kun Deng, Shengyu Fu, and Shujie Liu. 2021. CodeXGLUE: A Machine Learning Benchmark Dataset for Code Understanding and Generation. arXiv:2102.04664
- Luo et al. (2021) Yu Luo, Weifeng Xu, and Dianxiang Xu. 2021. Detecting Integer Overflow Errors in Java Source Code via Machine Learning. In International Conference on Tools with Artificial Intelligence (ICTAI). IEEE, Washington, DC, USA, 724–728. doi:10.1109/ICTAI52525.2021.00115
- Luo et al. (2025) Yu Luo, Han Zhou, Mengtao Zhang, Dylan De La Rosa, Hafsa Ahmed, Weifeng Xu, and Dianxiang Xu. 2025. HALURust: Exploiting Hallucinations of Large Language Models to Detect Vulnerabilities in Rust. arXiv:2503.10793
- Mahmud et al. (2025) Adiba Mahmud, Yasmeen Rawajfih, and Fan Wu. 2025. An Ensemble Transformer Approach with Cross-Attention for Automated Code Security Vulnerability Detection and Documentation. In International Symposium on Digital Forensics and Security (ISDFS). IEEE, Boston, MA, USA, 1–6. doi:10.1109/ISDFS65363.2025.11012039
- Maligazhdarova et al. (2025) Nargiz Maligazhdarova, Avinash B. M., Assel Mukasheva, Didar Yedilkhan, Aidos Askhatuly, and A. Berdyshev. 2025. A Comparative Study of Machine Learning and Large Language Models for SQL and NoSQL Injection Vulnerability Detection. In IEEE International Conference on Smart Information Systems and Technologies (SIST). IEEE, Astana, Kazakhstan, 1–7. doi:10.1109/SIST61657.2025.11139190
- Mamede et al. (2022) Cláudia Mamede, Eduard Pinconschi, Rui Abreu, and José Campos. 2022. Exploring Transformers for Multi-Label Classification of Java Vulnerabilities. In International Conference on Software Quality, Reliability and Security (QRS). IEEE, Guangzhou, China, 43–52. doi:10.1109/QRS57517.2022.00015
- Mao et al. (2025) Qiheng Mao, Zhenhao Li, Xing Hu, Kui Liu, Xin Xia, and Jianling Sun. 2025. Towards Explainable Vulnerability Detection With Large Language Models. IEEE Transactions on Software Engineering (TSE) 51, 10 (2025), 2957–2971. doi:10.1109/TSE.2025.3605442
- Mathews et al. (2024) Noble Saji Mathews, Yelizaveta Brus, Yousra Aafer, Meiyappan Nagappan, and Shane McIntosh. 2024. LLbezpeky: Leveraging Large Language Models for Vulnerability Detection. arXiv:2401.01269
- Mechri et al. (2025) Abdechakour Mechri, Mohamed Amine Ferrag, and Merouane Debbah. 2025. SecureQwen: Leveraging LLMs for Vulnerability Detection in Python Codebases. Computers & Security (COSE) 148 (2025), 104151. doi:10.1016/j.cose.2024.104151
- Mei et al. (2023) Huan Mei, Guanjun Lin, Da Fang, and Jun Zhang. 2023. Detecting Vulnerabilities in IoT Software: New Hybrid Model and Comprehensive Data Analysis. Journal of Information Security and Applications (JISA) 74 (2023), 103467. doi:10.1016/j.jisa.2023.103467
- Melo et al. (2025) Rui Melo, Claudia Mamede, Andre Catarino, Rui Abreu, and Henrique Lopes Cardoso. 2025. Are Sparse Autoencoders Useful for Java Function Bug Detection? arXiv:2505.10375
- Meng et al. (2024) Fanxu Meng, Zhaohui Wang, and Muhan Zhang. 2024. Pissa: Principal Singular Values and Singular Vectors Adaptation of Large Language Models. In Advances in Neural Information Processing Systems (NeurIPS), Vol. 37. Curran Associates, Inc., Vancouver, BC, Canada, 121038–121072.
- Metrics and Evaluation (2006) NIST Software Assurance Metrics and Tool Evaluation. 2006. Software Assurance Reference Dataset. https://samate.nist.gov/SARD.
- Metrics and Evaluation (2017a) NIST Software Assurance Metrics and Tool Evaluation. 2017a. Juliet C/C++ 1.3. https://samate.nist.gov/SARD/test-suites/112.
- Metrics and Evaluation (2017b) NIST Software Assurance Metrics and Tool Evaluation. 2017b. Juliet Java 1.3. https://samate.nist.gov/SARD/test-suites/111.
- Mhatre et al. (2025) Akshay Mhatre, Noujoud Nader, Patrick Diehl, and Deepti Gupta. 2025. LLM-GUARD: Large Language Model-Based Detection and Repair of Bugs and Security Vulnerabilities in C++ and Python. arXiv:2508.16419
- MITRE (2023a) MITRE. 2023a. CWE-1000: Research Concepts. https://cwe.mitre.org/data/definitions/1000.html.
- MITRE (2023b) MITRE. 2023b. CWE CATEGORY: Resource Management Errors. https://cwe.mitre.org/data/definitions/399.html.
- MITRE (2025a) MITRE. 2025a. CWE - Common Weakness Enumeration. https://cwe.mitre.org/.
- MITRE (2025b) MITRE. 2025b. CWE-119: Improper Restriction of Operations within the Bounds of a Memory Buffer. https://cwe.mitre.org/data/definitions/119.html.
- MITRE (2025c) MITRE. 2025c. CWE-120: Buffer Copy without Checking Size of Input (’Classic Buffer Overflow’). https://cwe.mitre.org/data/definitions/120.html.
- Mock et al. (2025) Moritz Mock, Thomas Forrer, and Barbara Russo. 2025. Cross-Domain Evaluation of Transformer-Based Vulnerability Detection on Open and Industry Data. In International Conference on Product-Focused Software Process Improvement (PROFES). Springer, Salerno, Italy, 36–52. doi:10.1007/978-3-032-12092-2_3
- Mohajer et al. (2024) Mohammad Mahdi Mohajer, Reem Aleithan, Nima Shiri Harzevili, Moshi Wei, Alvine Boaye Belle, Hung Viet Pham, and Song Wang. 2024. Effectiveness of ChatGPT for Static Analysis: How Far Are We?. In International Conference on AI-Powered Software (AIware). ACM, Porto de Galinhas, Brazil, 151–160. doi:10.1145/3664646.3664777
- Moonen and Vidziunas (2024) Leon Moonen and Linas Vidziunas. 2024. CVEfixes Dataset: Automatically Collected Vulnerabilities and Their Fixes from Open-Source Software. https://zenodo.org/records/13118970.
- Munson et al. (2025) Amy Munson, Juanita Gomez, and Álvaro A. Cárdenas. 2025. ♪ With a Little Help from My (LLM) Friends: Enhancing Static Analysis with LLMs to Detect Software Vulnerabilities. In International Workshop on Large Language Models for Code (LLM4Code). IEEE/ACM, Ottawa, ON, Canada, 25–32. doi:10.1109/LLM4Code66737.2025.00008
- Mächtle et al. (2025) Felix Mächtle, Nils Loose, Tim Schulz, Florian Sieck, Jan-Niclas Serr, Ralf Möller, and Thomas Eisenbarth. 2025. Trace Gadgets: Minimizing Code Context for Machine Learning-Based Vulnerability Prediction. arXiv:2504.13676
- Ngan et al. (2025) Nguyen Dac Thien Ngan, Nghi Hoang Khoa, Van-Hau Pham, and Phan The Duy. 2025. Multimodal Fusion for Vulnerability Detection: Integrating Sequence and Graph-Based Analysis with LLM Augmentation. In International Conference on Multimedia Analysis and Pattern Recognition (MAPR). IEEE, Nha Trang, Vietnam, 1–6. doi:10.1109/MAPR67746.2025.11133833
- Nguyen et al. (2025a) Hong Quy Nguyen, Thong Hoang, Hoa Khanh Dam, Guoxin Su, Zhenchang Xing, Qinghua Lu, and Jiamou Sun. 2025a. Human-Understandable Explanation for Software Vulnerability Prediction. Journal of Systems and Software (JSS) 228 (2025), 112455. doi:10.1016/j.jss.2025.112455
- Nguyen et al. (2025c) Van Nguyen, Surya Nepal, Xingliang Yuan, Tingmin Wu, Fengchao Chen, and Carsten Rudolph. 2025c. MulVuln: Enhancing Pre-trained LMs with Shared and Language-Specific Knowledge for Multilingual Vulnerability Detection. arXiv:2510.04397
- Nguyen et al. (2025b) Van Nguyen, Surya Nepal, Xingliang Yuan, Tingmin Wu, and Carsten Rudolph. 2025b. SAFE: A Novel Approach For Software Vulnerability Detection from Enhancing The Capability of Large Language Models. In ACM Asia Conference on Computer and Communications Security (ASIACCS). ACM, Ha Noi, Vietnam, 392–406. doi:10.1145/3708821.3736208
- Ni et al. (2023a) Chao Ni, Xinrong Guo, Yan Zhu, Xiaodan Xu, and Xiaohu Yang. 2023a. Function-Level Vulnerability Detection Through Fusing Multi-Modal Knowledge. In International Conference on Automated Software Engineering (ASE). IEEE/ACM, Kirchberg, Luxembourg, 1911–1918. doi:10.1109/ASE56229.2023.00084
- Ni et al. (2024) Chao Ni, Liyu Shen, Xiaohu Yang, Yan Zhu, and Shaohua Wang. 2024. MegaVul: A C/C++ Vulnerability Dataset with Comprehensive Code Representations. In International Conference on Mining Software Repositories (MSR). ACM, Lisbon, Portugal, 738–742. doi:10.1145/3643991.3644886
- Ni et al. (2025) Chao Ni, Xin Yin, Liyu Shen, and Shaohua Wang. 2025. Learning-based Models for Vulnerability Detection: An Extensive Study. Empirical Software Engineering (EMSE) 31, 18 (2025), 38. doi:10.1007/s10664-025-10734-x
- Ni et al. (2023b) Chao Ni, Xin Yin, Kaiwen Yang, Dehai Zhao, Zhenchang Xing, and Xin Xia. 2023b. Distinguishing Look-Alike Innocent and Vulnerable Code by Subtle Semantic Representation Learning and Explanation. In Joint European Software Engineering Conference and Symposium on the Foundations of Software Engineering (ESEC/FSE). ACM, San Francisco, CA, USA, 1611–1622. doi:10.1145/3611643.3616358
- Nikitopoulos et al. (2021a) Georgios Nikitopoulos, Konstantina Dritsa, Panos Louridas, and Dimitris Mitropoulos. 2021a. Cross-Language Vulnerability Dataset with File Changes and Commit Messages. https://zenodo.org/records/4734050.
- Nikitopoulos et al. (2021b) Georgios Nikitopoulos, Konstantina Dritsa, Panos Louridas, and Dimitris Mitropoulos. 2021b. CrossVul: A Cross-Language Vulnerability Dataset with Commit Data. In Joint Meeting on European Software Engineering Conference and Symposium on the Foundations of Software Engineering (ESEC/FSE). ACM, Athens, Greece, 1565–1569. doi:10.1145/3468264.3473122
- Noever (2023) David Noever. 2023. Can Large Language Models Find And Fix Vulnerable Software? arXiv:2308.10345
- Nong et al. (2024) Yu Nong, Mohammed Aldeen, Long Cheng, Hongxin Hu, Feng Chen, and Haipeng Cai. 2024. Chain-of-Thought Prompting of Large Language Models for Discovering and Fixing Software Vulnerabilities. arXiv:2402.17230
- Nong et al. (2023) Yu Nong, Rainy Sharma, Abdelwahab Hamou-Lhadj, Xiapu Luo, and Haipeng Cai. 2023. Open Science in Software Engineering: A Study on Deep Learning-Based Vulnerability Detection. IEEE Transactions on Software Engineering (TSE) 49, 4 (2023), 1983–2005. doi:10.1109/TSE.2022.3207149
- NVD (2025) NVD. 2025. National Vulnerability Database (NVD). https://nvd.nist.gov/.
- Omar (2023) Marwan Omar. 2023. VulDefend: A Novel Technique based on Pattern-exploiting Training for Detecting Software Vulnerabilities Using Language Models. In Jordan International Joint Conference on Electrical Engineering and Information Technology (JEEIT). IEEE, Amman, Jordan, 287–293. doi:10.1109/JEEIT58638.2023.10185860
- Omar and Shiaeles (2023) Marwan Omar and Stavros Shiaeles. 2023. VulDetect: A novel technique for detecting software vulnerabilities using Language Models. In International Conference on Cyber Security and Resilience (CSR). IEEE, Venice, Italy, 105–110. doi:10.1109/CSR57506.2023.10224924
- OpenAI et al. (2024) OpenAI, Josh Achiam, Steven Adler, Sandhini Agarwal, Lama Ahmad, Ilge Akkaya, Florencia Leoni Aleman, Diogo Almeida, Janko Altenschmidt, Sam Altman, et al. 2024. GPT-4 Technical Report. arXiv:2303.08774
- Ouchebara and Dupont (2025) Dyna Soumhane Ouchebara and Stéphane Dupont. 2025. Llama-Based Source Code Vulnerability Detection: Prompt Engineering vs Fine Tuning. In European Symposium on Research in Computer Security (ESORICS). Springer, Toulouse, France, 289–308. doi:10.1007/978-3-032-07884-1_15
- Ozturk et al. (2023) Omer Said Ozturk, Emre Ekmekcioglu, Orcun Cetin, Budi Arief, and Julio Hernandez-Castro. 2023. New Tricks to Old Codes: Can AI Chatbots Replace Static Code Analysis Tools?. In European Interdisciplinary Cybersecurity Conference (EICC). ACM, Stavanger, Norway, 13–18. doi:10.1145/3590777.3590780
- Pan et al. (2024) Shirui Pan, Linhao Luo, Yufei Wang, Chen Chen, Jiapu Wang, and Xindong Wu. 2024. Unifying Large Language Models and Knowledge Graphs: A Roadmap. IEEE Transactions on Knowledge and Data Engineering (TKDE) 36, 7 (2024), 3580–3599. doi:10.1109/TKDE.2024.3352100
- Panichella (2025) Annibale Panichella. 2025. Metamorphic-Based Many-Objective Distillation of LLMs for Code-Related Tasks. In International Conference on Software Engineering (ICSE). IEEE/ACM, Ottawa, ON, Canada, 1001–1013. doi:10.1109/ICSE55347.2025.00230
- Pearce et al. (2025) Hammond Pearce, Baleegh Ahmad, Benjamin Tan, Brendan Dolan-Gavitt, and Ramesh Karri. 2025. Asleep at the Keyboard? Assessing the Security of GitHub Copilot’s Code Contributions. Commun. ACM 68, 2 (2025), 96–105. doi:10.1145/3610721
- Pelofske et al. (2024) Elijah Pelofske, Vincent Urias, and Lorie M. Liebrock. 2024. Automated Software Vulnerability Static Code Analysis Using Generative Pre-Trained Transformer Models. arXiv:2408.00197
- Peng et al. (2023) Tao Peng, Shixu Chen, Fei Zhu, Junwei Tang, Junping Liu, and Xinrong Hu. 2023. PTLVD:Program Slicing and Transformer-based Line-level Vulnerability Detection System. In International Working Conference on Source Code Analysis and Manipulation (SCAM). IEEE, Bogotá, Colombia, 162–173. doi:10.1109/SCAM59687.2023.00026
- Peng et al. (2025) Yun Peng, Kisub Kim, Linghan Meng, and Kui Liu. 2025. iCodeReviewer: Improving Secure Code Review with Mixture of Prompts. arXiv:2510.12186
- Petersen et al. (2015) Kai Petersen, Sairam Vakkalanka, and Ludwik Kuzniarz. 2015. Guidelines for Conducting Systematic Mapping Studies in Software Engineering: An Update. Information and Software Technology (INFSOF) 64 (2015), 1–18. doi:10.1016/j.infsof.2015.03.007
- Pham et al. (2024) Van-Hau Pham, Do Thi Thu Hien, Hien Do Hoang, and Phan The Duy. 2024. Defect-Scanner: A Comparative Empirical Study on Language Model and Deep Learning Approach for Software Vulnerability Detection. International Journal of Information Security (IJIS) 23 (2024), 1–14. Issue 6. doi:10.1007/s10207-024-00901-4
- Pinconschi (2020) Eduard Pinconschi. 2020. Devign. https://github.com/epicosy/devign.
- Prabith et al. (2024) GS Prabith, Rohit Narayanan, A Arya, PK Binu, et al. 2024. BiT5: A Bidirectional NLP Approach for Advanced Vulnerability Detection in Codebase. Procedia Computer Science 233 (2024), 812–821.
- Prasad et al. (2025) Rupesh Prasad, Lokesh B, Swathika Nivedini D, Piyush Pratap Singh, Arvind Maurya, and Atul Singh. 2025. Sink Vulnerability Type Prediction Using Small Language Model (SLM). In International Conference on Cognitive Computing in Engineering, Communications, Sciences and Biomedical Health Informatics (IC3ECSBHI). IEEE, Greater Noida, India, 496–501. doi:10.1109/IC3ECSBHI63591.2025.10991300
- Purba et al. (2023) Moumita Das Purba, Arpita Ghosh, Benjamin J. Radford, and Bill Chu. 2023. Software Vulnerability Detection using Large Language Models. In International Symposium on Software Reliability Engineering Workshops (ISSREW). IEEE, Florence, Italy, 112–119. doi:10.1109/ISSREW60843.2023.00058
- Qi et al. (2024) Weiliang Qi, Jiahao Cao, Darsh Poddar, Sophia Li, and Xinda Wang. 2024. Software Vulnerability Detection using Large Language Models. In International Conference on Security and Privacy in Communication Systems (SecureComm). Springer, Dubai, United Arab Emirates, 184–203. doi:10.1007/978-3-031-94458-1_9
- Qian et al. (2025) Zhuoyun Qian, Fangtian Zhong, Qin Hu, Yili Jiang, Jiaqi Huang, Mengfei Ren, and Jiguo Yu. 2025. Software Vulnerability Analysis Across Programming Language and Program Representation Landscapes: A Survey. arXiv:2503.20244
- Quan et al. (2025) Vu Le Anh Quan, Chau Thuan Phat, Kiet Van Nguyen, Phan The Duy, and Van-Hau Pham. 2025. XGV-BERT: Leveraging Contextualized Language Model and Graph Neural Network for Efficient Software Vulnerability Detection. The Journal of Supercomputing 81, 6 (2025), 750. doi:10.1007/s11227-025-07198-7
- Qwen et al. (2025) Qwen, An Yang, Baosong Yang, Beichen Zhang, Binyuan Hui, Bo Zheng, Bowen Yu, Chengyuan Li, Dayiheng Liu, Fei Huang, et al. 2025. Qwen2.5 Technical Report. arXiv:2412.15115
- Radford et al. (2019) Alec Radford, Jeffrey Wu, Rewon Child, David Luan, Dario Amodei, Ilya Sutskever, et al. 2019. Language Models are Unsupervised Multitask Learners. OpenAI Blog 1, 8 (2019), 9.
- Raffel et al. (2020) Colin Raffel, Noam Shazeer, Adam Roberts, Katherine Lee, Sharan Narang, Michael Matena, Yanqi Zhou, Wei Li, and Peter J Liu. 2020. Exploring the Limits of Transfer Learning with a Unified Text-to-text Transformer. Journal of Machine Learning Research (JMLR) 21, 140 (2020), 1–67.
- Rahman et al. (2024) Md Mahbubur Rahman, Ira Ceka, Chengzhi Mao, Saikat Chakraborty, Baishakhi Ray, and Wei Le. 2024. Towards Causal Deep Learning for Vulnerability Detection. In International Conference on Software Engineering (ICSE). IEEE/ACM, Lisbon, Portugal, Article 153, 11 pages. doi:10.1145/3597503.3639170
- Rangapuram and Ratha (2025) Deekshith Sagar Rangapuram and Nalini Ratha. 2025. GraphCodeBERT-Augmented Graph Attention Networks for Code Vulnerability Detection. In Conference on Artificial Intelligence (CAI). IEEE, Santa Clara, CA, USA, 912–917. doi:10.1109/CAI64502.2025.00161
- Ren et al. (2024) Zilong Ren, Xiaolin Ju, Xiang Chen, and Hao Shen. 2024. ProRLearn: Boosting Prompt Tuning-based Vulnerability Detection by Reinforcement Learning. Automated Software Engineering (ASE) 31, 2 (2024), 38. doi:10.1007/s10515-024-00438-9
- Ridley et al. (2023) Norah Ridley, Enrico Branca, Jadyn Kimber, and Natalia Stakhanova. 2023. Enhancing Code Security Through Open-source Large Language Models: A Comparative Study. In International Symposium on Foundations and Practice of Security (FPS). Springer, Bordeaux, France, 233–249. doi:10.1007/978-3-031-57537-2_15
- Ridoy et al. (2024) Shahriyar Zaman Ridoy, Md. Shazzad Hossain Shaon, Alfredo Cuzzocrea, and Mst Shapna Akter. 2024. EnStack: An Ensemble Stacking Framework of Large Language Models for Enhanced Vulnerability Detection in Source Code. In International Conference on Big Data (BigData). IEEE, Washington DC, USA, 6356–6364. doi:10.1109/BigData62323.2024.10825609
- Risse and B”̈ohme (2024) Niklas Risse and Marcel B”̈ohme. 2024. Uncovering the Limits of Machine Learning for Automatic Vulnerability Detection. In USENIX Security Symposium (USENIX Security). USENIX, Philadelphia, PA, USA, 4247–4264.
- Risse et al. (2025) Niklas Risse, Jing Liu, and Marcel Böhme. 2025. Top Score on the Wrong Exam: On Benchmarking in Machine Learning for Vulnerability Detection. Proceedings of the ACM on Software Engineering (ASE) 2, ISSTA, Article ISSTA018 (2025), 23 pages. doi:10.1145/3728887
- Rozière et al. (2024) Baptiste Rozière, Jonas Gehring, Fabian Gloeckle, Sten Sootla, Itai Gat, Xiaoqing Ellen Tan, Yossi Adi, Jingyu Liu, Romain Sauvestre, Tal Remez, et al. 2024. Code Llama: Open Foundation Models for Code. arXiv:2308.12950
- Ruohonen (2021) Jukka Ruohonen. 2021. The Similarities of Software Vulnerabilities for Interpreted Programming Languages. In IEEE International Conference on Progress in Informatics and Computing (PIC). IEEE, Shanghai, China, 304–307. doi:10.1109/PIC53636.2021.9687053
- Russell et al. (2018) Rebecca Russell, Louis Kim, Lei Hamilton, Tomo Lazovich, Jacob Harer, Onur Ozdemir, Paul Ellingwood, and Marc McConley. 2018. Automated Vulnerability Detection in Source Code Using Deep Representation Learning. In International Conference on Machine Learning and Applications (ICMLA). IEEE, Orlando, FL, USA, 757–762. doi:10.1109/ICMLA.2018.00120
- Russo et al. (2025) Barbara Russo, Jorge Melegati, and Moritz Mock. 2025. Leveraging Multi-Task Learning to Improve the Detection of SATD and Vulnerability. In International Conference on Program Comprehension (ICPC). IEEE/ACM, Ottawa, ON, Canada, 1–12. doi:10.1109/ICPC66645.2025.00017
- Safdar et al. (2025b) Rijha Safdar, Danyail Mateen, Syed Taha Ali, M. Umer Ashfaq, and Wajahat Hussain. 2025b. Data and Context Matter: Towards Generalizing AI-based Software Vulnerability Detection. arXiv:2508.16625
- Safdar et al. (2025a) Rijha Safdar, Danyail Mateen, Syed Taha Ali, and Wajahat Hussain. 2025a. Real-VulLLM: An LLM Based Assessment Framework in the Wild. arXiv:2510.04056
- Saimbhi and Akpinar (2024) Sohan Singh Saimbhi and Kevser Ovaz Akpinar. 2024. VulnerAI: GPT Based Web Application Vulnerability Detection. In International Conference on Artificial Intelligence, Metaverse and Cybersecurity (ICAMAC). IEEE, Dubai, UAE, 1–6. doi:10.1109/ICAMAC62387.2024.10828788
- Sanh et al. (2020) Victor Sanh, Lysandre Debut, Julien Chaumond, and Thomas Wolf. 2020. DistilBERT, a distilled version of BERT: smaller, faster, cheaper and lighter. arXiv:1910.01108
- Sayagh and Ghafari (2025) Mohammed Sayagh and Mohammad Ghafari. 2025. Think Broad, Act Narrow: CWE Identification with Multi-Agent Large Language Models. arXiv:2508.01451
- Security & Software Engineering Research Lab at University of Notre Dame (2022) Security & Software Engineering Research Lab at University of Notre Dame. 2022. SecurityEval. https://github.com/s2e-lab/SecurityEval.
- Shahzad et al. (2025) Murtuza Shahzad, Joseph Wilson, Ibrahim Al Azher, Hamed Alhoori, and Mona Rahimi. 2025. From Theory to Practice: Code Generation Using LLMs for CAPEC and CWE Frameworks. In International Workshop on Large Language Models for Code (LLM4Code). IEEE/ACM, Ottawa, ON, Canada, 137–144. doi:10.1109/LLM4Code66737.2025.00022
- Shashwat et al. (2024) K Shashwat, F Hahn, X Ou, D Goldgof, J Ligatti, L Hall, R Rajagopalan, and A Tabari. 2024. A Preliminary Study on Using Large Language Models in Software Pentesting. In Network and Distributed System Security (NDSS) Symposium: Workshop on SOC Operations and Construction. NDSS Symposium, Marina del Rey, CA, USA.
- Sheng et al. (2025) Ze Sheng, Zhicheng Chen, Shuning Gu, Heqing Huang, Guofei Gu, and Jeff Huang. 2025. LLMs in Software Security: A Survey of Vulnerability Detection Techniques and Insights. Comput. Surveys 58, 5, Article 134 (2025), 35 pages. doi:10.1145/3769082
- Sheng et al. (2024) Ze Sheng, Fenghua Wu, Xiangwu Zuo, Chao Li, Yuxin Qiao, and Hang Lei. 2024. Research on the LLM-Driven Vulnerability Detection System Using LProtector. In International Conference on Data Science and Computer Application (ICDSCA). IEEE, Ottawa, ON, Canada, 192–196. doi:10.1109/ICDSCA63855.2024.10859408
- Shereen et al. (2024) Ezzeldin Shereen, Dan Ristea, Sanyam Vyas, Shae McFadden, Madeleine Dwyer, Chris Hicks, and Vasilios Mavroudis. 2024. SoK: On Closing the Applicability Gap in Automated Vulnerability Detection. arXiv:2412.11194
- Shestov et al. (2025) Aleksei Shestov, Rodion Levichev, Ravil Mussabayev, Evgeny Maslov, Pavel Zadorozhny, Anton Cheshkov, Rustam Mussabayev, Alymzhan Toleu, Gulmira Tolegen, and Alexander Krassovitskiy. 2025. Finetuning Large Language Models for Vulnerability Detection. IEEE Access 13 (2025), 38889–38900. doi:10.1109/ACCESS.2025.3546700
- Shi et al. (2024) Jieke Shi, Zhou Yang, Hong Jin Kang, Bowen Xu, Junda He, and David Lo. 2024. Greening Large Language Models of Code. In International Conference on Software Engineering: Software Engineering in Society (ICSE-SEIS). ACM, Lisbon, Portugal, 142–153. doi:10.1145/3639475.3640097
- Shimmi et al. (2025) Samiha Shimmi, Hamed Okhravi, and Mona Rahimi. 2025. AI-Based Software Vulnerability Detection: A Systematic Literature Review. arXiv:2506.10280
- Shimmi et al. (2024a) Samiha Shimmi, Ashiqur Rahman, Mohan Gadde, Hamed Okhravi, and Mona Rahimi. 2024a. VulSim: Leveraging Similarity of Multi-Dimensional Neighbor Embeddings for Vulnerability Detection. In USENIX Security Symposium (USENIX Security). USENIX, Philadelphia, PA, USA, 1777–1794.
- Shimmi et al. (2024b) Samiha Shimmi, Yash Saini, Mark Schaefer, Hamed Okhravi, and Mona Rahimi. 2024b. Software Vulnerability Detection Using LLM: Does Additional Information Help?. In Annual Computer Security Applications Conference Workshops (ACSAC Workshops). IEEE, Honolulu, Hawaii, USA, 216–223. doi:10.1109/ACSACW65225.2024.00031
- Shiri Harzevili et al. (2024) Nima Shiri Harzevili, Alvine Boaye Belle, Junjie Wang, Song Wang, Zhen Ming (Jack) Jiang, and Nachiappan Nagappan. 2024. A Systematic Literature Review on Automated Software Vulnerability Detection Using Machine Learning. Comput. Surveys 57, 3, Article 55 (2024), 36 pages. doi:10.1145/3699711
- Shu et al. (2025) Honglin Shu, Michael Fu, Junji Yu, Dong Wang, Chakkrit Tantithamthavorn, Junjie Chen, and Yasutaka Kamei. 2025. Large Language Models for Multilingual Vulnerability Detection: How Far Are We? arXiv:2506.07503
- Siddiq and Santos (2022) Mohammed Latif Siddiq and Joanna C. S. Santos. 2022. SecurityEval Dataset: Mining Vulnerability Examples to Evaluate Machine Learning-Based Code Generation Techniques. In International Workshop on Mining Software Repositories Applications for Privacy and Security (MSR4P&S). ACM, Singapore, 29–33. doi:10.1145/3549035.3561184
- Simoni et al. (2025) Marco Simoni, Aleksandar Fontana, Giulio Rossolini, and Andrea Saracino. 2025. Improving LLM Reasoning for Vulnerability Detection via Group Relative Policy Optimization. arXiv:2507.03051
- Singh et al. (2022) Kanchan Singh, Sakshi S Grover, and Ranjini Kishen Kumar. 2022. Cyber Security Vulnerability Detection Using Natural Language Processing. In World AI IoT Congress (AIIoT). IEEE, Seattle, WA, USA, 174–178. doi:10.1109/AIIoT54504.2022.9817336
- Sovrano et al. (2025) Francesco Sovrano, Adam Bauer, and Alberto Bacchelli. 2025. Large Language Models for In-File Vulnerability Localization Can Be ””Lost in the End””. Proceedings of the ACM on Software Engineering (PACMSE) 2, FSE, Article FSE041 (2025), 23 pages. doi:10.1145/3715758
- SRI Lab, ETH Zurich (2023) SRI Lab, ETH Zurich. 2023. SVEN: Security Hardening and Adversarial Testing for Code LLMs. https://github.com/eth-sri/sven.
- Steenhoek et al. (2023a) Benjamin Steenhoek, Md Mahbubur Rahman, Richard Jiles, and Wei Le. 2023a. An Empirical Study of Deep Learning Models for Vulnerability Detection. In International Conference on Software Engineering (ICSE). IEEE/ACM, Melbourne, Australia, 2237–2248. doi:10.1109/ICSE48619.2023.00188
- Steenhoek et al. (2025a) Benjamin Steenhoek, Md Mahbubur Rahman, Monoshi Kumar Roy, Mirza Sanjida Alam, Hengbo Tong, Swarna Das, Earl T. Barr, and Wei Le. 2025a. To Err is Machine: Vulnerability Detection Challenges LLM Reasoning. arXiv:2403.17218
- Steenhoek et al. (2023b) Benjamin Steenhoek, Md Mahbubur Rahman, Shaila Sharmin, and Wei Le. 2023b. Do Language Models Learn Semantics of Code? A Case Study in Vulnerability Detection. arXiv:2311.04109
- Steenhoek et al. (2025b) Benjamin Steenhoek, Kalpathy Sivaraman, Renata Saldivar Gonzalez, Yevhen Mohylevskyy, Roshanak Zilouchian Moghaddam, and Wei Le. 2025b. Closing the Gap: A User Study on the Real-world Usefulness of AI-powered Vulnerability Detection & Repair in the IDE. In International Conference on Software Engineering (ICSE). IEEE/ACM, Ottawa, ON, Canada, 1–13. doi:10.1109/ICSE55347.2025.00126
- Su and Wu (2023) Jingyi Su and Yan Wu. 2023. Optimizing Pre-trained Language Models for Efficient Vulnerability Detection in Code Snippets. In International Conference on Computer and Communications (ICCC). IEEE, Chengdu, China, 2139–2143. doi:10.1109/ICCC59590.2023.10507456
- Sultan et al. (2024) Md. Fahim Sultan, Tasmin Karim, Md. Shazzad Hossain Shaon, Mohammad Wardat, and Mst Shapna Akter. 2024. Enhanced LLM-Based Framework for Predicting Null Pointer Dereference in Source Code. arXiv:2412.00216
- Sultana et al. (2024) Shaznin Sultana, Sadia Afreen, and Nasir U. Eisty. 2024. Code Vulnerability Detection: A Comparative Analysis of Emerging Large Language Models. arXiv:2409.10490
- Sun et al. (2023) Hao Sun, Yongji Liu, Zhenquan Ding, Yang Xiao, Zhiyu Hao, and Hongsong Zhu. 2023. An Enhanced Vulnerability Detection in Software Using a Heterogeneous Encoding Ensemble. In Symposium on Computers and Communications (ISCC). IEEE, Tunis, Tunisia, 1214–1220. doi:10.1109/ISCC58397.2023.10217978
- Sun et al. (2025c) Xiaobing Sun, Mingxuan Zhou, Sicong Cao, Xiaoxue Wu, Lili Bo, Di Wu, Bin Li, and Yang Xiang. 2025c. HgtJIT: Just-in-Time Vulnerability Detection Based on Heterogeneous Graph Transformer. IEEE Transactions on Dependable and Secure Computing (TDSC) 22, 6 (2025), 1–17. doi:10.1109/TDSC.2025.3586669
- Sun et al. (2025b) Yuqiang Sun, Daoyuan Wu, Yue Xue, Han Liu, Wei Ma, Lyuye Zhang, Yang Liu, and Yingjiu Li. 2025b. LLM4Vuln: A Unified Evaluation Framework for Decoupling and Enhancing LLMs’ Vulnerability Reasoning. arXiv:2401.16185
- Sun et al. (2025a) Zhihong Sun, Jia Li, Yao Wan, Chuanyi Li, Hongyu Zhang, Zhi jin, Ge Li, Hong Liu, Chen Lyu, and Songlin Hu. 2025a. Ensembling Large Language Models for Code Vulnerability Detection: An Empirical Evaluation. arXiv:2509.12629
- Sun et al. (2024) Zhequ Sun, Ke Liu, and Yu Yang. 2024. Enhancing Source Code Vulnerability Detection Using Flattened Code Graph Structures. In International Conference on Frontier Technologies of Information and Computer (ICFTIC). IEEE, Qingdao, China, 209–213. doi:10.1109/ICFTIC64248.2024.10913325
- Taghavi Far and Feyzi (2025) Seyed Mohammad Taghavi Far and Farid Feyzi. 2025. Large Language Models for Software Vulnerability Detection: A Guide for Researchers on Models, Methods, Techniques, Datasets, and Metrics. International Journal of Information Security (IJIS) 24, 2 (2025), 78. doi:10.1007/s10207-025-00992-7
- Takieldeen et al. (2025) Ali E. Takieldeen, Merna Gamal, Abdelrahman Ahmed Salah, Mohamed Walid, Abdelrahman Mohamed Mahmoud, khaled Ahmed, Saleh Tamer Abdallah, Nada Samir Elsayed, and Aida A. Nasr. 2025. AI-Powered Vulnerability Detection in Code Using BERT-Based LLM with Transparency Measures. In International Telecommunications Conference (ITC-Egypt). IEEE, Cairo, Egypt, 297–304. doi:10.1109/ITC-Egypt66095.2025.11186618
- Tamberg and Bahsi (2025) Karl Tamberg and Hayretdin Bahsi. 2025. Harnessing Large Language Models for Software Vulnerability Detection: A Comprehensive Benchmarking Study. IEEE Access 13 (2025), 29698–29717. doi:10.1109/ACCESS.2025.3541146
- Team et al. (2024d) CodeGemma Team, Heri Zhao, Jeffrey Hui, Joshua Howland, Nam Nguyen, Siqi Zuo, Andrea Hu, Christopher A. Choquette-Choo, Jingyue Shen, Joe Kelley, et al. 2024d. CodeGemma: Open Code Models Based on Gemma. arXiv:2406.11409
- Team et al. (2024a) Gemini Team, Petko Georgiev, Ving Ian Lei, Ryan Burnell, Libin Bai, Anmol Gulati, Garrett Tanzer, Damien Vincent, Zhufeng Pan, Shibo Wang, et al. 2024a. Gemini 1.5: Unlocking multimodal understanding across millions of tokens of context. arXiv:2403.05530
- Team et al. (2024b) Gemma Team, Thomas Mesnard, Cassidy Hardin, Robert Dadashi, Surya Bhupatiraju, Shreya Pathak, Laurent Sifre, Morgane Rivière, Mihir Sanjay Kale, Juliette Love, et al. 2024b. Gemma: Open Models Based on Gemini Research and Technology. arXiv:2403.08295
- Team et al. (2024c) Gemma Team, Morgane Riviere, Shreya Pathak, Pier Giuseppe Sessa, Cassidy Hardin, Surya Bhupatiraju, Léonard Hussenot, Thomas Mesnard, Bobak Shahriari, Alexandre Ramé, et al. 2024c. Gemma 2: Improving Open Language Models at a Practical Size. arXiv:2408.00118
- Thapa et al. (2022) Chandra Thapa, Seung Ick Jang, Muhammad Ejaz Ahmed, Seyit Camtepe, Josef Pieprzyk, and Surya Nepal. 2022. Transformer-Based Language Models for Software Vulnerability Detection. In Annual Computer Security Applications Conference (ACSAC). ACM, Austin, TX, USA, 481–496. doi:10.1145/3564625.3567985
- Tian et al. (2025) Bowen Tian, Zhengyang Xu, Mingqiang Wu, Songning Lai, and Yutai Yue. 2025. You Only Train Once: A Flexible Training Framework for Code Vulnerability Detection Driven by Vul-Vector. arXiv:2506.10988
- Tian and Zhang (2025) Lei Tian and Cheng Zhang. 2025. EFVD: A Framework of Source Code Vulnerability Detection via Fusion of Enhanced Graph Representation Learning and Pre-trained Transformer-Based Model. In International Conference on Computer Network Security and Software Engineering (CNSSE). ACM, Qingdao, China, 316–320. doi:10.1145/3732365.3732421
- Tian et al. (2024) Zonghang Tian, Shuting Lou, Yingying Zhang, Zhengdan Jiang, and Xingkai Cheng. 2024. SQL Injection Vulnerability Detection Based on Pissa-Tuned Llama 3 Large Language Model. In International Conference on Frontier Technologies of Information and Computer (ICFTIC). IEEE, Qingdao, China, 255–259. doi:10.1109/ICFTIC64248.2024.10912886
- Tihanyi et al. (2023) Norbert Tihanyi, Tamas Bisztray, Ridhi Jain, Mohamed Amine Ferrag, Lucas C. Cordeiro, and Vasileios Mavroeidis. 2023. The FormAI Dataset: Generative AI in Software Security through the Lens of Formal Verification. In International Conference on Predictive Models and Data Analytics in Software Engineering (PROMISE). ACM, San Francisco, CA, USA, 33–43. doi:10.1145/3617555.3617874
- Torkamani et al. (2025) Mohammad Jalili Torkamani, Joey NG, Nikita Mehrotra, Mahinthan Chandramohan, Padmanabhan Krishnan, and Rahul Purandare. 2025. Streamlining Security Vulnerability Triage with Large Language Models. arXiv:2501.18908
- Touvron et al. (2023a) Hugo Touvron, Thibaut Lavril, Gautier Izacard, Xavier Martinet, Marie-Anne Lachaux, Timothée Lacroix, Baptiste Rozière, Naman Goyal, Eric Hambro, Faisal Azhar, et al. 2023a. LLaMA: Open and Efficient Foundation Language Models. arXiv:2302.13971
- Touvron et al. (2023b) Hugo Touvron, Louis Martin, Kevin Stone, Peter Albert, Amjad Almahairi, Yasmine Babaei, Nikolay Bashlykov, Soumya Batra, Prajjwal Bhargava, Shruti Bhosale, et al. 2023b. Llama 2: Open Foundation and Fine-Tuned Chat Models. arXiv:2307.09288
- Trad and Chehab (2025) Fouad Trad and Ali Chehab. 2025. Manual Prompt Engineering is Not Dead: A Case Study on Large Language Models for Code Vulnerability Detection with DSPy. In International Conference on Data Science and Machine Learning Applications (CDMA). IEEE, Riyadh, Saudi Arabia, 168–173. doi:10.1109/CDMA61895.2025.00034
- Tsai et al. (2025a) Chung-Nan Tsai, Xin Wang, Cheng-Hsiung Lee, and Ching-Sheng Lin. 2025a. A Sequential Multi-Stage Approach for Code Vulnerability Detection via Confidence- and Collaboration-based Decision Making. In Conference on Empirical Methods in Natural Language Processing (EMNLP). Association for Computational Linguistics, Suzhou, China, 21162–21168. doi:10.18653/v1/2025.emnlp-main.1071
- Tsai et al. (2025b) Chung-Nan Tsai, Jingnan Xie, Chun-Ming Lai, and Ching-Sheng Lin. 2025b. Leveraging Intra-and Inter-References in Vulnerability Detection using Multi-Agent Collaboration Based on LLMs. Cluster Computing 28, 15 (2025), 1–12. doi:10.1007/s10586-025-05721-2
- Ullah et al. (2024) Saad Ullah, Mingji Han, Saurabh Pujar, Hammond Pearce, Ayse Coskun, and Gianluca Stringhini. 2024. LLMs Cannot Reliably Identify and Reason About Security Vulnerabilities (Yet?): A Comprehensive Evaluation, Framework, and Benchmarks. In Symposium on Security and Privacy (SP). IEEE, San Francisco, CA, USA, 862–880. doi:10.1109/SP54263.2024.00210
- Vaswani et al. (2017) Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones, Aidan N Gomez, Ł ukasz Kaiser, and Illia Polosukhin. 2017. Attention is All you Need. In Advances in Neural Information Processing Systems (NeurIPS), Vol. 30. Curran Associates, Inc., Long Beach, CA, USA, 5998–6008.
- VulDetProject (2020) VulDetProject. 2020. Deep Learning based Vulnerability Detection:Are We There Yet? https://github.com/VulDetProject/ReVeal.
- Wagner Research Group (2025) Wagner Research Group. 2025. DiverseVul: A New Vulnerable Source Code Dataset for Deep Learning Based Vulnerability Detection. https://github.com/wagner-group/diversevul.
- Wang et al. (2025b) Huanting Wang, Dejice Jacob, David Kelly, Yehia Elkhatib, Jeremy Singer, and Zheng Wang. 2025b. SecureMind: A Framework for Benchmarking Large Language Models in Memory Bug Detection and Repair. In International Symposium on Memory Management (ISMM). ACM, Seoul, South Korea, 27–40. doi:10.1145/3735950.3735954
- Wang et al. (2023a) Jiexin Wang, Liuwen Cao, Xitong Luo, Zhiping Zhou, Jiayuan Xie, Adam Jatowt, and Yi Cai. 2023a. Enhancing Large Language Models for Secure Code Generation: A Dataset-driven Study on Vulnerability Mitigation. arXiv:2310.16263
- Wang et al. (2023b) Jin Wang, Zishan Huang, Hengli Liu, Nianyi Yang, and Yinhao Xiao. 2023b. DefectHunter: A Novel LLM-Driven Boosted-Conformer-based Code Vulnerability Detection Mechanism. arXiv:2309.15324
- Wang et al. (2024b) Rongcun Wang, Senlei Xu, Yuan Tian, Xingyu Ji, Xiaobing Sun, and Shujuang Jiang. 2024b. SCL-CVD: Supervised Contrastive Learning for Code Vulnerability Detection via GraphCodeBERT. Computers & Security (COSE) 145 (2024), 103994. doi:10.1016/j.cose.2024.103994
- Wang et al. (2025g) Weizhou Wang, Eric Liu, Xiangyu Guo, Xiao Hu, Ilya Grishchenko, and David Lie. 2025g. ANVIL: Anomaly-based Vulnerability Identification without Labelled Training Data. arXiv:2408.16028
- Wang (2024) Xinchen Wang. 2024. [ICSE 2024 Industry Challenge Track] Official Implementation of ”ReposVul: A Repository-Level High-Quality Vulnerability Dataset”. https://github.com/Eshe0922/ReposVul.
- Wang et al. (2024a) Xinchen Wang, Ruida Hu, Cuiyun Gao, Xin-Cheng Wen, Yujia Chen, and Qing Liao. 2024a. ReposVul: A Repository-Level High-Quality Vulnerability Dataset. In International Conference on Software Engineering: Companion Proceedings (ICSE-Companion). IEEE/ACM, Lisbon, Portugal, 472–483. doi:10.1145/3639478.3647634
- Wang et al. (2023d) Xuezhi Wang, Jason Wei, Dale Schuurmans, Quoc Le, Ed Chi, Sharan Narang, Aakanksha Chowdhery, and Denny Zhou. 2023d. Self-Consistency Improves Chain of Thought Reasoning in Language Models. arXiv:2203.11171
- Wang et al. (2025a) Yuxuan Wang, Jingshu Chen, and Qingyang Wang. 2025a. Leveraging Large Language Models for Command Injection Vulnerability Analysis in Python: An Empirical Study on Popular Open-Source Projects. arXiv:2505.15088
- Wang et al. (2023c) Yue Wang, Hung Le, Akhilesh Deepak Gotmare, Nghi D. Q. Bui, Junnan Li, and Steven C. H. Hoi. 2023c. CodeT5+: Open Code Large Language Models for Code Understanding and Generation. arXiv:2305.07922
- Wang et al. (2025f) Yifan Wang, Penghao Liu, and Yang Zhang. 2025f. VuL-MCBERT: A Vulnerability Detection Method Based on Self-Supervised Contrastive Learning. In International Conference on Artificial Intelligence, Big Data and Algorithms (CAIBDA). IEEE, Beijing, China, 923–927. doi:10.1109/CAIBDA65784.2025.11183103
- Wang et al. (2021) Yue Wang, Weishi Wang, Shafiq Joty, and Steven C. H. Hoi. 2021. CodeT5: Identifier-aware Unified Pre-trained Encoder-Decoder Models for Code Understanding and Generation. arXiv:2109.00859
- Wang et al. (2025c) Ziliang Wang, Ge Li, Jia Li, Yihong Dong, Yingfei Xiong, and Zhi Jin. 2025c. Line-level Semantic Structure Learning for Code Vulnerability Detection. In International Conference on Internetware (Internetware). ACM, Trondheim, Norway, 269–280. doi:10.1145/3755881.3755894
- Wang et al. (2025d) Ziliang Wang, Ge Li, Jia Li, Jia Li, Meng Yan, Yingfei Xiong, and Zhi Jin. 2025d. M2CVD: Enhancing Vulnerability Understanding through Multi-Model Collaboration for Code Vulnerability Detection. ACM Transactions on Software Engineering and Methodology (TOSEM) (2025). doi:10.1145/3771923 Just Accepted.
- Wang et al. (2025e) Ziliang Wang, Ge Li, Jia Li, Hao Zhu, and Zhi Jin. 2025e. VulAgent: Hypothesis-Validation based Multi-Agent Vulnerability Detection. arXiv:2509.11523
- Wartschinski et al. (2022) Laura Wartschinski, Yannic Noller, Thomas Vogel, Timo Kehrer, and Lars Grunske. 2022. VUDENC: Vulnerability Detection with Deep Learning on a Natural Codebase for Python. Information and Software Technology (INFSOF) 144 (2022), 106809. doi:10.1016/j.infsof.2021.106809
- Wei et al. (2022) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Fei Xia, Ed Chi, Quoc V Le, Denny Zhou, et al. 2022. Chain-of-Thought Prompting Elicits Reasoning in Large Language Models. In Advances in Neural Information Processing Systems (NeurIPS), Vol. 35. Curran Associates, Inc., New Orleans, LA, USA, 24824–24837.
- Wei et al. (2025) Zhiyuan Wei, Xiaoxuan Yang, Jing Sun, and Zijian Zhang. 2025. Distilling Lightweight Language Models for C/C++ Vulnerabilities. arXiv:2510.06645
- Wen et al. (2024a) Xin-Cheng Wen, Cuiyun Gao, Shuzheng Gao, Yang Xiao, and Michael R. Lyu. 2024a. SCALE: Constructing Structured Natural Language Comment Trees for Software Vulnerability Detection. In International Symposium on Software Testing and Analysis (ISSTA). ACM, Vienna, Austria, 235–247. doi:10.1145/3650212.3652124
- Wen et al. (2024b) Xin-Cheng Wen, Xinchen Wang, Yujia Chen, Ruida Hu, David Lo, and Cuiyun Gao. 2024b. VulEval: Towards Repository-Level Evaluation of Software Vulnerability Detection. arXiv:2404.15596
- Wen et al. (2023) Xin-Cheng Wen, Xinchen Wang, Cuiyun Gao, Shaohua Wang, Yang Liu, and Zhaoquan Gu. 2023. When Less is Enough: Positive and Unlabeled Learning Model for Vulnerability Detection. In International Conference on Automated Software Engineering (ASE). IEEE/ACM, Kirchberg, Luxembourg, 345–357. doi:10.1109/ASE56229.2023.00144
- Wen et al. (2025) Xin-Cheng Wen, Yijun Yang, Cuiyun Gao, Yang Xiao, and Deheng Ye. 2025. Boosting Vulnerability Detection of LLMs via Curriculum Preference Optimization with Synthetic Reasoning Data. arXiv:2506.07390
- Weyssow et al. (2025) Martin Weyssow, Chengran Yang, Junkai Chen, Yikun Li, Huihui Huang, Ratnadira Widyasari, Han Wei Ang, Frank Liauw, Eng Lieh Ouh, Lwin Khin Shar, and David Lo. 2025. R2Vul: Learning to Reason about Software Vulnerabilities with Reinforcement Learning and Structured Reasoning Distillation. arXiv:2504.04699
- Widyasari et al. (2024) Ratnadira Widyasari, David Lo, and Lizi Liao. 2024. Beyond ChatGPT: Enhancing Software Quality Assurance Tasks with Diverse LLMs and Validation Techniques. arXiv:2409.01001
- Widyasari et al. (2025) Ratnadira Widyasari, Martin Weyssow, Ivana Clairine Irsan, Han Wei Ang, Frank Liauw, Eng Lieh Ouh, Lwin Khin Shar, Hong Jin Kang, and David Lo. 2025. Let the Trial Begin: A Mock-Court Approach to Vulnerability Detection using LLM-Based Agents. arXiv:2505.10961
- Wu et al. (2025) Bozhi Wu, Chengjie Liu, Zhiming Li, Yushi Cao, Jun Sun, and Shang-Wei Lin. 2025. Enhancing Vulnerability Detection via Inter-procedural Semantic Completion. Proceedings of the ACM on Software Engineering (ASE) 2, ISSTA, Article ISSTA037 (2025), 23 pages. doi:10.1145/3728912
- Wu et al. (2023) Fangzhou Wu, Qingzhao Zhang, Ati Priya Bajaj, Tiffany Bao, Ning Zhang, Ruoyu ””Fish”” Wang, and Chaowei Xiao. 2023. Exploring the Limits of ChatGPT in Software Security Applications. arXiv:2312.05275
- Wu and Zhang (2024) Jie Wu and Cheng Zhang. 2024. A Software Vulnerability Detection Model Combined with Graph Simplification. In Asia-Pacific Artificial Intelligence and Big Data Forum (AIBDF). ACM, Ganzhou Shaanxi, China, 202–207. doi:10.1145/3718491.3718525
- Wu and Xiao (2025) Yifan Wu and Lei Xiao. 2025. Sparse-MoE: Syntax-Aware Multi-view Mixture of Experts for Long-Sequence Software Vulnerability Detection. In International Conference on Advanced Data Mining and Applications (ADMA). Springer, Kyoto, Japan, 348–362. doi:10.1007/978-981-95-3456-2_24
- Xia et al. (2025) Yifan Xia, Zichen Xie, Peiyu Liu, Kangjie Lu, Yan Liu, Wenhai Wang, and Shouling Ji. 2025. Beyond Static Pattern Matching? Rethinking Automatic Cryptographic API Misuse Detection in the Era of LLMs. Proceedings of the ACM on Software Engineering (PACMSE) 2, ISSTA, Article ISSTA006 (2025), 24 pages. doi:10.1145/3728875
- Xiong and Dong (2024) Zhe Xiong and Weiyu Dong. 2024. VulD-CodeBERT: CodeBERT-Based Vulnerability Detection Model for C/C++ Code. In International Conference on Communications, Information System and Computer Engineering (CISCE). IEEE, Guangzhou, China, 914–919. doi:10.1109/CISCE62493.2024.10653337
- Xu et al. (2025) Hanxiang Xu, Shenao Wang, Ningke Li, Kailong Wang, Yanjie Zhao, Kai Chen, Ting Yu, Yang Liu, and Haoyu Wang. 2025. Large Language Models for Cyber Security: A Systematic Literature Review. ACM Transactions on Software Engineering and Methodology (2025). doi:10.1145/3769676 Just Accepted.
- Xu et al. (2023) Wenlin Xu, Tong Li, Jinsong Wang, Haibo Duan, and Yahui Tang. 2023. Software Vulnerabilities Detection Based on a Pre-trained Language Model. In International Conference on Trust, Security and Privacy in Computing and Communications (TrustCom). IEEE, Exeter, UK, 904–911. doi:10.1109/TrustCom60117.2023.00129
- Xue et al. (2025) Shiying Xue, Lin Li, Tao Li, Haodong Chen, Jiapan Li, and Yangqing Qin. 2025. AIDetectVul: Software Vulnerability Detection Method Based on Feature Fusion of Pre-trained Models. In International Conference on Consumer Electronics and Computer Engineering (ICCECE). IEEE, Dongguan, China, 258–263. doi:10.1109/ICCECE65250.2025.10985370
- Yadav and Wilson (2024) Anurag Swarnim Yadav and Joseph N. Wilson. 2024. R+R: Security Vulnerability Dataset Quality Is Critical. In Annual Computer Security Applications Conference (ACSAC). IEEE, Honolulu, HI, USA, 1047–1061. doi:10.1109/ACSAC63791.2024.00086
- Yang et al. (2024b) An Yang, Baosong Yang, Binyuan Hui, Bo Zheng, Bowen Yu, Chang Zhou, Chengpeng Li, Chengyuan Li, Dayiheng Liu, Fei Huang, et al. 2024b. Qwen2 Technical Report. arXiv:2407.10671
- Yang et al. (2024a) Aidan Z. H. Yang, Haoye Tian, He Ye, Ruben Martins, and Claire Le Goues. 2024a. Security Vulnerability Detection with Multitask Self-Instructed Fine-Tuning of Large Language Models. arXiv:2406.05892
- Yang et al. (2025a) Xu Yang, Shaowei Wang, Jiayuan Zhou, and Wenhan Zhu. 2025a. One-for-All Does Not Work! Enhancing Vulnerability Detection by Mixture-of-Experts (MoE). Proceedings of the ACM on Software Engineering (PACMSE) 2, FSE, Article FSE021 (2025), 19 pages. doi:10.1145/3715736
- Yang et al. (2025b) Yixin Yang, Bowen Xu, Xiang Gao, and Hailong Sun. 2025b. Context-Enhanced Vulnerability Detection Based on Large Language Models. ACM Transactions on Software Engineering and Methodology (TOSEM) (2025). doi:10.1145/3779222 Just Accepted.
- Yang et al. (2025c) Yanjing Yang, Xin Zhou, Runfeng Mao, Jinwei Xu, Lanxin Yang, Yu Zhang, Haifeng Shen, and He Zhang. 2025c. DLAP: A Deep Learning Augmented Large Language Model Prompting Framework for Software Vulnerability Detection. Journal of Systems and Software (JSS) 219 (2025), 112234. doi:10.1016/j.jss.2024.112234
- Yao et al. (2023) Shunyu Yao, Jeffrey Zhao, Dian Yu, Nan Du, Izhak Shafran, Karthik Narasimhan, and Yuan Cao. 2023. ReAct: Synergizing Reasoning and Acting in Language Models. In International Conference on Learning Representations (ICLR). NSF PAR, Kigali, Rwanda, 1–33.
- Yıldırım et al. (2024) Recep Yıldırım, Kerem Aydın, and Orçun Çetin. 2024. Evaluating the Impact of Conventional Code Analysis Against Large Language Models in API Vulnerability Detection. In European Interdisciplinary Cybersecurity Conference (EICC). ACM, Xanthi, Greece, 57–64. doi:10.1145/3655693.3655701
- Yildiz et al. (2025) Alperen Yildiz, Sin G Teo, Yiling Lou, Yebo Feng, Chong Wang, and Dinil Mon Divakaran. 2025. Benchmarking LLMs and LLM-based Agents in Practical Vulnerability Detection for Code Repositories. In Annual Meeting of the Association for Computational Linguistics (ACL). Association for Computational Linguistics, Vienna, Austria, 30848–30865. doi:10.18653/v1/2025.acl-long.1490
- Yin et al. (2024) Xin Yin, Chao Ni, and Shaohua Wang. 2024. Multitask-Based Evaluation of Open-Source LLM on Software Vulnerability. IEEE Transactions on Software Engineering (TSE) 50, 11 (2024), 3071–3087. doi:10.1109/TSE.2024.3470333
- Yong et al. (2025) Heng Yong, Zhong Li, Minxue Pan, Tian Zhang, Jianhua Zhao, and Xuandong Li. 2025. GVI: Guided Vulnerability Imagination for Boosting Deep Vulnerability Detectors. In International Conference on Software Engineering (ICSE). IEEE/ACM, Ottawa, ON, Canada, 2867–2879. doi:10.1109/ICSE55347.2025.00214
- Yu et al. (2025) Jiaxin Yu, Peng Liang, Yujia Fu, Amjed Tahir, Mojtaba Shahin, Chong Wang, and Yangxiao Cai. 2025. An Insight into Security Code Review with LLMs: Capabilities, Obstacles, and Influential Factors. arXiv:2401.16310
- Yu et al. (2022) Weiye Yu, Guang Kou, Qihe Liu, Haoyu Zhang, and Qiang Wang. 2022. PATVD: Vulnerability Detection Based on Pre-training Techniques and Adversarial Training. In Smartworld, Ubiquitous Intelligence & Computing, Scalable Computing & Communications, Digital Twin, Privacy Computing, Metaverse, Autonomous & Trusted Vehicles (SmartWorld/UIC/ScalCom/DigitalTwin/PriComp/Meta). IEEE, Haikou, China, 1774–1781. doi:10.1109/SmartWorld-UIC-ATC-ScalCom-DigitalTwin-PriComp-Metaverse56740.2022.00253
- Yu et al. (2024) Xiao Yu, Lei Liu, Xing Hu, Jacky Wai Keung, Jin Liu, and Xin Xia. 2024. Fight Fire With Fire: How Much Can We Trust ChatGPT on Source Code-Related Tasks? IEEE Transactions on Software Engineering (TSE) 50, 12 (2024), 3435–3453. doi:10.1109/TSE.2024.3492204
- Yu and Zhang (2025) Yalong Yu and Cheng Zhang. 2025. A Vulnerability Detection Framework Based on Graph Decomposition Fusion and Augmented Abstract Syntax Tree. In International Conference on Big Data, Information and Computer Network (BDICN). ACM, Guangzhou, China, 734–738. doi:10.1145/3727353.3727471
- Yuan et al. (2022) Xue Yuan, Guanjun Lin, Yonghang Tai, and Jun Zhang. 2022. Deep Neural Embedding for Software Vulnerability Discovery: Comparison and Optimization. Security and Communication Networks 2022, 1 (2022), 5203217. doi:10.1155/2022/5203217
- Yusuf and Jiang (2024) Imam Nur Bani Yusuf and Lingxiao Jiang. 2024. Your Instructions Are Not Always Helpful: Assessing the Efficacy of Instruction Fine-tuning for Software Vulnerability Detection. arXiv:2401.07466
- Zeng et al. (2023) Peng Zeng, Guanjun Lin, Jun Zhang, and Ying Zhang. 2023. Intelligent Detection of Vulnerable Functions in Software through Neural Embedding-based Code Analysis. International Journal of Network Management (IJNM) 33, 3 (2023), e2198. doi:10.1002/nem.2198
- Zhang et al. (2024c) Boyu Zhang, Triet H. M. Le, and M. Ali Babar. 2024c. MVD: A Multi-Lingual Software Vulnerability Detection Framework. arXiv:2412.06166
- Zhang et al. (2023a) Junwei Zhang, Zhongxin Liu, Xing Hu, Xin Xia, and Shanping Li. 2023a. Vulnerability Detection by Learning From Syntax-Based Execution Paths of Code. IEEE Transactions on Software Engineering (TSE) 49, 8 (2023), 4196–4212. doi:10.1109/TSE.2023.3286586
- Zhang et al. (2024a) Quanjun Zhang, Chunrong Fang, Yang Xie, Yaxin Zhang, Yun Yang, Weisong Sun, Shengcheng Yu, and Zhenyu Chen. 2024a. A Survey on Large Language Models for Software Engineering. arXiv:2312.15223
- Zhang et al. (2025a) Shaobo Zhang, Qianzhi Wang, Qin Liu, Entao Luo, and Tao Peng. 2025a. VulTrLM: LLM-assisted Vulnerability Detection via AST Decomposition and Comment Enhancement. Empirical Software Engineering (EMSE) 31, 1 (2025), 1–36.
- Zhang et al. (2025b) Ting Zhang, Chengran Yang, Yindu Su, Martin Weyssow, Hung Nguyen, Tan Bui, Hong Jin Kang, Yikun Li, Eng Lieh Ouh, Lwin Khin Shar, and David Lo. 2025b. Benchmarking Large Language Models for Multi-Language Software Vulnerability Detection. arXiv:2503.01449
- Zhang et al. (2024b) Yingying Zhang, Zhengdan Jiang, Xingkai Cheng, Han Wu, Zonghang Tian, and Shuting Lou. 2024b. A Method of SQL Injection Attack Detection Based on Large Language Models. In International Conference on Computer Network Technology and Electronic and Information Engineering (CNTEIE). IEEE, Changchun, China, 154–158. doi:10.1109/CNTEIE66268.2024.00035
- Zhang et al. (2023b) Yuting Zhang, Jiahao Zhu, Yixin Yang, Ming Wen, and Hai Jin. 2023b. Comparing the Performance of Different Code Representations for Learning-based Vulnerability Detection. In Asia-Pacific Symposium on Internetware (Internetware). ACM, Hangzhou, China, 174–184. doi:10.1145/3609437.3609464
- Zhao et al. (2024c) Jiawei Zhao, Zhenyu Zhang, Beidi Chen, Zhangyang Wang, Anima Anandkumar, and Yuandong Tian. 2024c. GaLore: Memory-efficient LLM Training by Gradient Low-rank Projection. In International Conference on Machine Learning (ICML). JMLR, Vienna, Austria, Article 2528, 23 pages.
- Zhao et al. (2024a) Kunpeng Zhao, Shuya Duan, Ge Qiu, Jinyuan Zhai, Mingze Li, and Long Liu. 2024a. Python Source Code Vulnerability Detection Based on CodeBERT Language Model. In International Conference on Algorithms, Computing and Artificial Intelligence (ACAI). IEEE, Guangzhou, China, 1–6. doi:10.1109/ACAI63924.2024.10899694
- Zhao et al. (2024b) Yu Zhao, Lina Gong, Zhiqiu Huang, Yongwei Wang, Mingqiang Wei, and Fei Wu. 2024b. Coding-PTMs: How to Find Optimal Code Pre-trained Models for Code Embedding in Vulnerability Detection?. In International Conference on Automated Software Engineering (ASE). IEEE/ACM, Sacramento, CA, USA, 1732–1744. doi:10.1145/3691620.3695539
- Zhao et al. (2023) Yu Zhao, Lina Gong, Haoxiang Zhang, Yaoshen Yu, and Zhiqiu Huang. 2023. How to Get Better Embeddings with Code Pre-Trained Models? An Empirical Study. arXiv:2311.08066
- Zhao and Guan (2025) Ying Zhao and Xin Guan. 2025. Adversarial Training for Robustness Enhancement in LLM-Based Code Vulnerability Detection. In International Conference on Communications, Information System and Computer Engineering (CISCE). IEEE, Guangzhou, China, 1147–1152. doi:10.1109/CISCE65916.2025.11065803
- Zheng et al. (2024) Weining Zheng, Xiaohong Su, Hongwei Wei, and Wenxin Tao. 2024. SVulDetector: Vulnerability Detection based on Similarity using Tree-based Attention and Weighted Graph Embedding Mechanisms. Computers & Security (COSE) 144 (2024), 103930. doi:10.1016/j.cose.2024.103930
- Zheng et al. (2025) Xinran Zheng, Xingzhi Qian, Huichi Zhou, Shuo Yang, Yiling He, Suman Jana, and Lorenzo Cavallaro. 2025. Learning to Focus: Context Extraction for Efficient Code Vulnerability Detection with Language Models. arXiv:2505.17460
- Zheng et al. (2021) Yunhui Zheng, Saurabh Pujar, Burn Lewis, Luca Buratti, Edward Epstein, Bo Yang, Jim Laredo, Alessandro Morari, and Zhong Su. 2021. D2A: A Dataset Built for AI-Based Vulnerability Detection Methods Using Differential Analysis. In International Conference on Software Engineering: Software Engineering in Practice (ICSE-SEIP). IEEE/ACM, Madrid, Spain, 111–120. doi:10.1109/ICSE-SEIP52600.2021.00020
- Zhou et al. (2025a) Xin Zhou, Sicong Cao, Xiaobing Sun, and David Lo. 2025a. Large Language Model for Vulnerability Detection and Repair: Literature Review and the Road Ahead. ACM Transactions on Software Engineering and Methodology (TOSEM) 34, 5, Article 145 (2025), 31 pages. doi:10.1145/3708522
- Zhou et al. (2024a) Xin Zhou, Duc-Manh Tran, Thanh Le-Cong, Ting Zhang, Ivana Clairine Irsan, Joshua Sumarlin, Bach Le, and David Lo. 2024a. Comparison of Static Application Security Testing Tools and Large Language Models for Repo-level Vulnerability Detection. arXiv:2407.16235
- Zhou et al. (2024b) Xin Zhou, Ting Zhang, and David Lo. 2024b. Large Language Model for Vulnerability Detection: Emerging Results and Future Directions. In International Conference on Software Engineering: New Ideas and Emerging Results (ICSE-NIER). ACM/IEEE, Lisbon, Portugal, 47–51. doi:10.1145/3639476.3639762
- Zhou et al. (2019) Yaqin Zhou, Shangqing Liu, Jingkai Siow, Xiaoning Du, and Yang Liu. 2019. Devign: Effective Vulnerability Identification by Learning Comprehensive Program Semantics via Graph Neural Networks. In International Conference on Neural Information Processing Systems (NeurIPS). Curran Associates, Inc., Vancouver, BC, Canada, Article 915, 11 pages.
- Zhou et al. (2025b) Yuan Zhou, Enze Wang, and Shuoyoucheng Ma. 2025b. SSRFSeek: An LLM-based Static Analysis Framework for Detecting SSRF Vulnerabilities in PHP Applications. In International Seminar on Artificial Intelligence, Networking and Information Technology (AINIT). IEEE, Shenzhen, China, 939–944. doi:10.1109/AINIT65432.2025.11035424
- Zhu et al. (2022) Chenguang Zhu, Gewangzi Du, Tongshuai Wu, Ningning Cui, Liwei Chen, and Gang Shi. 2022. BERT-Based Vulnerability Type Identification with Effective Program Representation. In International Conference on Wireless Algorithms, Systems, and Applications (WASA). Springer, Dalian, China, 271–282. doi:10.1007/978-3-031-19208-1_23
- Zhu et al. (2025a) Hao Zhu, Jia Li, Cuiyun Gao, Jiaru Qian, Yihong Dong, Huanyu Liu, Lecheng Wang, Ziliang Wang, Xiaolong Hu, and Ge Li. 2025a. Specification-Guided Vulnerability Detection with Large Language Models. arXiv:2511.04014
- Zhu et al. (2024) Jin Zhu, Hui Ge, Yun Zhou, Xiao Jin, Rui Luo, and Yanchen Sun. 2024. Detecting Source Code Vulnerabilities Using Fine-Tuned Pre-Trained LLMs. In International Conference on Signal Processing (ICSP). IEEE, Xian, China, 238–242. doi:10.1109/ICSP62129.2024.10846595
- Zhu et al. (2025b) Qiang Zhu, Zhengying Wang, Yufei Zhuang, Guo Chen, He Luo, Meiqi Wu, Hongxiang Chen, and Zhi Liu. 2025b. An Automated Code Review Framework Based on BERT and Qianwen Large Model. In IEEE International Conference on Computer Communication and Artificial Intelligence (CCAI). IEEE, Haikou, China, 960–964. doi:10.1109/CCAI65422.2025.11189422
- Zibaeirad and Vieira (2024) Arastoo Zibaeirad and Marco Vieira. 2024. VulnLLMEval: A Framework for Evaluating Large Language Models in Software Vulnerability Detection and Patching. arXiv:2409.10756
- Zibaeirad and Vieira (2025) Arastoo Zibaeirad and Marco Vieira. 2025. Reasoning with LLMs for Zero-Shot Vulnerability Detection. arXiv:2503.17885
- Ziems and Wu (2021) Noah Ziems and Shaoen Wu. 2021. Security Vulnerability Detection Using Deep Learning Natural Language Processing. In Conference on Computer Communications Workshops (INFOCOM WKSHPS). IEEE, Virtual Conference, 1–6. doi:10.1109/INFOCOMWKSHPS51825.2021.9484500